### v1

# HyperPg - Prototypical Gaussians on the Hypersphere for Interpretable Deep Learning

Abstract
Prototype Learning methods provide an interpretable alternative to black-box deep learning models. Approaches such as ProtoPNet learn, which part of a test image “look like” known prototypical parts from training images, combining predictive power with the inherent interpretability of case-based reasoning. However, existing approaches have two main drawbacks: A) They rely solely on deterministic similarity scores without statistical confidence. B) The prototypes are learned in a black-box manner without human input. This work introduces HyperPg, a new prototype representation leveraging Gaussian distributions on a hypersphere in latent space, with learnable mean and variance. HyperPg prototypes adapt to the spread of clusters in the latent space and output likelihood scores. The new architecture, HyperPgNet, leverages HyperPg to learn prototypes aligned with human concepts from pixel-level annotations. Consequently, each prototype represents a specific concept such as color, image texture, or part of the image subject. A concept extraction pipeline built on foundation models provides pixel-level annotations, significantly reducing human labeling effort. Experiments on CUB-200-2011 and Stanford Cars datasets demonstrate that HyperPgNet outperforms other prototype learning architectures while using fewer parameters and training steps. Additionally, the concept-aligned HyperPg prototypes are learned transparently, enhancing model interpretability.
1 Introduction
Deep Learning has achieved high accuracy in many computer vision tasks. However, the decision-making processes of these models lack transparency and interpretability, making deployment in safety-critical areas challenging. Explainable Artificial Intelligence (XAI) seeks to develop interpretability methods to open the black-box reasoning processes of these models and increase trust in their decisions.
XAI methods can be broadly divided into two categories: First, Post-Hoc Methods like LIME (Ribeiro et al., 2016), SHAP (Lundberg & Lee, 2017) or GradCAM Selvaraju et al. (2017) offer explanations for predictions without requiring retraining. While applicable in many scenarios, post-hoc methods may not actually align with the models’ decision making processes, potentially leading to interpretations that are not entirely faithful (Rudin, 2019). Second, inherently interpretable methods provide built-in, case-based reasoning processes. For instance, small decision trees are inherently interpretable because their reasoning can be easily understood as a series of if-else statements (Molnar, 2020). However, they are constrained in their representational power.
Deep Prototype Learning Architectures such as ProtoPNet (Chen et al., 2019) and its derivatives (e.g., Rymarczyk et al., 2020; Donnelly et al., 2021; Sacha et al., 2023) integrate inherent interpretability into deep learning models through a prototype layer. Each neuron in this layer represents a prototype, storing a latent feature vector. The model’s predictions are based on the distances between sample features and prototype parameters, for example by computing the $L_{2}$-distance. However, these deterministic similarity scores do not include statistical information like confidence. To include such contextual information, prototypes could be modeled as probability distributions, like Gaussian distributions. Yet, recent prototype architectures favor the cosine hypersphere as feature space for its classification advantages (Mettes et al., 2019), where defining a Gaussian distribution analogue is challenging (Hillen et al., 2017).
The prototypes in ProtoPNet and its successors are learned in an opaque manner via backpropagation. After training, the model “pushes” the learned prototypes on the embedding of the most similar known training sample (Chen et al., 2019). However, these prototypes are learned without human control and do not incorporate domain knowledge. A new training regime is needed to aligned prototypes with human-defined concepts. The main contributions of this paper are as follows:
  • HyperPg: A new prototype representation with learned parameters anchor $\alpha$, mean $\mu$ and standard deviation $\sigma$. This representation models a Gaussian distribution over cosine similarities, thereby projecting a Gaussian distribution on the surface of a hypersphere. HyperPg’s similarity score is based on the Gaussian’s probability density function and adapts its size through a learned standard deviation.
  • HyperPgNet: A new prototype learning architecture built on HyperPg prototypes. HyperPgNet learns prototypes aligned with human-defined concepts using pixel level annotations. These concepts describe features such as color and patterns of birds or car parts.
  • Concept Extraction Pipeline: Based on Grounding DINO (Liu et al., 2023) and SAM2 (Ravi et al., 2024), this pipeline provides pixel-level concept annotations at scale.
  • Classification Experiments: HyperPgNet is compared with other prototype learning architectures like ProtoPNet on the CUB-200-2011 (Wah et al., 2011) and Stanford Cars (Krause et al., 2013) datasets. HyperPgNet outperforms the other models with fewer learned prototypes, while infusing each prototype with more meaning.
2 Related Work
Prototype Learning. In image classification, prototype learning approaches using autoencoders provide high interpretability by reconstructing learned prototypes from latent space back to the image space (Li et al., 2018). However, these approaches are limited in their performance because each prototype must represent the entire image. Alternatively, parts-based approaches like ProtoPNet (Chen et al., 2019; Donnelly et al., 2021) learn a predefined number of prototypical parts per class. Each class-specific prototype is an image patch represented by a $1\times 1$ spatial unit in the latent space. The network classifies images by summarizing the similarities of all the latent image patches to all prototypes using a fully connected layer. PIPNet (Nauta et al., 2023) aligns learned prototypes with human perception by optimizing the model to assign the same prototype to two different views of the same augmented image patch. ProtoPShare (Rymarczyk et al., 2020) merges similar prototypes after training is complete, while ProtoPool (Rymarczyk et al., 2022) directly learns class-shared prototypes. Other approaches additionally replace the linear output layer with other interpretable models. ProtoKNN (Ukai et al., 2023) classifies the similarity scores using a k-nearest neighbor (KNN) and ProtoTree Nauta et al. (2021) learn a decision tree.
HyperPgNet learns a predefined number of prototypes per concept, not per class. These concepts are provided as pixel-level annotations extracted by a new labeling pipeline based on the foundation models DINOv2 (Oquab et al., 2024) and SAM2 (Ravi et al., 2024). Similar to part-based prototypes, the concept prototypes span a $1\times 1$ spatial unit in the latent space. The prototypes are exclusive to each concept but shared among the classes.
Prototype learning is also applied to image segmentation. ProtoSeg (Sacha et al., 2023) uses prototypical parts a latent space with high spatial resolution for image segmentation. Another method proposes non-learnable prototypes (Zhou et al., 2022), which are trained by clustering in the latent space instead of being optimized via Backpropagation. ProtoGMM (Moradinasab et al., 2024) builds on this by updating class-specific prototypes like a Gaussian Mixture Model (GMM) using an Expectation-Maximization (EM) algorithm (Liang et al., 2022) with fixed mixture weights. MGProto (Wang et al., 2023) also learns mixture weights for their prototype GMM model.
HyperPgNet employs HyperPg, a novel prototype representation using Gaussian distributions on a hypersphere’s surface with learned parameters: direction $\bm{{\alpha}}$, mean $\mu$ and standard deviation (std) $\sigma$. Unlike ProtoGMM, which learns a multivariate Gaussian distribution in Euclidian space, each HyperPg prototype is defined by a directional vector $\bm{{\alpha}}$ and a $1D$-Gaussian distribution of cosine similarity values with mean $\mu$ and standard-deviation $\sigma$ around this vector. The use of cosine similarity for prototype learning has been shown to improve classification (Mettes et al., 2019) and has been widely adopted. (e.g., Donnelly et al., 2021; Rymarczyk et al., 2022; Sacha et al., 2023). HyperPgNet advances this by computing prototype activations as the probability density of each HyperPg prototype’s Gaussian distribution, allowing each prototype to adapt to the spread of the latent cluster.
Concept-Based Methods. Testing with Concept Activation Vectors (TCAV) (Kim et al., 2018) is a post-hoc method that aligns model activations with human-understandable concepts. In image classification, TCAV measures model activations on images representing a specific concept like “striped” and compares them to activations on a random image set. Concept Activation Regions (CAR) (Crabbé & van der Schaar, 2022) relax TCAV’s assumption of linear separability in latent space by using support vector machines and the kernel trick. Another post-hoc approach, CRAFT (Fel et al., 2023), generates concept attribution maps by identifying relevant concepts from the training set via Non-Negative Matrix Factorization (NMF) and backpropagating with GradCAM (Selvaraju et al., 2017).
Instead of analyzing the learned concepts post-hoc, some approaches integrate concept learning into the classifier. MCPNet (Wang et al., 2024) extracts concepts from multiple layers by splitting feature maps along the channel dimension, learning concepts during training and at multiple scales. Concept Bottleneck Models (CBMs) (Koh et al., 2020) augment black-box models to predict pre-defined concept labels in an intermediate step. In this manner, CBMs behave like two sequential black box models and their internal reasoning remains opaque.
HyperPgNet combines the concept-aligned learning approach of CBM with interpretable prototype learning. HyperPgNet learns multiple HyperPg prototypes per concept, instead of a fully connected concept prediction layer. A novel concept extraction pipeline, built on foundation models, provides pixel level annotations at scale. By using those annotations to learn concept prototypes, HyperPgNet avoids the ambiguity of aligning post-hoc explanations with the model’s internal reasoning.
3 Prototypical Gaussians on the Hypersphere
Figure 1: Illustration on how the different prototype formulations compute the similarity between a prototype ${\bm{p}}$ and latent vector ${\bm{z}}$. $L_{2}$ prototypes compute the Euclidian distance between two points in latent space. Hyperspherical prototypes use the cosine similarity of normalized vectors, which corresponds to the angle between two points on a hypersphere. Gaussian prototypes model a Gaussian distribution in Euclidian space and compute a probability density function (PDF). HyperPg prototypes learn a Gaussian distribution of cosine similarities, thereby projecting a Gaussian distribution onto the surface of a hypersphere.
Prototype Learning is an inherently interpretable machine learning method. The reasoning process is based on the similarity scores of the inputs to the prototypes, retained representations of the training data. For example, a K-Nearest Neighbor (KNN) model is a prototype learning approach with the identity function for representation and an unlimited number of prototypes. In contrast, a Gaussian Mixture Model (GMM) uses a mean representation but restricts the number of prototypes to the number of mixture components.
Prototype learning for deep neural networks involves finding structures in latent space representations. This section provides an overview of existing methods, which are also illustrated in Fig. 1. Prior work uses point-based prototypes, computing similarity scores relative to a single point in latent space. Aligning prototypes with human-labeled concepts requires a more powerful prototype representation, leading to the introduction of HyperPg - Prototypical Gaussians on the Hypersphere. HyperPg prototypes are able to adapt their shape in the latent space by modeling Gaussian distributions with mean and standard deviation, adapting to the variance in encoded concept features.
3.1 Point Based Prototypes
The general formulation of prototypes, as defined in previous work (e.g., Chen et al., 2019), is discussed first. Let $\mathcal{D}=[{\bm{X}},{\bm{Y}}]=\{({\bm{x}}_{i},y_{i})\}_{i=1}^{N}$ denote the training set, e.g., a set of labeled images, with classes $C$. Each class $c\in C$ is represented by $Q$ many prototypes ${\bm{P}}_{c}=\{{\bm{p}}_{c,j}\}_{j=1}^{Q}$.
Some feature encoder $\mathrm{Enc}$ projects the inputs into a $D$-dimensional latent space $\mathcal{Z}$, with
${\bm{z}}_{i}=\mathrm{Enc}({\bm{x}}_{i})$ being a feature map of shape $\zeta_{w}\times\zeta_{h}\times D$ with spatial size $\zeta=\zeta_{w}\zeta_{h}$. Commonly, the prototypes ${\bm{p}}$ are also part of $\mathcal{Z}$ with shape $\rho_{w}\times\rho_{h}\times D$, i.e., spatial size $\rho=\rho_{w}\rho_{h}$.
Autoencoder approaches use $\rho=\zeta$, meaning the prototype represents the entire image and can be reconstructed from latent space (Li et al., 2018). Part-based approaches like ProtoPNet and segmentation models like ProtoSeg use $\rho=1$ (Chen et al., 2019; Zhou et al., 2022), meaning each prototype represents some part of the image. Notable exceptions include Deformable ProtoPNet (Donnelly et al., 2021), where each prototype has spatial size $\rho=3\times 3$ and MCPNet (Wang et al., 2024), where prototypes are obtained by dividing the channels of the latent space into chunks.
The prediction is computed by comparing each prototype ${\bm{p}}$ to the latent feature map ${\bm{z}}$. For simplicity’s sake lets assume the spatial dimensions $\rho=\zeta=1$. The following equations can be adapted for higher spatial dimensions by summing over $\sum_{\rho_{w}}\sum_{\rho_{h}}$ for each chunk of the latent map.
ProtoPNet’s prototypes leverage the $L_{2}$ similarity. The $L_{2}$ similarity measure is defined as
$s_{L_{2}}({\bm{z}}|{\bm{p}})=\log\left(\frac{\|{\bm{z}}-{\bm{p}}\|^{2}_{2}+1}{\|{\bm{z}}-{\bm{p}}\|^{2}_{2}+\epsilon}\right)$ (1)
and is based on the inverted $L_{2}$ distance between a latent vector ${\bm{z}}$ and a prototype vector ${\bm{p}}$. This similarity is a point-based measure, as only the two vectors are compared, without any additional context like the expected variance of the cluster represented by the prototype.
Hyperspherical prototypes using the cosine similarity have been shown to perform well in classification tasks (Mettes et al., 2019) and have been widely used since (e.g., Zhou et al., 2022; Ukai et al., 2023). The cosine similarity is defined as
$s_{\cos}({\bm{z}}|{\bm{p}})=\frac{{\bm{z}}^{\top}{\bm{p}}}{\|{\bm{z}}\|_{2}\|{\bm{p}}\|_{2}},$ (2)
which is based on the angle between two normalized vectors of unit length. By normalizing $D$ dimensional vectors to unit length, they are projected onto the surface of a $D$ dimensional hypershere. The cosine similarity is defined on the interval $[-1,1]$ and measures: $1$ for two vectors pointing in the same direction, 00 for orthogonal vectors, and $-1$ for vectors pointing in opposite directions. Like the $L_{2}$ similarity, the cosine similarity is a point-based measure comparing only two vectors.
Both the $L_{2}$ and cosine similarity can be used for classification. The similarity scores are processed by a fully connected layer (e.g., Chen et al., 2019; Donnelly et al., 2021), or a winner-takes-all approach assigns the class of the most similar prototype (e.g., Sacha et al., 2023). Prototypes can be learned by optimizing a task-specific loss, such as cross-entropy, via backpropagation. Alternatively, Zhou et al. (2022) propose “non-learnable” prototypes, whose parameters are obtained via clustering in the latent space rather than backpropagation.
3.2 Gaussian Prototypes
Gaussian prototypes model prototypes as a Gaussian distribution with mean and covariance. They adapt to the spread of the associated latent cluster by adjusting their covariance matrix. Thus, a Gaussian prototype with a wide covariance can still have a relatively high response even for larger distances from the mean vector.
Let the formal definition of a Gaussian prototype be ${\bm{p}}^{G}=(\bm{{\mu}},\bm{\mathrm{\Sigma}})$. The parameters of ${\bm{p}}^{G}_{c,j}$ now track both the mean and covariance of latent vector distribution ${\bm{Z}}_{c,j}$. Each Gaussian prototype ${\bm{p}}^{G}$ thus defines a multivariate Gaussian Distribution ${\mathcal{N}}(\bm{{\mu}},\bm{\mathrm{\Sigma}})$. Gaussian prototypes can be trained using EM for clustering in the latent space (Zhou et al., 2022; Wang et al., 2023; Moradinasab et al., 2024) or by directly optimizing the parameters via Backpropagation. The similarity measure of Gaussian prototypes is defined as the probability density function (PDF) for $D$-dimensional multivariate Gaussians, namely
$s_{\mathrm{Gauss}}({\bm{z}}|{\bm{p}}^{G})$ $=\mathcal{N}({\bm{z}};\bm{{\mu}},\bm{\mathrm{\sigma}})$ (3)
This formulation as a PDF has several advantages: A) The similarity can be interpreted as the likelihood of being sampled from the Gaussian prototypes, which is more meaningful than a distance metric in a high-dimensional latent space. B) Prototypes can adapt their shape using a full covariance matrix, allowing different variances along various feature dimensions, offering more flexibility in shaping the latent space. However, this increases computational requirements, especially with EM clustering. Gaussians prototypes lose the advantages of hyperspherical prototypes for classification and regression Mettes et al. (2019).
3.3 Gaussian Prototypes on the Hypersphere - HyperPg
Prototypical Gaussians on the Hypersphere (HyperPg) combine the advantages of Gaussian and hyperspherical prototypes. HyperPg prototypes are defined as ${\bm{p}}^{H}=(\bm{{\alpha}},\mu,\sigma)$ with a directional anchor vector $\bm{{\alpha}}$, scalar mean similarity $\mu$ and scalar standard deviation (std) $\sigma$. HyperPg prototypes learn a $1D$ Gaussian distribution over the cosine similarities to the anchor vector $\bm{{\alpha}}$. Because the cosine similarity is bounded to $[-1,1]$, HyperPg’s similarity measure is defined as the PDF of the truncated Gaussian distribution within these bounds. Let $\mathcal{G}(x,\mu,\sigma)$ be the cumulative Gaussian distribution function. Then HyperPg’s similarity measure based on the truncated Gaussian distribution is defined as
$s_{\mathrm{HyperPg}}({\bm{z}}|{\bm{p}}^{H})$ $=\mathcal{T}_{G}(s_{\cos}({\bm{z}}|\bm{{\alpha}});\mu,\sigma,-1,1)$ (4)
$=\frac{{\mathcal{N}}\left(s_{\cos}({\bm{z}}|\bm{{\alpha}});\mu,\sigma\right)}{\mathcal{G}(1,\mu,\sigma)-\mathcal{G}(-1,\mu,\sigma).}$ (5)
(a) $\mu=1$
(b) $\mu=0.6$
(c) $\mu=0$
(d) $\mu=-1$
Figure 2: HyperPg learns a Gaussian distribution with mean $\mu$ and std $\sigma$ of cosine similarities with the anchor $\bm{{\alpha}}$. The plots visualize HyperPg’s activations on the $3D$ hypersphere with anchor $\bm{{\alpha}}=(0,0,1)$, std $\sigma=0.1$ and and various mean cosine similarities $\mu\in[-1,1]$. The anchor $\bm{{\alpha}}$ is shown as a red arrow, and std $\sigma$ governs the width of the distribution. For $\mu=1$, the distribution is aligned with the vector; for $\mu=-1$ it is on the opposite side of the hypersphere. Due to the cosine similarity, setting $11>\mu>-$ 1 interpolates the area of highest probability density between both poles, leading to a ring shape on the hypersphere’s surface.
Fig. 2 illustrates the activations of HyperPg’s similarity function on the surface of a $3D$ hypersphere with anchor $\bm{{\alpha}}=(0,0,1)$, fixed std $\sigma=0.1$ and various mean values $\mu\in[-1,1]$. The anchor $\bm{{\alpha}}$ defines a prototypical direction vector in latent space $\mathcal{Z}$, similar to other hyperspherical prototypes, and is visualized as a red arrow. The learned Gaussian distribution of cosine similarities is projected onto the hypersphere’s surface with std $\sigma$ governing the spread of the distribution, and mean $\mu$ the expected distance to the anchor $\bm{{\alpha}}$. For $\mu=1$, the distribution centers around the anchor, as the cosine similarity is $1$ if two vectors point in the same direction. For $\mu=-1$, the distribution is the on opposite side of the hypersphere, as the cosine similarity is $-1$ for vectors pointing in opposite directions.
For values of $11>\mu>-$ 1, the distribution forms a hollow ring around the anchor vector $\bm{{\alpha}}$. This occurs because the cosine similarity for these $\mu$ values expects the activating vectors to point in a different direction than the anchor, without specifying the direction. Imagining the interpolation between $\mu=1$ and $\mu=-1$, the probability mass moves from one pole of the hypersphere to the other, stretching like a rubber band over the surface. For $\mu=0$, the expected cosine similarity indicates that vectors with the highest activation are orthogonal to the anchor $\bm{{\alpha}}$. Since no specific direction is indicated, the entire hyperplanar segment orthogonal to the anchor has the highest activation. This activation pattern for the cosine similarity would typically require an infinite mixture of prototype vectors pointing in all directions in this hyperplane. HyperPg achieves the same effect by learning only one prototype vector (the anchor) and just two additional scalar parameters. This significantly increases HyperPg’s representational power compared to standard hyperspherical and Gaussian prototypes, as a single HyperPg prototype can cover multiple clusters simultaneously.
Combining the strengths of hyperspherical and Gaussian prototypes, HyperPg can learn more complex structures in the latent space, such as human-defined concepts, and provide a more meaningful similarity measure. HyperPg can be easily adapted to other probability distributions with additional desirable properties. Possible candidate distributions are elaborated on in Appendix A. Similarly, it is possible to exchange the cosine similarity to other similarity measures or functions, and learn an untruncated PDF over their output, making the HyperPg idea transferable to applications outside of prototype learning.
4 Training
The original ProtoPNet implementation uses three loss functions: a task specific loss like crossentropy for classification, a cluster loss to increase compactness within a class’s cluster, and a separation loss to increase distances between different prototype clusters. However, the prototypes of ProtoPNet and its successors are learned in a black-box manner.
HyperPgNet is a new inherently interpretable deep learning approach built on HyperPg prototypes. It introduces a “Right for the Right Concept” loss, inspired by “Right for the Right Reasons” (Ross et al., 2017), to restrict the learned prototypes to human-defined concepts. This focus enhances the interpretability and minimizes the influence of confounding factors. This section first provides an overview of the used prototype learning losses, then introduces the Right for the Right Concept loss, and finally discusses the overall network architecture and final multi-objective loss.
4.1 Prototype Losses
ProtoPNet defines a cluster loss function to shape the latent space such that all latent vectors ${\bm{z}}_{c}\in{\bm{Z}}_{c}$ with class label $c$ are clustered tightly around the semantically similar prototypes ${\bm{p}}_{c}\in{\bm{P}}_{c}$. The cluster loss function is defined as
$L_{\mathrm{Clst}}=-\frac{1}{N}\sum_{i=1}^{N}\frac{1}{|C|}\sum_{c\in C}\max_{{\bm{p}}_{c}\in{\bm{P}}_{c}}\max_{{\bm{z}}_{c,i}\in{\bm{Z}}_{c,i}}s({\bm{p}}_{c},{\bm{z}}_{c,i}),$ (6)
where $s(\cdot,\cdot)$ is some similarity measure. The $L_{\mathrm{Clst}}$-Loss function increases compactness by increasing the similarity between prototypes ${\bm{p}}_{c}$ of class $c$ latent embeddings ${\bm{z}}_{c}$ of class $c$ over all samples.
An additional separation loss increases the margin between different prototypes. The separation loss function is defined as
$L_{\mathrm{Sep}}=\frac{1}{N}\sum_{i=1}^{N}\frac{1}{|C|}\sum_{c\in C}\max_{{\bm{p}}_{\neg c}\notin{\bm{P}}_{c}}\max_{{\bm{z}}_{c,i}\in{\bm{Z}}_{c,i}}s({\bm{p}}_{\neg c},{\bm{z}}_{c,i}),$ (7)
The $L_{\mathrm{Sep}}$ function punishes high similarity values between a latent vector ${\bm{z}}_{c}$ of class c and prototypes ${\bm{p}}_{\neg c}$ not belonging to $c$, thereby separating the clusters in latent space. Please note, ProtoPNet (Chen et al., 2019) use a slightly different notation by working with the $L_{2}$ distance, instead of a similarity measure.
In HyperPgNet the learned prototypes are not class exclusive, but shared among different classes. Instead, each prototype is assigned to one human-defined concept $k\in K$. The Concept Activation Regions (CAR, Crabbé & van der Schaar, 2022) model proposes a concept density loss defined as
$L_{\mathrm{Den}}$ $=-\frac{1}{|K|}\sum_{k\in K}\frac{1}{|{\bm{P}}_{k}|}\sum_{{\bm{p}}_{k}\in{\bm{P}}_{k}}\phi({\bm{p}}_{k},{\bm{Z}}_{k})-\phi({\bm{p}}_{k},{\bm{Z}}_{\neg k})$ (8)
$\phi({\bm{p}},{\bm{Z}})$ $=\frac{1}{|{\bm{Z}}|}\sum_{{\bm{z}}\in{\bm{Z}}}s({\bm{p}},{\bm{z}}).$ (9)
The density loss uses the similarity aggregation $\phi({\bm{p}},{\bm{Z}})$ which computes the mean response of a prototype ${\bm{p}}$ with a set of latent features ${\bm{Z}}$. The density loss therefore computes the mean response of correctly assigned prototypes minus the mean response of incorrectly assigned prototypes. The loss functions proposed by ProtoPNet compute the maximum response of correctly and incorrectly assigned prototypes instead.
4.2 Right Concept Loss
To ensure prototypes actually correspond to the input pixels containing the concept, and do not respond to other factors in the background, HyperPgNet introduces the “Right for the Right Concept” (RRC) -Loss inspired by “Right for the Right Reasons” (RRR, Ross et al., 2017). The RRC loss is defined as
$L_{\mathrm{RRC}}=\frac{1}{N}\sum_{i=1}^{N}\sum_{k\in K}\left({\bm{A}}_{{\bm{x}}_{i},k}\frac{\partial}{\partial{\bm{x}}_{i}}\sum_{{\bm{p}}_{k}\in{\bm{P}}_{k}}s\left({\bm{p}}_{k},\mathrm{Enc}\left({\bm{x}}_{i}\right)\right)\right)^{2},$ (10)
with binary annotation matrix ${\bm{A}}_{{\bm{x}}_{i},k}\in\{0,1\}^{N\times\times W\times H}$ for each input sample ${\bm{x}}_{i}$ and concept $k$. This annotation matrix defines for each input image, which pixels contain which concept. In the original RRR paper, the annotation matrix specified relevant regions for the classification task, steering the model’s activations away from confounding factors in the background. In HyperPgNet the RRC-Loss further strengthens the prototype-concept association.
4.3 Network Architecture
Figure 3: HyperPgNet Architecture. The HyperPg module can be easily exchanged to other prototype formulations such as ProtoPNet. HyperPgNet uses the truncated Gaussian distribution as density estimator, but other PDFs are possible.
Fig. 3 illustrates HyperPgNet’s Architecture for interpretable image classification. HyperPgNet uses a pretrained, off-the-shelf feature encoder as a backbone, such as the convolution pyramid from a CNN like ConvNeXt-tiny (Liu et al., 2022) or a transformer architecture like Segformer (Xie et al., 2021). The high-dimensional feature maps provided by the backbone are passed through a neck, consisting of two $1\times 1$ convolution layers with ReLU activation in between, projecting into a smaller latent space with $256$ channels, followed by Layer Normalization. The hyperparameter of 256 channels aligns with previous work (e.g., Rymarczyk et al., 2022).
HyperPg’s similarity function computes the truncated Gaussian PDF over the cosine similarity of the latent feature maps with the HyperPg prototype anchors $\bm{{\alpha}}$. HyperPgNet implements the prototype learning part of the network as a module with two layers: First, a prototype learning layer computes the cosine similarity of learnable HyperPg prototype anchors $\alpha$ to the latent feature map $z$, similar to other prototype learning architectures. Second, the new Density Estimation layer with learnable parameters mean $\mu$ and std $\sigma$ computes the truncated Gaussian PDF over the activations of the previous layer.
This implementation of the HyperPg Module with two internal layers supports swapping out the prototype similarity measure to other measures in the future. Additionally, the Density Estimation Layer concept could be applied to other architectures to learn a Gaussian distribution over the previous layer’s activations. Finally, the output of the HyperPg Module is passed through a single fully connected layer (FCL) to produce the output logits or class scores.
HyperPgNet is trained for image classification with a multi-objective loss function. This loss function is defined as
$L=L_{\mathrm{CE}}+\lambda_{\mathrm{Den}}L_{\mathrm{Den}}+\lambda_{\mathrm{RRC}}L_{\mathrm{RRC}},$
where $L_{\mathrm{CE}}$ is the Crossentropy loss over network predictions and ground truth image labels. By choosing different values for $\lambda$, the model’s tradeoff between task performance and interpretability can be adjusted.
5 Concept Extraction at Scale
To learn prototypes based on human-defined concepts with RRC-Loss for image classification, pixel level annotations are required. Foundation models are leveraged to generate these annotations, reducing the need for human labeling.
The CUB-200-2011 dataset (Wah et al., 2011) contains 11,788 images of birds across 200 sub-species. Each image is annotated with up to 15 part locations, such as head, tail and back. The dataset includes 312 binary attributes describing patterns and colors of these parts, or additional shape information like “duck-like” or “owl-like”. For concept learning, 30 attributes are manually selected, excluding small concepts like eye-color or class exclusive attributes like “duck-like”.
To extract pixel level annotation masks, the point annotations for the attributes are given to a Segment Anything 2 (SAM2) model(Ravi et al., 2024). Using the point locations belonging to other attributes as negative points, segmentation masks are generated (see Fig. 4). The segmentation mask is then assigned the attribute from the dataset annotations to define a concept.
(a) Concept Segmentation Masks for CUB-200-2011 based on provided positive and negative part annotations.
(b) Concept Segmentation Mask for Stanford Cars generated from a list of human-defined car parts.
Figure 4: Examples for Automated Concept Extraction.
The Stanford Cars dataset (Krause et al., 2013) does not provide part annotations. This dataset contains 16,185 images of 196 car classes, categorized by make, model and year (e.g., BMW M3 coupe 2012). The dataset provides only bounding boxes for the main subject matter, useful for foreground-background separation. To obtain pixel-level annotations, a list of 10 car parts (e.g., “wheel”, “headlight”, “radiator”) is defined. A Grounding DINO model (Liu et al., 2023) uses this list to generate bounding boxes for parts in the image. A post-processing step removes possible outliers by eliminating bounding boxes outside the original bounding box provided by the dataset.Finally, SAM2 generates pixel-level segmentation masks (see Fig. 4). For Stanford Cars, only the part label is used as the concept label as the dataset does not provide additional part attributes. The automated pipeline labeled the entire Stanford Cars dataset within 2h on an NVidia 4060ti with 16GB VRAM and can be easily adapted to other datasets.
6 Experiments
The experiments were performed on two datasets: CUB-200-2011 bird species classification dataset (Wah et al., 2011) and the Stanford Cars dataset (Krause et al., 2013). Two variants of HyperPgNet were tested, with and without the $L_{\mathrm{RRC}}$ loss. HyperPgNet was compared to ProtoPNet with both standard $L_{2}$ and the newly introduced HyperPg prototypes, and Concept Bottleneck Models (CBM) with joint training scheme (Koh et al., 2020). Segformer and ConvNext served as Transformer and CNN baselines, respectively. All models were implemented with pretrained Segformer backbones in Pytorch. However, HyperPgNet’s performance is independent of the chosen backbone architecture (see Appendix B.1).
In contrast to prior work, the training and testing was done on full images, without crops to the bounding boxes provided by the datasets. For data augmentation RandomPerspective, RandomHorizontalFlip, and RandomAffine were used in an online manner. Thus, for CUB-200-2011, each class has 30 training images, instead of the offline augmentation to 1200 images used in prior work (e.g., Chen et al., 2019; Rymarczyk et al., 2020; Ukai et al., 2023). All models were trained with batch size 96 until convergence. The experiments were performed on a workstation with a single NVidia RTX 3090 GPU (24GB VRAM) per model.
6.1 Quantitative Results
Table 1: Test Top-1 Accuracy with Segformer Backbone. Models are grouped by their reasoning process. BB: Black Box. PP: Prototypical Parts. CAP: Concept Aligned Prototypes.
CUB-200-2011 Stanford Cars
# Prototypes Accuracy [%] # Prototypes Accuracy [%]
BB Segformer Baseline - 17.7 - 1.9
ConvNeXt Baseline - 74.2 - 57.5
CBM - 75.7 - 79.9
PP ProtoPNet 2000 68.0 1960 86.4
ProtoPNet + HyperPg 2000 70.5 1960 87.4
CAP HyperPgNet - $L_{\mathrm{RRC}}$ 300 76.5 180 88.6
HyperPgNet + $L_{\mathrm{RRC}}$ 300 74.1 180 81.2
Figure 5: Test Accuracy with Segformer Backbone.
Table 1 shows the top-1 accuracy for models on the CUB-200-2011 (Birds) and Stanford Cars (Cars) datasets using the Segformer backbone. The Test Accuracy per Epoch curves are visualized in Fig. 5. The Segformer baseline overfit to the small training datasets (17.7% and 1.9% test accuracy), while the ConvNeXt baseline performed better (74.2% and 57.5%). CBM achieved high results quickly, scoring 75.7% on Birds and 79.9% on Cars. ProtoPNet required the most training time, converging only after about 490 epochs, scoring 68% on Birds and 86.4% on Cars. Switching ProtoPNet’s prototypes from $L_{2}$ to HyperPg improved both training speed and accuracy, converging in 200 epochs with 70.5% and 87.4% test accuracy.
HyperPgNet further improved on this by learning human-defined concepts. HyperPgNet with concept aligned prototypes, but without $L_{\mathrm{RRC}}$, outperformed the other models. It achieved 76.5% accuracy on Birds and 88.6% on Cars after only 40 training epochs. With the “Right for the Right Concept” loss, the model retained the high learning speed and performance, although the test accuracy dropped slightly to 74.1% and 81.2%. This performance loss is offset by increased transparency due to more precise concept prototypes. For additional ablation studies, see Appendix B.
6.2 Qualitative Results
HyperPg prototypes do not only enhance the models learning speed and task performance, but also the interpretabilty. Prior work built activations by overlaying the up-sampled latent space activation over the image. Inspired by GradCAM, Fig. 6 shows the gradient of the prototype activations backpropagated to the input image. Class-based prototypes focus on regions relevant for the prediction, but do not follow human defined concepts. With concept-alignment, the prototype focuses more towards the white head of the bird. But as the concept alignment is only encouraged by $L_{\mathrm{Clst}}$ and $L_{\mathrm{Sep}}$ the prototype still activates on other bright areas on the birds body. With the addition of $L_{\mathrm{RRC}}$ to the training regime, the prototypes are forced to focus mainly on the annotated region. Thus, the head prototype ignores the wing area, which is learned by a different prototype. The overall learned concept prototypes fit more precisely to the images’ annotated areas. For additional examples and latent space analyses, please refer to Appendix C.
(a) ProtoPNet + HyperPg
(b) HyperPgNet - $L_{\mathrm{RRC}}$
(c) HyperPgNet + $L_{\mathrm{RRC}}$
Figure 6: Prototype Gradient Maps for different HyperPgNet models.
7 Conclusion
This work introduces HyperPg, a new prototype representation learning a probability distribution on the surface of a hypersphere in latent space. HyperPg adapt to the variance of clusters in latent space and improve training time and accuracy compared to other prototype formulations. HyperPgNet leverages HyperPg to learn human defined concept prototypes, instead of black-box optimized prototypes. The combination of probabilistic prototypes on the hypersphere and concept aligned prototypes allows HyperPgNet to outperform other prototype learning approaches. One limitation HyperPgNet faces are increased computational requirements due to the inclusion of concept annotations and $L_{\mathrm{RRC}}$ during training. However, this is offset by an acceleration of the training process regarding epochs and number of samples. Coupled with the increased transparency and interpretability of the model, this makes HyperPgNet a strong contender for scenarios with few training samples and high safety requirements, like medical applications or human-robot interaction.
Acknowledgments
This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 448648559
Appendix A Adapting HyperPg to other Probability Distributions
Subsection 3.3 defines HyperPg prototypes ${\bm{p}}^{H}=(\bm{{\alpha}},\mu,\sigma)$ as a Gaussian Distribution with mean $\mu$ and std $\sigma$ of cosine similarities around an anchor vector $\alpha$. This idea of learning a distribution of cosine similarity values around an anchor $\bm{{\alpha}}$ can be adapted to other distributions. This sections introduces some potential candidates. As early experiments on the CUB-200-2011 dataset showed no significant difference in performance, these sections are relegated to the appendix.
A.1 Cauchy Distribution
One theoretical disadvantage of the Gaussian distribution is the fast approach to zero, which is why a distribution with heavier tails such as the Cauchy distribution might be desirable. The Cauchy distribution’s PDF is defined as
$\mathcal{C}(x;x_{0},\gamma)=\frac{1}{\pi\gamma\left(1+\left(\frac{x-x_{0}}{\gamma}\right)^{2}\right)},$ (11)
with median $x_{0}$ and average absolute deviation $\gamma$. The HyperPg prototypes with Cauchy are defined as accordingly as ${\bm{p}}^{\mathrm{Cauchy}}=(\bm{{\alpha}},x_{0},\gamma)$.
Figure 7: PDF for the Gaussian and Cauchy distribution of cosine similarity values. The Cauchy distribution has heavier tails, avoiding vanishing gradients issues.
Fig. 7 illustrates the PDF of the Gaussian and Cauchy distributions with $\mu=x_{0}=1$ and $\sigma=\gamma=0.2$, i.e., the main probability mass is aligned with the anchor $\bm{{\alpha}}$. The Gaussian distributions PDF quicly approaches zero and stays near constant. This could potentially cause vanishing gradient issues during training. The heavier tails of the Cauchy distribution ensure that for virtually the entire value range of the cosine similarity, gradients could be propagated back through the model. However, experiments on CUB-200-2011 showed no significant performance difference between using HyperPg with the Gaussian or Cauchy distribution.
A.2 Truncated Distributions
The cosine similarity is defined only on the interval $[-1,1]$. This makes it attractive to also use truncated probability distributions, which are also only defined on this interval. The truncation imposes a limit on the range of the PDF, thereby limiting the influence of large values for the distribution’s $\sigma$ or $\gamma$ parameter, respectively. The truncated Gaussian pdf $\mathcal{T}_{\mathrm{Gauss}}$ requires the cumulative probability function $\mathcal{G}$ and error function $f_{\mathrm{err}}$, and is defined as
[…]
${\mathcal{N}}(x;\mu,\sigma)$ $=\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp\left(-\frac{(x-\mu)^{2}}{2\sigma^{2}}\right),$ (13)
$\mathcal{G}(x,\mu,\sigma)$ $=\frac{1}{2}\left(1+f_{\mathrm{err}}\left(\frac{x-\mu}{\sigma\sqrt{2}}\right)\right),$ (14)
$\mathcal{T}_{\mathrm{Gauss}}(x,\mu,\sigma,a,b)$ $=\frac{{\mathcal{N}}\left(s_{\cos}({\bm{z}}|\bm{{\alpha}});\mu,\sigma\right)}{\mathcal{G}(1,\mu,\sigma)-\mathcal{G}(-1,\mu,\sigma).},$ (15)
with lower bound $a$ and upper bound $b$, e.g., for the cosine similarity $a=-1$ and $b=1$. Similarly, the truncated Cauchy distribution can be applied, which is defined as
$\mathcal{T}_{\mathrm{Cauchy}}(x,x_{0},\gamma,a,b)=\frac{1}{\gamma}\left(1+\left(\frac{x-x_{0}}{\sigma}\right)^{2}\right)^{-1}\left(\arctan\left(\frac{b-x_{0}}{\gamma}\right)-\arctan\left(\frac{a-x_{0}}{\gamma}\right)\right)^{-1}.$ (16)
Fig. 8 shows the test accuracy of three HyperPgNet models with Gaussian, truncated Gaussian and truncated Cauchy distribution on the CUB-200-2011 dataset. While difference in test performance and learning speed were minimal on the CUB-200-2011 dataset, further exploration is necessary, as other experiments showed that the concept-alignment on CUB-200-2011 dominates the learning process, lessening the influence of the prototypes.
Figure 8: CUB-200-2011 Test Accuracy of HyperPgNet with different probability distributions. The difference in performance is only marginal.
A.3 von Mises-Fisher Distribution
The von Mises-Fisher distribution (vMF) is the analogue of the Gaussian distribution on the surface of a hypersphere Hillen et al. (2017). The density function $f_{d}$ of the vMF distribution for a $D$-dimensional unit-length vector ${\bm{v}}$ is defined as
$f_{d}({\bm{v}}|\bm{{\alpha}},\kappa)=C_{d}(\kappa)\exp\left(\kappa\bm{{\alpha}}^{\top}{\bm{v}}\right),$ (17)
with mean vector $\bm{{\alpha}}$, scalar concentration parameter $\kappa$ and normalization constant $C_{d}(\kappa)$. The normalization constant $C_{d}(\kappa)$ is a complex function and difficult to compute for higher dimensions, which is why, for example, Tensorflow [^1] Tensorflow API Documentation - Accessed 2024-09-20 only supports the vMF distribution for $D\leq 5$. However, the vMF distribution is a viable similarity measure when using the unnormalized density function with $C_{d}(\kappa)=1$. Working with unnormalized densities highlights the relationship between the normal distribution and the vMF distribution.
Let $\hat{G}$ be the unnormalized PDF of a multivariate Gaussian with normalized mean $\bm{{\alpha}}$ and isotropic covariance $\bm{{\sigma}}^{2}=\kappa^{-1}{\bm{I}}$, then it is proportional to the vMF distribution for normalized vectors ${\bm{v}}$ with $|{\bm{v}}|=1$, as shown by
$\hat{G}({\bm{v}}|\bm{{\alpha}},\kappa)$ $=\exp\left(-\kappa\frac{({\bm{v}}-\bm{{\alpha}})^{\top}({\bm{v}}-\bm{{\alpha}})}{2}\right)$ (18)
$=\exp\left(-\kappa\frac{{\bm{v}}^{\top}{\bm{v}}+\bm{{\alpha}}^{\top}\bm{{\alpha}}-2{\bm{v}}^{\top}\bm{{\alpha}}}{2}\right)$ (19)
$=\exp\left(-\kappa\frac{1+1-2{\bm{v}}^{\top}\bm{{\alpha}}}{2}\right)$ (20)
$=\exp\left(-\kappa\frac{2-2{\bm{v}}^{\top}\bm{{\alpha}}}{2}\right)$ (21)
$=\exp\left(-\kappa\frac{1-{\bm{v}}^{\top}\bm{{\alpha}}}{1}\right)$ (22)
$=\exp\left(\kappa({\bm{v}}^{\top}\bm{{\alpha}}-1)\right)$ (23)
$\propto\exp\left(\kappa{\bm{v}}^{\top}\bm{{\alpha}}\right).$ (24)
Eq. 23 also shows the relationship to the HyperPg similarity with an untruncated Gaussian distribution and prototype mean activation $\mu=1$.
A.4 Fisher-Bingham Distribution
As the vMF distribution is the equivalent of an isotropic Gaussian distribution on the surface of a hypersphere, the Fisher-Bingham (FB) distribution is the equivalent of a Gaussian with full covariance matrix. Similar to the vMF, the normalization constant is difficult to compute for higher dimensions, but the unnormalized density function remains feasible.
For a $D$ dimensional space, the FB distribution is by a $D\times D$ matrix ${\bm{A}}$ of orthogonal vectors $(\bm{{\alpha}}_{1},\bm{{\alpha}}_{2},\ldots,\bm{{\alpha}}_{D})$, concentration parameter $\kappa$ and ellipticity factors $[\beta]_{2:D}$ where $\sum_{j=2}^{D}\beta_{j}=1$ and $0\leq 2|\beta_{j}|<\kappa$. The FB unnormalzied PDF is defined as
$b({\bm{v}}|{\bm{A}},\kappa,\beta)=\exp\left(\kappa\bm{{\alpha}}_{1}^{\top}{\bm{v}}+\sum_{j=2}^{D}\beta_{j}\left(\bm{{\alpha}}_{j}^{\top}{\bm{v}}\right)^{2}\right).$ (25)
Figure 9: Illustration of the Fisher-Bingham Distribution in $D=3$
The FB distribution’s main advantage is the elliptic form of the distribution on the surface of the hypersphere, offering higher adaptability than the other formulations (see Fig. 9. However, the parameter count and constraints are higher.
A.5 Mixture Models
HyperPg’s probabilistic nature lends itself to a mixture formulation. Let the definition of a HyperPg Mixture Prototype be ${\bm{p}}^{M}=(\bm{{\alpha}},\mu,\sigma,\pi)$ with additionally learned mixture weight $\pi$. Further, let’s define the probability of a latent vector ${\bm{z}}$ belonging to a Gaussian HyperPg prototype ${\bm{p}}$ as
$\phi({\bm{z}}|{\bm{p}})=s_{\mathrm{HyperPg}}({\bm{z}}|{\bm{p}}).$ (26)
Then the probability of ${\bm{z}}$ belonging to class $c$ can be expressed through the mixture of all prototypes ${\bm{p}}_{c}\in{\bm{P}}_{c}$ of class $c$, i.e.,
$\phi({\bm{z}}|c)=\sum_{{\bm{p}}_{c}\in{\bm{P}}_{c}}\pi({\bm{p}}_{c})\phi({\bm{z}}|{\bm{p}}_{c}).$ (27)
First experiments with mixture of HyperPg prototypes did not show any improvement over the standard formulation. However, this might change with other datasets.
Appendix B Ablations
B.1 Backbone Performance
HyperPgNets performance is not dependent on a CNN or Transformer based backbone. Fig. 10 shows the average test accuracy curves for HyperPgNet with concept aligned prototypes on CUB-200-2011. The maximum differnece in test accuracy between both architectures is 2%.
Figure 10: CUB-200-2011 Test Accuracies with different Backbones. The maximum difference in performance is 2%.
B.2 Number of prototypes
The number of prototypes per class or concept is an important hyper-parameter for prototype learning models like HyperPgNet. The choice of the number of prototypes controls, how many clusters the model should fit in latent space. For example, ProtoPNet requires 10 prototypes per class or 2000 prototypes in total for the CUB-200-2011 dataset (Chen et al., 2019). ProtoPool is able to merge similar prototypical parts, requiring only 202 prototypes for its best performance (Rymarczyk et al., 2022).
The change from class specific to concept aligned prototypes improves overall test performance, and the added information during training leads to fewer training epochs. Concept Bottleneck Models (CBM) are able to outperform the other models with only 30 learned concepts, or 1 prototype per concept. This indicates that the concept annotations on CUB-200-2011 do not produce multi-modal distributions in latent space, which would require more prototypes per concept. This behavior is confirmed in ablation studies on the number of prototypes for HyperPgNet. Fig. 11 presents the test accuracies for HyperPgNet on full CUB-200-2011 images after 60 training epochs with different numbers of prototypes per concept.
Figure 11: Ablation on number of prototypes in HyperPgNet on the CUB-200-2011 dataset. CUB-200-2011’s concept annotations have sufficient granularity for modeling with one prototype per concept. This is in line with CBM’s perfomance, which also learns only one prototype per concept.
Appendix C Extended Interpretability Analysis
C.1 Right for the Right Concept
Figure 12: Concept Prototypes from HyperPgNet
Fig. 12 shows more example prototypes for HyperPgNet. Without $L_{\mathrm{RRC}}$, the prototype do not always correspond to the defined prototype. For the “Orange Body” concept, the prototype also activates on confounding factors in the background. With $L_{\mathrm{RRC}}$ during training, the model learns to ignore these factors.
With the car dataset, the effect is even more pronounced. First analyses on the Stanford Cars experiments indicate, that most models focus on the front bumper below the headlights. This part of the car seems to be highly predictive for the class label. Because of this, the HyperPgNet, even with context annotations but without $L_{\mathrm{RRC}}$ will focus more on these areas of the image. For example the “bonnet” prototype does not focus on the car bonnet at all. Only when trained with $L_{\mathrm{RRC}}$, the prototype actually focuses on the human defined concept.
C.2 Latent Space Structure
Prototype learning is based on learning structures in the latent space. HyperPg specifically learns Gaussian distributions on the surface of a hypersphere in a high-dimensional latent space. Dimensionality reduction techniques like UMAP (McInnes et al., 2018) aim to preserve global and local structures from a high dimensional space when projecting into a low dimensional one. UMAP supports multiple distance metrices, enabling the projection onto the surface of a 3D sphere.
Fig. 13 illustrates UMAP projections of HyperPg concept-aligned prototypes trained on Stanford Cars. This visualization indicates that HyperPgNet is able to disentangle the different concepts in latent space, as the projection shows no overlap of the different clusters. When trained with $L_{\mathrm{RRC}}$, the prototypes are packed closer together. This could indicate, that the chosen hyper parameter of 20 prototypes per concept is higher than required as not all concept embeddings have the same diversity in the latent space.
Fig. 14 illustrates one UMAP projection of HyperPg concept-aligend prototypes trained on CUB-200-2011. In comparison to Stanford Cars, the latent space appears less structured. This could explain the higher difficulty associated with this dataset.
(a) HyperPgNet - $L_{\mathrm{RRC}}$
(b) HyperPgNet + $L_{\mathrm{RRC}}$
Figure 13: UMAP projection of HyperPg concept prototypes learned on Stanford Cars. Each dot represents one prototype, colored by the concept it represents.
Figure 14: UMAP Projection of HyperPg concept prototypes learned on CUB-200-2011.

### v2

@@
 However, these approaches are limited in their performance because each prototype must represent the entire image.
-Alternatively, parts-based approaches like ProtoPNet (Chen et al., 2019; Donnelly et al., 2021) learn a predefined number of prototypical parts per class.
-Each class-specific prototype is an image patch represented by a $1\times 1$ spatial unit in the latent space.
-The network classifies images by summarizing the similarities of all the latent image patches to all prototypes using a fully connected layer.
-PIPNet (Nauta et al., 2023) aligns learned prototypes with human perception by optimizing the model to assign the same prototype to two different views of the same augmented image patch.
-ProtoPShare (Rymarczyk et al., 2020) merges similar prototypes after training is complete, while ProtoPool (Rymarczyk et al., 2022) directly learns class-shared prototypes.
-Other approaches additionally replace the linear output layer with other interpretable models.
-ProtoKNN (Ukai et al., 2023) classifies the similarity scores using a k-nearest neighbor (KNN) and ProtoTree Nauta et al.
-(2021) learn a decision tree.
-HyperPgNet learns a predefined number of prototypes per concept, not per class.
-These concepts are provided as pixel-level annotations extracted by a new labeling pipeline based on the foundation models DINOv2 (Oquab et al., 2024) and SAM2 (Ravi et al., 2024).
-Similar to part-based prototypes, the concept prototypes span a $1\times 1$ spatial unit in the latent space.
-The prototypes are exclusive to each concept but shared among the classes.
-Prototype learning is also applied to image segmentation.
-ProtoSeg (Sacha et al., 2023) uses prototypical parts a latent space with high spatial resolution for image segmentation.
-Another method proposes non-learnable prototypes (Zhou et al., 2022), which are trained by clustering in the latent space instead of being optimized via Backpropagation.
-ProtoGMM (Moradinasab et al., 2024) builds on this by updating class-specific prototypes like a Gaussian Mixture Model (GMM) using an Expectation-Maximization (EM) algorithm (Liang et al., 2022) with fixed mixture weights.
-MGProto (Wang et al., 2023) also learns mixture weights for their prototype GMM model.
-HyperPgNet employs HyperPg, a novel prototype representation using Gaussian distributions on a hypersphere’s surface with learned parameters: direction $\bm{{\alpha}}$, mean $\mu$ and standard deviation (std) $\sigma$.
-Unlike ProtoGMM, which learns a multivariate Gaussian distribution in Euclidian space, each HyperPg prototype is defined by a directional vector $\bm{{\alpha}}$ and a $1D$-Gaussian distribution of cosine similarity values with mean $\mu$ and standard-deviation $\sigma$ around this vector.
-The use of cosine similarity for prototype learning has been shown to improve classification (Mettes et al., 2019) and has been widely adopted.
-(e.g., Donnelly et al., 2021; Rymarczyk et al., 2022; Sacha et al., 2023).
-HyperPgNet advances this by computing prototype activations as the probability density of each HyperPg prototype’s Gaussian distribution, allowing each prototype to adapt to the spread of the latent cluster.
+ProtoPNet (Chen et al., 2019) introduced the idea of prototypical parts.
+In this setting, each prototype is a latent patch of the input image, commonly a $1\times 1$ latent patch.
+The prototypes are each associated with a single class and learned via backpropagation without additional information.
+HyperPgNet builds on the idea of prototypical parts, but learns the prototypes in a transparent, concept-aligned manner.
+The prototypes are not class exclusive, but class shared.
+Each prototype corresponds to a human defined concept.
+To provide the required concept annotations, we propose a labeling pipeline based on foundation models such as DINOv2 (Oquab et al., 2024) and SAM2 (Ravi et al., 2024).
+HyperPgNet learns the concept-aligned prototypes by only adapting the loss functions and providing additional annotations, without requiring changes to the feature encoder like PIPNet (Nauta & Seifert, 2023) or Lucid-PPN (Pach et al., 2024), which uses a hybrid input head to disentangle prototype color and shape.
+Multiple successors build on the idea of ProtoPNet.
+ProtoPShare (Rymarczyk et al., 2020) and ProtoPool (Rymarczyk et al., 2022) use additional optimizations to learn class-shared prototypical parts.
+Deformable ProtoPNet (Donnelly et al., 2021) learns a mixture of prototypical parts with dynamic spatial arrangement.
+Other work replace the linear output layer of ProtoPNet with other interpretable models: ProtoKNN (Ukai et al., 2023) uses a k-nearest neighbor classifier and ProtoTree (Nauta et al., 2021) employs a decision tree.
+All these models change how the prototype activations are further processed.
+However, they all rely on the point-based prototype formulation introduced by ProtoPNet based on the $L_{2}$ or cosine similarity.
+HyperPgNet proposes a novel prototype formulation based on Gaussian distributions on the hypersphere, without changing the downstream processing.
+This probabilistic view with learned mean and variance allows HyperPgNet to learn prototypes with different degrees of specialization.
+Furthermore, the statistical confidence derived from the learned variance could be integrate in further downstream processing.
+The view of prototype learning as clustering in latent space was introduced for image segmentation (Zhou et al., 2022).
+The image patches are points in latent space with the prototypes as cluster centers.
+ProtoGMM (Moradinasab et al., 2024) and MGProto (Wang et al., 2023) build on this idea to model the prototypes as Gaussian mixture models for image segmentation.
+HyperPg adapts the idea of Gaussian prototypes used in image segmentation to hyperspherical latent spaces based on the cosine similarity.
+The cosine similarity has been shown to perform well in classification tasks (Mettes et al., 2019) and also been applied in recent prototype learning works (Zhou et al., 2022; Ukai et al., 2023).
+This makes HyperPgNet the first model to employ probabilistic prototypes with learnable mean and variance in a hyperspherical space.
 Concept-Based Methods.
@@
 Figure 1: Illustration on how the different prototype formulations compute the similarity between a prototype ${\bm{p}}$ and latent vector ${\bm{z}}$.
-$L_{2}$ prototypes compute the Euclidian distance between two points in latent space.
+$L_{2}$ prototypes compute the Euclidean distance between two points in latent space.
 Hyperspherical prototypes use the cosine similarity of normalized vectors, which corresponds to the angle between two points on a hypersphere.
-Gaussian prototypes model a Gaussian distribution in Euclidian space and compute a probability density function (PDF).
+Gaussian prototypes model a Gaussian distribution in Euclidean space and compute the probability density function (PDF).
 HyperPg prototypes learn a Gaussian distribution of cosine similarities, thereby projecting a Gaussian distribution onto the surface of a hypersphere.
@@
 Let $\mathcal{G}(x,\mu,\sigma)$ be the cumulative Gaussian distribution function.
-Then HyperPg’s similarity measure based on the truncated Gaussian distribution is defined as
+Then, HyperPg’s similarity measure based on the truncated Gaussian distribution is defined as
 $s_{\mathrm{HyperPg}}({\bm{z}}|{\bm{p}}^{H})$ $=\mathcal{T}_{G}(s_{\cos}({\bm{z}}|\bm{{\alpha}});\mu,\sigma,-1,1)$ (4)
 $=\frac{{\mathcal{N}}\left(s_{\cos}({\bm{z}}|\bm{{\alpha}});\mu,\sigma\right)}{\mathcal{G}(1,\mu,\sigma)-\mathcal{G}(-1,\mu,\sigma).}$ (5)
-(a) $\mu=1$
-(b) $\mu=0.6$
-(c) $\mu=0$
-(d) $\mu=-1$
 Figure 2: HyperPg learns a Gaussian distribution with mean $\mu$ and std $\sigma$ of cosine similarities with the anchor $\bm{{\alpha}}$.
-The plots visualize HyperPg’s activations on the $3D$ hypersphere with anchor $\bm{{\alpha}}=(0,0,1)$, std $\sigma=0.1$ and and various mean cosine similarities $\mu\in[-1,1]$.
+The plots visualize HyperPg’s activations on the $3D$ hypersphere with anchor $\bm{{\alpha}}=(0,0,1)$, std $\sigma=0.1$ and and various mean cosine similarities $\mu\in[-1,1]$ for 10k random samples.
 The anchor $\bm{{\alpha}}$ is shown as a red arrow, and std $\sigma$ governs the width of the distribution.
@@
 HyperPg achieves the same effect by learning only one prototype vector (the anchor) and just two additional scalar parameters.
-This significantly increases HyperPg’s representational power compared to standard hyperspherical and Gaussian prototypes, as a single HyperPg prototype can cover multiple clusters simultaneously.
+This significantly increases HyperPg’s representational power compared to standard hyperspherical and Gaussian prototypes and is a major difference to the von Mises-Fisher distribution (see also Appendix A).
 Combining the strengths of hyperspherical and Gaussian prototypes, HyperPg can learn more complex structures in the latent space, such as human-defined concepts, and provide a more meaningful similarity measure.
@@
 The Concept Activation Regions (CAR, Crabbé & van der Schaar, 2022) model proposes a concept density loss defined as
-$L_{\mathrm{Den}}$ $=-\frac{1}{|K|}\sum_{k\in K}\frac{1}{|{\bm{P}}_{k}|}\sum_{{\bm{p}}_{k}\in{\bm{P}}_{k}}\phi({\bm{p}}_{k},{\bm{Z}}_{k})-\phi({\bm{p}}_{k},{\bm{Z}}_{\neg k})$ (8)
+$L_{Density}$ $=-\frac{1}{|K|}\sum_{k\in K}\frac{1}{|{\bm{P}}_{k}|}\sum_{{\bm{p}}_{k}\in{\bm{P}}_{k}}\phi({\bm{p}}_{k},{\bm{Z}}_{k})-\phi({\bm{p}}_{k},{\bm{Z}}_{\neg k})$ (8)
 $\phi({\bm{p}},{\bm{Z}})$ $=\frac{1}{|{\bm{Z}}|}\sum_{{\bm{z}}\in{\bm{Z}}}s({\bm{p}},{\bm{z}}).$ (9)
@@
 In HyperPgNet the RRC-Loss further strengthens the prototype-concept association.
+4.3 Multi-Objective Loss Function
+To train a prototype learning network like HyperPgNet for downstream tasks like image classification, a multi-objective loss function is employed. This multi-objective loss function is defined as
+$L=L_{\mathrm{CE}}+\lambda_{\mathrm{Clst}}L_{\mathrm{Clst}}+\lambda_{\mathrm{Sep}}L_{\mathrm{Sep}}+\lambda_{\mathrm{RRC}}L_{\mathrm{RRC}},$
+where $L_{\mathrm{CE}}$ is the cross-entropy loss over network predictions and ground truth image labels.
+ProtoPNet uses $\lambda_{\mathrm{Clst}}=0.8$ and $\lambda_{\mathrm{Sep}}=0.08$ (Chen et al., 2019).
+For HyperPgNet, the different loss terms are weighted equally, meaning all $\lambda=1$.
 4.4 Network Architecture
@@
 3 illustrates HyperPgNet’s Architecture for interpretable image classification.
-HyperPgNet uses a pretrained, off-the-shelf feature encoder as a backbone, such as the convolution pyramid from a CNN like ConvNeXt-tiny (Liu et al., 2022) or a transformer architecture like Segformer (Xie et al., 2021).
-The high-dimensional feature maps provided by the backbone are passed through a neck, consisting of two $1\times 1$ convolution layers with ReLU activation in between, projecting into a smaller latent space with $256$ channels, followed by Layer Normalization.
-The hyperparameter of 256 channels aligns with previous work (e.g., Rymarczyk et al., 2022).
-HyperPg’s similarity function computes the truncated Gaussian PDF over the cosine similarity of the latent feature maps with the HyperPg prototype anchors $\bm{{\alpha}}$.
-HyperPgNet implements the prototype learning part of the network as a module with two layers: First, a prototype learning layer computes the cosine similarity of learnable HyperPg prototype anchors $\alpha$ to the latent feature map $z$, similar to other prototype learning architectures.
-Second, the new Density Estimation layer with learnable parameters mean $\mu$ and std $\sigma$ computes the truncated Gaussian PDF over the activations of the previous layer.
-This implementation of the HyperPg Module with two internal layers supports swapping out the prototype similarity measure to other measures in the future.
-Additionally, the Density Estimation Layer concept could be applied to other architectures to learn a Gaussian distribution over the previous layer’s activations.
+HyperPgNet uses a pretrained feature encoder such as ConvNext-tiny (Liu et al., 2022) or MiT-B4 (Xie et al., 2021) as a backbone model.
+A neck consisting of two $1\times 1$ convolution layers with ReLU activation in between projects the high dimensional feature map of the backbone into a lower dimensional feature space.
+The HyperPg prototype module is implemented with two layers: First, a prototype learning layer computes the cosine similarity of the learnable HyperPg anchors $\bm{{\alpha}}$ to the latent vectors produced by the neck.
+Second, a Density Estimation layer with learnable parameters mean $\mu$ and std $\sigma$ computes the Gaussian PDF over the activations of the previous layer.
+Future extensions can adapt both components independently, for example by implementing a hyperbolic similarity measure or a multi-modal probability distribution.
 Finally, the output of the HyperPg Module is passed through a single fully connected layer (FCL) to produce the output logits or class scores.
-HyperPgNet is trained for image classification with a multi-objective loss function. This loss function is defined as
-$L=L_{\mathrm{CE}}+\lambda_{\mathrm{Den}}L_{\mathrm{Den}}+\lambda_{\mathrm{RRC}}L_{\mathrm{RRC}},$
-where $L_{\mathrm{CE}}$ is the Crossentropy loss over network predictions and ground truth image labels.
-By choosing different values for $\lambda$, the model’s tradeoff between task performance and interpretability can be adjusted.
 5 Concept Extraction at Scale
@@
 Foundation models are leveraged to generate these annotations, reducing the need for human labeling.
-The CUB-200-2011 dataset (Wah et al., 2011) contains 11,788 images of birds across 200 sub-species.
-Each image is annotated with up to 15 part locations, such as head, tail and back.
-The dataset includes 312 binary attributes describing patterns and colors of these parts, or additional shape information like “duck-like” or “owl-like”.
-For concept learning, 30 attributes are manually selected, excluding small concepts like eye-color or class exclusive attributes like “duck-like”.
-To extract pixel level annotation masks, the point annotations for the attributes are given to a Segment Anything 2 (SAM2) model(Ravi et al., 2024).
-Using the point locations belonging to other attributes as negative points, segmentation masks are generated (see Fig.
-4).
-The segmentation mask is then assigned the attribute from the dataset annotations to define a concept.
+The CUB-200-2011 (Wah et al., 2011) annotates each image with up to 15 part locations (e.g., head, tail, wing).
+Each part is labeled with some attribute describing the coloring, pattern or shape.
+A Segment Anything 2 (SAM2) model (Ravi et al., 2024) creates pixel-level segmentation masks based on the part locations.
+The Stanford Cars dataset (Krause et al., 2013) does not provide part level annotations.
+We defined a list of 10 car parts (e.g., wheel, headlight, radiator) and use a Grounding DINO model (Liu et al., 2023) to generate bounding boxes for each part.
+Finally, SAM2 creates pixel-level segmentation masks based on the bounding boxes.
+This automated pipeline labeled the 16k images of Stanford Cars within 2h on a NVIDIA 4060Ti with 16GB VRAM, and can be easily adapted to other datasets.
+Fig.
+4 presents some examples for CUB-200-2011 and Stanford Cars.
 (a) Concept Segmentation Masks for CUB-200-2011 based on provided positive and negative part annotations.
@@
 Figure 4: Examples for Automated Concept Extraction.
-The Stanford Cars dataset (Krause et al., 2013) does not provide part annotations.
-This dataset contains 16,185 images of 196 car classes, categorized by make, model and year (e.g., BMW M3 coupe 2012).
-The dataset provides only bounding boxes for the main subject matter, useful for foreground-background separation.
-To obtain pixel-level annotations, a list of 10 car parts (e.g., “wheel”, “headlight”, “radiator”) is defined.
-A Grounding DINO model (Liu et al., 2023) uses this list to generate bounding boxes for parts in the image.
-A post-processing step removes possible outliers by eliminating bounding boxes outside the original bounding box provided by the dataset.Finally, SAM2 generates pixel-level segmentation masks (see Fig.
-4).
-For Stanford Cars, only the part label is used as the concept label as the dataset does not provide additional part attributes.
-The automated pipeline labeled the entire Stanford Cars dataset within 2h on an NVidia 4060ti with 16GB VRAM and can be easily adapted to other datasets.
 6 Experiments
-The experiments were performed on two datasets: CUB-200-2011 bird species classification dataset (Wah et al., 2011) and the Stanford Cars dataset (Krause et al., 2013).
-Two variants of HyperPgNet were tested, with and without the $L_{\mathrm{RRC}}$ loss.
-HyperPgNet was compared to ProtoPNet with both standard $L_{2}$ and the newly introduced HyperPg prototypes, and Concept Bottleneck Models (CBM) with joint training scheme (Koh et al., 2020).
-Segformer and ConvNext served as Transformer and CNN baselines, respectively.
-All models were implemented with pretrained Segformer backbones in Pytorch.
-However, HyperPgNet’s performance is independent of the chosen backbone architecture (see Appendix B.1).
-In contrast to prior work, the training and testing was done on full images, without crops to the bounding boxes provided by the datasets.
-For data augmentation RandomPerspective, RandomHorizontalFlip, and RandomAffine were used in an online manner.
-Thus, for CUB-200-2011, each class has 30 training images, instead of the offline augmentation to 1200 images used in prior work (e.g., Chen et al., 2019; Rymarczyk et al., 2020; Ukai et al., 2023).
-All models were trained with batch size 96 until convergence.
-The experiments were performed on a workstation with a single NVidia RTX 3090 GPU (24GB VRAM) per model.
+The experiments were performed on two datasets: CUB-200-2011 (CUB) with 200 bird species Wah et al.
+(2011) and Stanford Cars (CARS) with 196 car models Krause et al.
+(2013).
+Implementation details such as data preprocessing or model implementation are provided in Appendix B.
+Birds Cars
+# Prototypes MiT-B4 ConvNeXt # Prototypes MiT-B4 ConvNeXt
+BB Baseline - 17.7 74.2 - 1.9 57.5
+CBM - 75.7 77.9 - 79.9 81.0
+PP ProtoPNet (PPN) 2000 68.0 68.1 1960 86.4 87.0
+PPN + HyperPG 2000 70.5 65.0 1960 87.4 77.3
+CAP HyperPgNet $-L_{\mathrm{RRC}}$ 300 76.5 76.9 180 88.6 88.9
+(Ours) HyperPgNet $+L_{\mathrm{RRC}}$ 300 74.1 71.5 180 81.2 86.0
+Table 1: Test Top-1 Accuracy. Models are grouped by their reasoning process. BB: Black Box. PP: Prototypical Parts. CAP: Concept Aligned Prototypes.
 6.1 Quantitative Results
-Table 1: Test Top-1 Accuracy with Segformer Backbone. Models are grouped by their reasoning process. BB: Black Box. PP: Prototypical Parts. CAP: Concept Aligned Prototypes.
-CUB-200-2011 Stanford Cars
-# Prototypes Accuracy [%] # Prototypes Accuracy [%]
-BB Segformer Baseline - 17.7 - 1.9
-ConvNeXt Baseline - 74.2 - 57.5
-CBM - 75.7 - 79.9
-PP ProtoPNet 2000 68.0 1960 86.4
-ProtoPNet + HyperPg 2000 70.5 1960 87.4
-CAP HyperPgNet - $L_{\mathrm{RRC}}$ 300 76.5 180 88.6
-HyperPgNet + $L_{\mathrm{RRC}}$ 300 74.1 180 81.2
-Figure 5: Test Accuracy with Segformer Backbone.
-Table 1 shows the top-1 accuracy for models on the CUB-200-2011 (Birds) and Stanford Cars (Cars) datasets using the Segformer backbone.
-The Test Accuracy per Epoch curves are visualized in Fig.
-5.
-The Segformer baseline overfit to the small training datasets (17.7% and 1.9% test accuracy), while the ConvNeXt baseline performed better (74.2% and 57.5%).
-CBM achieved high results quickly, scoring 75.7% on Birds and 79.9% on Cars.
-ProtoPNet required the most training time, converging only after about 490 epochs, scoring 68% on Birds and 86.4% on Cars.
-Switching ProtoPNet’s prototypes from $L_{2}$ to HyperPg improved both training speed and accuracy, converging in 200 epochs with 70.5% and 87.4% test accuracy.
-HyperPgNet further improved on this by learning human-defined concepts.
-HyperPgNet with concept aligned prototypes, but without $L_{\mathrm{RRC}}$, outperformed the other models.
-It achieved 76.5% accuracy on Birds and 88.6% on Cars after only 40 training epochs.
-With the “Right for the Right Concept” loss, the model retained the high learning speed and performance, although the test accuracy dropped slightly to 74.1% and 81.2%.
-This performance loss is offset by increased transparency due to more precise concept prototypes.
-For additional ablation studies, see Appendix B.
-6.2 Qualitative Results
-HyperPg prototypes do not only enhance the models learning speed and task performance, but also the interpretabilty.
-Prior work built activations by overlaying the up-sampled latent space activation over the image.
-Inspired by GradCAM, Fig.
-6 shows the gradient of the prototype activations backpropagated to the input image.
-Class-based prototypes focus on regions relevant for the prediction, but do not follow human defined concepts.
-With concept-alignment, the prototype focuses more towards the white head of the bird.
-But as the concept alignment is only encouraged by $L_{\mathrm{Clst}}$ and $L_{\mathrm{Sep}}$ the prototype still activates on other bright areas on the birds body.
-With the addition of $L_{\mathrm{RRC}}$ to the training regime, the prototypes are forced to focus mainly on the annotated region.
-Thus, the head prototype ignores the wing area, which is learned by a different prototype.
-The overall learned concept prototypes fit more precisely to the images’ annotated areas.
-For additional examples and latent space analyses, please refer to Appendix C.
-(a) ProtoPNet + HyperPg
-(b) HyperPgNet - $L_{\mathrm{RRC}}$
-(c) HyperPgNet + $L_{\mathrm{RRC}}$
-Figure 6: Prototype Gradient Maps for different HyperPgNet models.
+Figure 5: CUB Test Accuracy per Epoch.
+Table 1 reports the top-1 accuracy for the tested models on the CUB and CARS datasets with different backbones.
+HyperPgNet achieves comparable results with the black box CBM and baseline models.
+Notably, HyperPgNet outperforms CBM on CARS, as it can differentiate between different modes for the same concept by learning multiple prototypes per concept.
+CBM is restricted to a binary concept presence prediction.
+Including ${\mathcal{L}}_{\mathrm{RRC}}$ slightly lowers HyperPgNet’s performance but is still comparable to the black-box models.
+Fig.
+5 shows the test accuracy over epochs for CUB with ConvNeXt backbone.
+The concept aligned models CBM and HyperPgNet are the fastest to converge at around 100 epochs.
+ProtoPNet requires over 400 training epochs until the best performance level is reached.
+However, switching from the standard $L_{2}$ based prototypes to HyperPg prototypes effectively halves the required training epochs.
+6.2 Qualitative Comparison
+Fig.
+7 shows the pixel attribution of the highest activated prototype on each image for the tested prototype learning models with ConvNeXt backbone.
+In contrast to prior work (e.g., Chen et al., 2019; Ukai et al., 2023), this visualization is based on the prototype gradients, similar to saliency maps such as GradCAM (Selvaraju et al., 2017), which should avoid spatial misalignment often observed in prototype explanations (Sacha et al., 2024).
+The concept aligned prototypes learned by HyperPgNet focus more on the relevant parts of the image subject than the class-based prototypes of ProtoPNet.
+The network seems to focus correctly on either the overall color of the bird and ignoring the background (first three examples) or on specific markings (last three examples).
+The addition of the ${\mathcal{L}}_{\mathrm{RRC}}$ loss seems to further increase the main focus of the prototype activation on the relevant image regions.
+For ProtoPNet, no qualitative difference can be detected between the standard $L_{2}$ or HyperPg prototypes.
+With both prototype formulations, ProtoPNet is likely to also activate on parts of the image background.
+This is a fundamental weakness of training ProtoPNet in an opaque manner without alignment to relevant concepts, which HyperPgNet addresses.
+On the CARS dataset, the effect of ${\mathcal{L}}_{\mathrm{RRC}}$ is more apparent.
+The defined concepts for this dataset are limited to car parts without fine-grained differentiation between patterns or shapes, such as the concepts on CUB.
+For this dataset, one of the most predictive image regions across models seems to be the front bumper of the cars, which has not been annotated with its own concept.
+As shown in Fig.
+6, the concept-aligned prototypes still activate in this region.
+However, by training the model with the ${\mathcal{L}}_{\mathrm{RRC}}$ loss, the prototypes focus more on the annotated image regions.
+Figure 6: Concept prototype activation after training with and without ${\mathcal{L}}_{\mathrm{RRC}}$ with MiT-B4 backbone.
+Figure 7: Highest prototype activation per image with ConvNeXt backbone.
+6.3 Ablation
+# Prototypes CUB Accuracy
+CBM 71.6
+1 70.5
+5 68.9
+10 68.0
+20 67.3
+Table 2: Test Top-1 Accuracy on CUB with ResNet50 backbone with varying numbers of prototypes per concept.
+We conduct an ablation on the number of prototypes per concept.
+Table 2 presents the test accuracies on CUB for HyperPgNet with ResNet50 backbone.
+The difference in task performance over the different model configurations is negligible.
+However, looking at some qualitative examples as depicted in Fig.
+8 a difference in prototype quality can be observed.
+With fewer prototypes per concept, the model is more likely to also react to parts of the mage background in order to still achieve a low training error.
+With too many prototypes per concept, the latent space becomes too segmented.
+The prototypes start to focus on smaller subparts of the image subject, potentially decreasing the interpretability.
+The chosen configuration of 10 prototypes per concept seems to be a sweet spot regarding predictive performance and prototype quality for the experiments.
+Figure 8: Qualitative Results for HyperPgNet with ResNet50 backbone and varying numbers of prototypes per concept.
 7 Conclusion
 This work introduces HyperPg, a new prototype representation learning a probability distribution on the surface of a hypersphere in latent space.
-HyperPg adapt to the variance of clusters in latent space and improve training time and accuracy compared to other prototype formulations.
-HyperPgNet leverages HyperPg to learn human defined concept prototypes, instead of black-box optimized prototypes.
-The combination of probabilistic prototypes on the hypersphere and concept aligned prototypes allows HyperPgNet to outperform other prototype learning approaches.
-One limitation HyperPgNet faces are increased computational requirements due to the inclusion of concept annotations and $L_{\mathrm{RRC}}$ during training.
-However, this is offset by an acceleration of the training process regarding epochs and number of samples.
-Coupled with the increased transparency and interpretability of the model, this makes HyperPgNet a strong contender for scenarios with few training samples and high safety requirements, like medical applications or human-robot interaction.
-Acknowledgments
-This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 448648559
+HyperPg prototypes adapt to the variance of clusters in latent space and improve training time and accuracy compared to other prototype formulations.
+HyperPgNet leverages HyperPg to learn human-defined concept prototypes instead of black-box optimized prototypes.
+The combination of probabilistic prototypes on the hypersphere and concept-aligned prototypes allows HyperPgNet to outperform other prototype learning approaches with regards to accuracy and interpretability.
+One limitation HyperPgNet faces are slightly higher computational requirements due to the inclusion of concept annotations and $L_{\mathrm{RRC}}$ during training.
+However, this is offset by faster convergence.
+Coupled with the increased transparency and interpretability of the model, this makes HyperPgNet a strong contender for scenarios with higher requirements for model trust and safety, like medical applications or human-robot interaction.
+Reproducibility Statement
+For reproducing the experiments, please refer to the following sections: Sec.
+4 describes the required loss functions and the overall network architecture.
+Subsection 3.3 details the activation function of the newly introduced HyperPg prototypes.
+Appendix B provides detailed information on the experiment implementations, including training hyperparameters, processing of the datasets and model implementation details.
+Furthermore, we will publish the code on GitHub after the double-blind review period.
 Appendix A Adapting HyperPg to other Probability Distributions
@@
 with median $x_{0}$ and average absolute deviation $\gamma$. The HyperPg prototypes with Cauchy are defined as accordingly as ${\bm{p}}^{\mathrm{Cauchy}}=(\bm{{\alpha}},x_{0},\gamma)$.
-Figure 7: PDF for the Gaussian and Cauchy distribution of cosine similarity values. The Cauchy distribution has heavier tails, avoiding vanishing gradients issues.
 Fig.
 9 illustrates the PDF of the Gaussian and Cauchy distributions with $\mu=x_{0}=1$ and $\sigma=\gamma=0.2$, i.e., the main probability mass is aligned with the anchor $\bm{{\alpha}}$.
-The Gaussian distributions PDF quicly approaches zero and stays near constant.
+The Gaussian distributions PDF quickly approaches zero and stays near constant.
 This could potentially cause vanishing gradient issues during training.
@@
 However, experiments on CUB-200-2011 showed no significant performance difference between using HyperPg with the Gaussian or Cauchy distribution.
+Figure 9: PDF for the Gaussian and Cauchy distribution of cosine similarity values. The Cauchy distribution has heavier tails, avoiding vanishing gradients issues.
 A.2 Truncated Distributions
@@
 $=\exp\left(\kappa({\bm{v}}^{\top}\bm{{\alpha}}-1)\right)$ (23)
-$\propto\exp\left(\kappa{\bm{v}}^{\top}\bm{{\alpha}}\right).$ (24)
-Eq. 23 also shows the relationship to the HyperPg similarity with an untruncated Gaussian distribution and prototype mean activation $\mu=1$.
+$=\exp(\kappa{\bm{v}}^{\top}\bm{{\alpha}}-\kappa)$ (24)
+$=\exp(\kappa)^{-1}\exp(\kappa{\bm{v}}^{\top}\bm{{\alpha}})$ (25)
+$\sim\exp\left(\kappa{\bm{v}}^{\top}\bm{{\alpha}}\right).$ (26)
+Eq.
+23 also shows the relationship to the HyperPg similarity with an untruncated Gaussian distribution and prototype mean activation $\mu=1$.
+Fig.
+11 presents a simulation of the vMF distribution on a 3D sphere.
+While both the vMF distribution and HyperPg activation can produce a spherical, gaussian-like activation pattern on the surface of a hypersphere, the vMF distribution cannot produce the ring pattern shown in Fig.
+2.
+The ring pattern produced by adapting HyperPg’s mean similarity $\mu$ could be approximated by a mixture of vMF distributions.
+Figure 11: Changing the concentration parameter $\kappa$ is akin to changing HyperPg’s std $\sigma$.
 A.4 Fisher-Bingham Distribution
@@
 $b({\bm{v}}|{\bm{A}},\kappa,\beta)=\exp\left(\kappa\bm{{\alpha}}_{1}^{\top}{\bm{v}}+\sum_{j=2}^{D}\beta_{j}\left(\bm{{\alpha}}_{j}^{\top}{\bm{v}}\right)^{2}\right).$ (27)
+The FB distribution’s main advantage is the elliptic form of the distribution on the surface of the hypersphere, offering higher adaptability than the other formulations (see Fig.
+12.
+However, the parameter count and constraints are higher.
 Figure 12: Illustration of the Fisher-Bingham Distribution in $D=3$
-The FB distribution’s main advantage is the elliptic form of the distribution on the surface of the hypersphere, offering higher adaptability than the other formulations (see Fig.
-9.
-However, the parameter count and constraints are higher.
 A.5 Mixture Models
@@
 First experiments with mixture of HyperPg prototypes did not show any improvement over the standard formulation. However, this might change with other datasets.
-Appendix B Ablations
-B.1 Backbone Performance
-HyperPgNets performance is not dependent on a CNN or Transformer based backbone.
-Fig.
-10 shows the average test accuracy curves for HyperPgNet with concept aligned prototypes on CUB-200-2011.
-The maximum differnece in test accuracy between both architectures is 2%.
-Figure 10: CUB-200-2011 Test Accuracies with different Backbones. The maximum difference in performance is 2%.
-B.2 Number of prototypes
-The number of prototypes per class or concept is an important hyper-parameter for prototype learning models like HyperPgNet.
-The choice of the number of prototypes controls, how many clusters the model should fit in latent space.
-For example, ProtoPNet requires 10 prototypes per class or 2000 prototypes in total for the CUB-200-2011 dataset (Chen et al., 2019).
-ProtoPool is able to merge similar prototypical parts, requiring only 202 prototypes for its best performance (Rymarczyk et al., 2022).
-The change from class specific to concept aligned prototypes improves overall test performance, and the added information during training leads to fewer training epochs.
-Concept Bottleneck Models (CBM) are able to outperform the other models with only 30 learned concepts, or 1 prototype per concept.
-This indicates that the concept annotations on CUB-200-2011 do not produce multi-modal distributions in latent space, which would require more prototypes per concept.
-This behavior is confirmed in ablation studies on the number of prototypes for HyperPgNet.
-Fig.
-11 presents the test accuracies for HyperPgNet on full CUB-200-2011 images after 60 training epochs with different numbers of prototypes per concept.
-Figure 11: Ablation on number of prototypes in HyperPgNet on the CUB-200-2011 dataset.
-CUB-200-2011’s concept annotations have sufficient granularity for modeling with one prototype per concept.
-This is in line with CBM’s perfomance, which also learns only one prototype per concept.
+Appendix B Implementation Details
+B.1 Data Preprocessing
+In contrast to prior work (e.g.
+Chen et al., 2019; Rymarczyk et al., 2020; Ukai et al., 2023) the experiment used an online augmentation process, resulting in 30 training images per class and epoch.
+The input images were first resized to a resolution of $224\times 224$ without cropping to bounding box annotations.
+The augmentations consisted of RandomPerspective, RandomHorizontalFlip and RandomAffine.
+B.2 Hyperparameters
+The prototypical part networks ProtoPNet and HyperPgNet use a convolutional neck after the feature encoder and work on a latent feature map of size $7\times 7\times 256$.
+The models were trained with a minibatch size of 96 images using AdamW optimizer with learning rate 1e-4 and weight decay 1e-4.
+B.3 Compute Resources
+All experiments were performed on a workstation with a single NVIDIA RTX 3090 GPU (24 GB VRAM) per model.
+B.4 Model Implementation
+The models were implement with different feature encoding backbones.
+ConvNeXt-tiny (Liu et al., 2022) and ResNet50 (He et al., 2016) with pretrained weights on ImageNet provide CNN based backbones.
+As a transformer based backbone MiT-B4 (Xie et al., 2021) as provided by HuggingFace is used.
+MiT-B4 is pretrained on ImageNet and finetuned on ADE20k, Cityscapes and COCO-stuff.
+The models are implemented as follows:
+  • The black box baseline model uses the pretrained feature encoder as backbone followed by the classification head: a single linear output layer and softmax activation.
+  • Concept Bottleneck Models (Koh et al., 2020) integrate the concept bottleneck, a linear layer with one neuron per concept, between the backbone and classification head.
+  • ProtoPNet (Chen et al., 2019) use a $L_{2}$ based prototype layer between backbone and classification head.
+  • ProtoPNet + HyperPg use a HyperPg based prototype module between backbone and classification head.
+  • HyperPgNet $-{\mathcal{L}}_{\mathrm{RRC}}$ uses a HyperPg prototype module between backbone and classification head.
+The prototypes are trained with concept alignment, but without the ${\mathcal{L}}_{\mathrm{RRC}}$ loss.
+  • HyperPgNet $+{\mathcal{L}}_{\mathrm{RRC}}$ uses a HyperPg prototype module between backbone and classification head.
+The prototypes are trained with concept alignment and including the ${\mathcal{L}}_{\mathrm{RRC}}$ loss.
 Appendix C Extended Interpretability Analysis
-C.1 Right for the Right Concept
-Figure 12: Concept Prototypes from HyperPgNet
-Fig.
-12 shows more example prototypes for HyperPgNet.
-Without $L_{\mathrm{RRC}}$, the prototype do not always correspond to the defined prototype.
-For the “Orange Body” concept, the prototype also activates on confounding factors in the background.
-With $L_{\mathrm{RRC}}$ during training, the model learns to ignore these factors.
-With the car dataset, the effect is even more pronounced.
-First analyses on the Stanford Cars experiments indicate, that most models focus on the front bumper below the headlights.
-This part of the car seems to be highly predictive for the class label.
-Because of this, the HyperPgNet, even with context annotations but without $L_{\mathrm{RRC}}$ will focus more on these areas of the image.
-For example the “bonnet” prototype does not focus on the car bonnet at all.
-Only when trained with $L_{\mathrm{RRC}}$, the prototype actually focuses on the human defined concept.
-C.2 Latent Space Structure
+C.1 Latent Space Structure
 Prototype learning is based on learning structures in the latent space.
@@
 Dimensionality reduction techniques like UMAP (McInnes et al., 2018) aim to preserve global and local structures from a high dimensional space when projecting into a low dimensional one.
-UMAP supports multiple distance metrices, enabling the projection onto the surface of a 3D sphere.
+UMAP supports multiple distance metrics, including hyperspheric manifold distances, thus retaining some of the high dimensional structure when projecting onto a 3D sphere.
 Fig.
@@
 Fig.
-14 illustrates one UMAP projection of HyperPg concept-aligend prototypes trained on CUB-200-2011.
+14 illustrates one UMAP projection of HyperPg concept-aligned prototypes trained on CUB-200-2011.
 In comparison to Stanford Cars, the latent space appears less structured.
@@
 (b) HyperPgNet + $L_{\mathrm{RRC}}$
-Figure 13: UMAP projection of HyperPg concept prototypes learned on Stanford Cars. Each dot represents one prototype, colored by the concept it represents.
+Figure 13: UMAP projection of HyperPg concept prototypes learned on Stanford Cars. Each dot represents one prototype with the color indicating a concept.
 Figure 14: UMAP Projection of HyperPg concept prototypes learned on CUB-200-2011.

### v3

@@
-HyperPg - Prototypical Gaussians on the Hypersphere for Interpretable Deep Learning
+An Overview of Prototype Formulations for Interpretable Deep Learning
 Abstract
-Prototype Learning methods provide an interpretable alternative to black-box deep learning models.
-Approaches such as ProtoPNet learn, which part of a test image “look like” known prototypical parts from training images, combining predictive power with the inherent interpretability of case-based reasoning.
-However, existing approaches have two main drawbacks: A) They rely solely on deterministic similarity scores without statistical confidence.
-B) The prototypes are learned in a black-box manner without human input.
-This work introduces HyperPg, a new prototype representation leveraging Gaussian distributions on a hypersphere in latent space, with learnable mean and variance.
-HyperPg prototypes adapt to the spread of clusters in the latent space and output likelihood scores.
-The new architecture, HyperPgNet, leverages HyperPg to learn prototypes aligned with human concepts from pixel-level annotations.
-Consequently, each prototype represents a specific concept such as color, image texture, or part of the image subject.
-A concept extraction pipeline built on foundation models provides pixel-level annotations, significantly reducing human labeling effort.
-Experiments on CUB-200-2011 and Stanford Cars datasets demonstrate that HyperPgNet outperforms other prototype learning architectures while using fewer parameters and training steps.
-Additionally, the concept-aligned HyperPg prototypes are learned transparently, enhancing model interpretability.
+Prototypical part networks offer interpretable alternatives to black-box deep learning models.
+However, many of these networks rely on Euclidean prototypes, which may limit their flexibility.
+This work provides a comprehensive overview of various prototype formulations.
+Experiments conducted on the CUB-200-2011, Stanford Cars, and Oxford Flowers datasets demonstrate the effectiveness and versatility of these different formulations.
+XAI, Prototype Learning, Image Classification, Deep Learning
 1 Introduction
@@
 However, the decision-making processes of these models lack transparency and interpretability, making deployment in safety-critical areas challenging.
-Explainable Artificial Intelligence (XAI) seeks to develop interpretability methods to open the black-box reasoning processes of these models and increase trust in their decisions.
-XAI methods can be broadly divided into two categories: First, Post-Hoc Methods like LIME (Ribeiro et al., 2016), SHAP (Lundberg & Lee, 2017) or GradCAM Selvaraju et al.
-(2017) offer explanations for predictions without requiring retraining.
-While applicable in many scenarios, post-hoc methods may not actually align with the models’ decision making processes, potentially leading to interpretations that are not entirely faithful (Rudin, 2019).
+Explainable Artificial Intelligence (XAI) seeks to develop interpretability methods to open the black-box reasoning processes of such deep models and increase trust in their decisions.
+XAI methods can be broadly divided into two categories: First, Post-Hoc Methods like LIME (Ribeiro et al., 2016), SHAP (Lundberg & Lee, 2017) or GradCAM (Selvaraju et al., 2017) offer explanations for predictions without requiring retraining.
+While applicable in many scenarios, post-hoc methods may not actually align with the models’ decision-making processes, potentially leading to interpretations that are not entirely faithful (Rudin, 2019).
 Second, inherently interpretable methods provide built-in, case-based reasoning processes.
@@
 However, they are constrained in their representational power.
+Figure 1: Different Prototype Formulations. HyperPG is a novel formulation for probabilistic prototypes on a Hypersphere.
 Deep Prototype Learning Architectures such as ProtoPNet (Chen et al., 2019) and its derivatives (e.g., Rymarczyk et al., 2020; Donnelly et al., 2021; Sacha et al., 2023) integrate inherent interpretability into deep learning models through a prototype layer.
 Each neuron in this layer represents a prototype, storing a latent feature vector.
-The model’s predictions are based on the distances between sample features and prototype parameters, for example by computing the $L_{2}$-distance.
-However, these deterministic similarity scores do not include statistical information like confidence.
-To include such contextual information, prototypes could be modeled as probability distributions, like Gaussian distributions.
-Yet, recent prototype architectures favor the cosine hypersphere as feature space for its classification advantages (Mettes et al., 2019), where defining a Gaussian distribution analogue is challenging (Hillen et al., 2017).
-The prototypes in ProtoPNet and its successors are learned in an opaque manner via backpropagation.
-After training, the model “pushes” the learned prototypes on the embedding of the most similar known training sample (Chen et al., 2019).
-However, these prototypes are learned without human control and do not incorporate domain knowledge.
-A new training regime is needed to aligned prototypes with human-defined concepts.
-The main contributions of this paper are as follows:
-  • HyperPg: A new prototype representation with learned parameters anchor $\alpha$, mean $\mu$ and standard deviation $\sigma$.
+The model’s predictions are based on the distances between sample features and prototype parameters, for example, by computing the $L_{2}$-distance.
+This makes prototype learning essentially a clustering task in latent space (Zhou et al., 2022).
+Different prototype formulations, such as prototypes using the cosine similarity, induce a hyperspherical structure to the latent space with performance advantages to classification (Mettes et al., 2019).
+Probabilistic formulations such as Gaussian prototypes enable further downstream tasks relying on probability values such as Bayesian approaches.
+With the increased popularity of prototypical part models, there has been a corresponding rise in the number of architectural choices.
+However, to the best of our knowledge, there is no comprehensive overview of the design choices and prototype formulations for learning interpretable prototypes and their effect on predictive performance.
+Table 1: Overview of 14 related models and their prototype learning configurations. 8/14 models use Euclidean prototypes.
+Model Similarity Shape Assignment Clf. Head
+Prototype Decoder Euclidean Entire Image Class Exclusive FCL Li et al. (2018)
+ProtoPNet Euclidean Patch Class Exclusive FCL Chen et al. (2019)
+Def. ProtoPNet Cosine Spatial Arrangement Class Exclusive FCL Donnelly et al. (2021)
+ProtoPShare Euclidean Patch Merged after Training FCL Rymarczyk et al. (2020)
+ProtoPool Euclidean Patch Pooling Shared FCL Rymarczyk et al. (2022)
+TesNet Grassman Patch Class Exclusive FCL Wang et al. (2021)
+ProtoTree Euclidean Patch Shared Decision Tree Nauta et al. (2021)
+ProtoKNN Cosine Patch Class Exclusive KNN Clf Ukai et al. (2023)
+PIPNet Cosine Patch Shared FCL Nauta et al. (2023)
+LucidPPN Euclidean Separate Color & Texture Class Exclusive Branch Aggregation Pach et al. (2024)
+ProtoSeg Cosine Patch Class Exclusive FCL Zhou et al. (2022)
+ProtoGMM Gaussian Patch Class Exclusive FCL Moradinasab et al. (2024)
+MGProto Gaussian Patch Class Exclusive Bayesian Likelihood Wang et al. (2024b)
+ProtoPFormer Euclidean Transformer Token Class Exclusive Branch Aggregration Xue et al. (2024)
+This paper explores different prototype formulations. Its contributions are as follows:
+Prototype Formulation Overview: A comprehensive overview of different, learnable prototype definitions, including both point based and probabilistic prototypes as shown in Figure 1.
+HyperPG: A new probabilistic prototype representation with learned parameters anchor $\alpha$, mean $\mu$ and variance $\sigma^{2}$.
 This representation models a Gaussian distribution over cosine similarities, thereby projecting a Gaussian distribution on the surface of a hypersphere.
 HyperPG’s similarity score is based on the Gaussian’s probability density function and adapts its size through a learned standard deviation.
-  • HyperPgNet: A new prototype learning architecture built on HyperPg prototypes.
-HyperPgNet learns prototypes aligned with human-defined concepts using pixel level annotations.
-These concepts describe features such as color and patterns of birds or car parts.
-  • Concept Extraction Pipeline: Based on Grounding DINO (Liu et al., 2023) and SAM2 (Ravi et al., 2024), this pipeline provides pixel-level concept annotations at scale.
-  • Classification Experiments: HyperPgNet is compared with other prototype learning architectures like ProtoPNet on the CUB-200-2011 (Wah et al., 2011) and Stanford Cars (Krause et al., 2013) datasets.
-HyperPgNet outperforms the other models with fewer learned prototypes, while infusing each prototype with more meaning.
+Benchmarking Protoypes: Extensive image classification experiments based on prototypical part network architectures like ProtoPNet on the CUB-200-2011 (Wah et al., 2011), Stanford Cars (Krause et al., 2013) and Oxford Flowers (Nilsback & Zisserman, 2008) datasets.
 2 Related Work
@@
 The prototypes are each associated with a single class and learned via backpropagation without additional information.
-HyperPgNet builds on the idea of prototypical parts, but learns the prototypes in a transparent, concept-aligned manner.
-The prototypes are not class exclusive, but class shared.
-Each prototype corresponds to a human defined concept.
-To provide the required concept annotations, we propose a labeling pipeline based on foundation models such as DINOv2 (Oquab et al., 2024) and SAM2 (Ravi et al., 2024).
-HyperPgNet learns the concept-aligned prototypes by only adapting the loss functions and providing additional annotations, without requiring changes to the feature encoder like PIPNet (Nauta & Seifert, 2023) or Lucid-PPN (Pach et al., 2024), which uses a hybrid input head to disentangle prototype color and shape.
+The similarity of the prototypes to the image patch is based on the Euclidean distance.
 Multiple successors build on the idea of ProtoPNet.
-ProtoPShare (Rymarczyk et al., 2020) and ProtoPool (Rymarczyk et al., 2022) use additional optimizations to learn class-shared prototypical parts.
+ProtoPShare (Rymarczyk et al., 2020) merges the class-exclusive prototypes to class-shared ones.
+ProtoPool (Rymarczyk et al., 2022) directly learns class-shared patch prototypes and pools them by learning a slot assignment.
 Deformable ProtoPNet (Donnelly et al., 2021) learns a mixture of prototypical parts with dynamic spatial arrangement.
-Other work replace the linear output layer of ProtoPNet with other interpretable models: ProtoKNN (Ukai et al., 2023) uses a k-nearest neighbor classifier and ProtoTree (Nauta et al., 2021) employs a decision tree.
-All these models change how the prototype activations are further processed.
-However, they all rely on the point-based prototype formulation introduced by ProtoPNet based on the $L_{2}$ or cosine similarity.
-HyperPgNet proposes a novel prototype formulation based on Gaussian distributions on the hypersphere, without changing the downstream processing.
-This probabilistic view with learned mean and variance allows HyperPgNet to learn prototypes with different degrees of specialization.
-Furthermore, the statistical confidence derived from the learned variance could be integrate in further downstream processing.
-The view of prototype learning as clustering in latent space was introduced for image segmentation (Zhou et al., 2022).
-The image patches are points in latent space with the prototypes as cluster centers.
-ProtoGMM (Moradinasab et al., 2024) and MGProto (Wang et al., 2023) build on this idea to model the prototypes as Gaussian mixture models for image segmentation.
-HyperPg adapts the idea of Gaussian prototypes used in image segmentation to hyperspherical latent spaces based on the cosine similarity.
-The cosine similarity has been shown to perform well in classification tasks (Mettes et al., 2019) and also been applied in recent prototype learning works (Zhou et al., 2022; Ukai et al., 2023).
-This makes HyperPgNet the first model to employ probabilistic prototypes with learnable mean and variance in a hyperspherical space.
-Concept-Based Methods.
-Testing with Concept Activation Vectors (TCAV) (Kim et al., 2018) is a post-hoc method that aligns model activations with human-understandable concepts.
-In image classification, TCAV measures model activations on images representing a specific concept like “striped” and compares them to activations on a random image set.
-Concept Activation Regions (CAR) (Crabbé & van der Schaar, 2022) relax TCAV’s assumption of linear separability in latent space by using support vector machines and the kernel trick.
-Another post-hoc approach, CRAFT (Fel et al., 2023), generates concept attribution maps by identifying relevant concepts from the training set via Non-Negative Matrix Factorization (NMF) and backpropagating with GradCAM (Selvaraju et al., 2017).
-Instead of analyzing the learned concepts post-hoc, some approaches integrate concept learning into the classifier.
-MCPNet (Wang et al., 2024) extracts concepts from multiple layers by splitting feature maps along the channel dimension, learning concepts during training and at multiple scales.
-Concept Bottleneck Models (CBMs) (Koh et al., 2020) augment black-box models to predict pre-defined concept labels in an intermediate step.
-In this manner, CBMs behave like two sequential black box models and their internal reasoning remains opaque.
-HyperPgNet combines the concept-aligned learning approach of CBM with interpretable prototype learning.
-HyperPgNet learns multiple HyperPg prototypes per concept, instead of a fully connected concept prediction layer.
-A novel concept extraction pipeline, built on foundation models, provides pixel level annotations at scale.
-By using those annotations to learn concept prototypes, HyperPgNet avoids the ambiguity of aligning post-hoc explanations with the model’s internal reasoning.
-3 Prototypical Gaussians on the Hypersphere
+For computing the similarity between the prototypes and image patches, Deformable ProtoPNet uses the cosine similarity.
+TesNet (Wang et al., 2021) proposes to compute the prototype similarity on the Grassman Manifold.
+All these approaches share the use of a linear classification layer.
+However, ProtoTree (Nauta et al., 2021) builds a Decision Tree on the learned Euclidean prototypes and ProtoKNN (Ukai et al., 2023) proposes to use a $k$-nearest-neighbor classifier with cosine prototype similarities.
+Recent work proposes to learn more interesting or robust features for the prototypes from the image encoders.
+PIPNet (Nauta et al., 2023) use augmentations during training to align the learned prototypes with more meaningfull image content.
+LucidPPN (Pach et al., 2024) propose a hybrid architecture with texture only grayscale prototypes and low resolution color only prototypes.
+Zhou et al.
+(2022) propose to view prototype learning as a clustering task in the network’s latent space for image segmentation.
+In this setting, the clustering can be done outside of the network’s backpropagation path with an external clustering algorithm, so-called non-learnable prototypes.
+Later work such as ProtoGMM (Moradinasab et al., 2024) and MGProto (Wang et al., 2024b) build on this idea and model the prototypes as Gaussian distributions.
+ProtoPFormer (Xue et al., 2024) adapts Euclidean prototypical part networks to Transformer based architectures by using a VisionTransformer (ViT, Dosovitskiy, 2020) backbone.
+To avoid the distraction problem due to the global attention mechanism of ViT architectures, ProtoPFormer uses a hybrid approach with global and local prototypes.
+The local prototypes are encouraged to focus on different parts of the image by modeling the spatial prototype activations as a 2D Gaussian.
+Table 1 presents a systematic overview of the discussed works.
+The cosine similarity has been shown to perform well in classification tasks (Mettes et al., 2019), but only few prototype learning works have used them so far (Donnelly et al., 2021; Zhou et al., 2022; Ukai et al., 2023).
+There is also a lack for probabilistic prototypes in a hyperspherical space.
+We introduce HyperPG, which combines the predictive performance gains from the cosine similarity with the probabilistic nature of Gaussian distributions.
+3 Overview: Prototype Formulations
 Figure 2: Illustration on how the different prototype formulations compute the similarity between a prototype ${\bm{p}}$ and latent vector ${\bm{z}}$.
-$L_{2}$ prototypes compute the Euclidean distance between two points in latent space.
+Euclidean prototypes compute the $L_{2}$ distance between two points in latent space.
 Hyperspherical prototypes use the cosine similarity of normalized vectors, which corresponds to the angle between two points on a hypersphere.
@@
 Prototype learning for deep neural networks involves finding structures in latent space representations.
-This section provides an overview of existing methods, which are also illustrated in Fig.
-1.
+This section provides an overview of existing methods, which are also illustrated in Figure 2.
 Prior work uses point-based prototypes, computing similarity scores relative to a single point in latent space.
-Aligning prototypes with human-labeled concepts requires a more powerful prototype representation, leading to the introduction of HyperPg - Prototypical Gaussians on the Hypersphere.
-HyperPg prototypes are able to adapt their shape in the latent space by modeling Gaussian distributions with mean and standard deviation, adapting to the variance in encoded concept features.
+On the other hand, probabilistic formulations such as Gaussian prototypes or HyperPG prototypes allow the model to adapt to the variance in the training data.
 3.1 Point Based Prototypes
@@
 Commonly, the prototypes ${\bm{p}}$ are also part of $\mathcal{Z}$ with shape $\rho_{w}\times\rho_{h}\times D$, i.e., spatial size $\rho=\rho_{w}\rho_{h}$.
-Autoencoder approaches use $\rho=\zeta$, meaning the prototype represents the entire image and can be reconstructed from latent space (Li et al., 2018).
+Early approaches based on Autoencoder architectures (Li et al., 2018) use $\rho=\zeta$, meaning the prototype represents the entire image and can therefore be decoded out from latent space back to the input space.
 Part-based approaches like ProtoPNet and segmentation models like ProtoSeg use $\rho=1$ (Chen et al., 2019; Zhou et al., 2022), meaning each prototype represents some part of the image.
-Notable exceptions include Deformable ProtoPNet (Donnelly et al., 2021), where each prototype has spatial size $\rho=3\times 3$ and MCPNet (Wang et al., 2024), where prototypes are obtained by dividing the channels of the latent space into chunks.
+Notable exceptions include Deformable ProtoPNet (Donnelly et al., 2021), where each prototype has is a $\rho=3\times 3$ arrangement of smaller patches, and MCPNet (Wang et al., 2024a), which learns concept prototypes across the latent features.
 The prediction is computed by comparing each prototype ${\bm{p}}$ to the latent feature map ${\bm{z}}$.
 For simplicity’s sake lets assume the spatial dimensions $\rho=\zeta=1$.
-The following equations can be adapted for higher spatial dimensions by summing over $\sum_{\rho_{w}}\sum_{\rho_{h}}$ for each chunk of the latent map.
-ProtoPNet’s prototypes leverage the $L_{2}$ similarity. The $L_{2}$ similarity measure is defined as
+The following equations can be adapted for higher spatial dimensions by summing over the height and width $\sum_{\rho_{w}}\sum_{\rho_{h}}$ for each patch of the latent map.
+3.1.1 Euclidean Prototypes
+ProtoPNet’s Euclidean prototypes leverage the $L_{2}$ similarity. The $L_{2}$ similarity measure is defined as
 $s_{L_{2}}({\bm{z}}|{\bm{p}})=\log\left(\frac{\|{\bm{z}}-{\bm{p}}\|^{2}_{2}+1}{\|{\bm{z}}-{\bm{p}}\|^{2}_{2}+\epsilon}\right)$ (1)
 and is based on the inverted $L_{2}$ distance between a latent vector ${\bm{z}}$ and a prototype vector ${\bm{p}}$.
-This similarity is a point-based measure, as only the two vectors are compared, without any additional context like the expected variance of the cluster represented by the prototype.
-Hyperspherical prototypes using the cosine similarity have been shown to perform well in classification tasks (Mettes et al., 2019) and have been widely used since (e.g., Zhou et al., 2022; Ukai et al., 2023).
+This similarity measure starts to perform worse with higher numbers of dimensions, as in a large enough space all points are distant to each other and the $L_{2}$ distance looses meaning.
+3.1.2 Cosine Prototypes
+Prototype formulations based on the cosine similarity create a hyperspherical space.
+They have been shown to perform well in classification tasks (Mettes et al., 2019) and recent works have started to apply them (e.g., Donnelly et al., 2021; Zhou et al., 2022; Nauta et al., 2021; Ukai et al., 2023).
 The cosine similarity is defined as
@@
 The cosine similarity is defined on the interval $[-1,1]$ and measures: $1$ for two vectors pointing in the same direction, 00 for orthogonal vectors, and $-1$ for vectors pointing in opposite directions.
-Like the $L_{2}$ similarity, the cosine similarity is a point-based measure comparing only two vectors.
-Both the $L_{2}$ and cosine similarity can be used for classification.
+Without the normalization to unit length, we can define a scaled dot-product prototype inspired by the Attention mechanism (Vaswani et al., 2017).
+However, this similarity measure is not bounded to any interval, as the vector length is proportionally preserved.
+For more details see Appendix C.
+Like the $L_{2}$ similarity, the cosine similarity is a point-based measure comparing two vectors directly.
+Both the $L_{2}$ and cosine similarity have been used for classification tasks.
 The similarity scores are processed by a fully connected layer (e.g., Chen et al., 2019; Donnelly et al., 2021), or a winner-takes-all approach assigns the class of the most similar prototype (e.g., Sacha et al., 2023).
-Prototypes can be learned by optimizing a task-specific loss, such as cross-entropy, via backpropagation.
+The prototypes can be learned by optimizing a task-specific loss, such as cross-entropy, via backpropagation.
 Alternatively, Zhou et al.
-(2022) propose “non-learnable” prototypes, whose parameters are obtained via clustering in the latent space rather than backpropagation.
-3.2 Gaussian Prototypes
-Gaussian prototypes model prototypes as a Gaussian distribution with mean and covariance.
-They adapt to the spread of the associated latent cluster by adjusting their covariance matrix.
-Thus, a Gaussian prototype with a wide covariance can still have a relatively high response even for larger distances from the mean vector.
+(2022) propose “non-learnable” prototypes, whose parameters are obtained via a clustering operation in the latent space instead of backpropagation.
+3.2 Probabilistic Prototypes
+Viewing prototype learning as clustering problem, point-based prototypes compute the similarity of a latent vector to the center coordinate of each cluster, as represented by the prototype vector.
+Probabilistic prototypes aim to model the cluster as a probability distribution.
+Future work could use probabilistic prototypes for extended downstream tasks such as outlier detection or interventions by sampling from generative distributions.
+3.2.1 Gaussian Prototypes
+Gaussian prototypes model the clusters in latent space as a Gaussian distribution with mean and covariance.
+By adapting the covariance matrix to the training data a Gaussian prototype with a wide covariance can still have a relatively high response even for larger $L_{2}$ distances from the mean vector.
+On the other hand, with a very small covariance matrix a Gaussian prototype could show no response unless the mean value is met nearly exactly.
 Let the formal definition of a Gaussian prototype be ${\bm{p}}^{G}=(\bm{{\mu}},\bm{\mathrm{\Sigma}})$.
@@
 Each Gaussian prototype ${\bm{p}}^{G}$ thus defines a multivariate Gaussian Distribution ${\mathcal{N}}(\bm{{\mu}},\bm{\mathrm{\Sigma}})$.
-Gaussian prototypes can be trained using EM for clustering in the latent space (Zhou et al., 2022; Wang et al., 2023; Moradinasab et al., 2024) or by directly optimizing the parameters via Backpropagation.
+Gaussian prototypes can be trained using EM for clustering in the latent space (Zhou et al., 2022; Moradinasab et al., 2024; Wang et al., 2024b) or naively by directly optimizing the parameters via Backpropagation.
 The similarity measure of Gaussian prototypes is defined as the probability density function (PDF) for $D$-dimensional multivariate Gaussians, namely
 $s_{\mathrm{Gauss}}({\bm{z}}|{\bm{p}}^{G})$ $=\mathcal{N}({\bm{z}};\bm{{\mu}},\bm{\mathrm{\sigma}})$ (3)
+$=\frac{1}{(2\pi)^{\frac{D}{2}}|\bm{\mathrm{\Sigma}}|^{\frac{1}{2}}}$ $\exp\left(-\frac{1}{2}({\bm{z}}-\bm{{\mu}})^{\top}\bm{\mathrm{\Sigma}}^{-1}({\bm{z}}-\bm{{\mu}})\right).$ (4)
+It is worth noting that the term inside the exponent is based on the $L_{2}$ distance between the latent vector ${\bm{z}}$ and the mean $\bm{{\mu}}$.
 This formulation as a PDF has several advantages: A) The similarity can be interpreted as the likelihood of being sampled from the Gaussian prototypes, which is more meaningful than a distance metric in a high-dimensional latent space.
 B) Prototypes can adapt their shape using a full covariance matrix, allowing different variances along various feature dimensions, offering more flexibility in shaping the latent space.
-However, this increases computational requirements, especially with EM clustering.
-Gaussians prototypes lose the advantages of hyperspherical prototypes for classification and regression Mettes et al.
-(2019).
-3.3 Gaussian Prototypes on the Hypersphere - HyperPg
-Prototypical Gaussians on the Hypersphere (HyperPg) combine the advantages of Gaussian and hyperspherical prototypes.
+However, learning a full covariance matrix increases computational requirements, especially with EM clustering.
+(a) $\mu=1$
+(b) $\mu=0.5$
+(c) $\mu=0$
+Figure 3: HyperPG prototypes learn a distribution of cosine similarities.
+They are parameterized by a learnable anchor vector $\bm{{\alpha}}$, scalar mean $\mu$ and scalar variance $\sigma^{2}$.
+HyperPG projects a Gaussian distribution of cosine similarities on the surface of a hypersphere, resulting in ring shaped activation patterns around the anchor vector.
+Figure 4: Prototype Learning Architecture.
+The HyperPG module can be easily exchanged to other prototype formulations such as Euclidian or Cosine prototypes.
+HyperPG uses a Gaussian distribution as Density Estimator, but other PDFs are possible.
+3.2.2 Gaussian Prototypes on the Hypersphere - HyperPG
+Just as Gaussian prototypes provide a probabilistic formulation for Euclidean prototypes, we aim to develop a probabilistic formulation for hyperspherical prototypes based on the cosine similarity.
+We name this new formulation Prototypical Gaussians on the Hypersphere (HyperPG).
 HyperPG prototypes are defined as ${\bm{p}}^{H}=(\bm{{\alpha}},\mu,\sigma)$ with a directional anchor vector $\bm{{\alpha}}$, scalar mean similarity $\mu$ and scalar standard deviation (std) $\sigma$.
@@
 $=\frac{{\mathcal{N}}\left(s_{\cos}({\bm{z}}|\bm{{\alpha}});\mu,\sigma\right)}{\mathcal{G}(1,\mu,\sigma)-\mathcal{G}(-1,\mu,\sigma)}.$ (6)
-Figure 2: HyperPg learns a Gaussian distribution with mean $\mu$ and std $\sigma$ of cosine similarities with the anchor $\bm{{\alpha}}$.
-The plots visualize HyperPg’s activations on the $3D$ hypersphere with anchor $\bm{{\alpha}}=(0,0,1)$, std $\sigma=0.1$ and and various mean cosine similarities $\mu\in[-1,1]$ for 10k random samples.
-The anchor $\bm{{\alpha}}$ is shown as a red arrow, and std $\sigma$ governs the width of the distribution.
-For $\mu=1$, the distribution is aligned with the vector; for $\mu=-1$ it is on the opposite side of the hypersphere.
-Due to the cosine similarity, setting $11>\mu>-$ 1 interpolates the area of highest probability density between both poles, leading to a ring shape on the hypersphere’s surface.
-Fig.
-2 illustrates the activations of HyperPg’s similarity function on the surface of a $3D$ hypersphere with anchor $\bm{{\alpha}}=(0,0,1)$, fixed std $\sigma=0.1$ and various mean values $\mu\in[-1,1]$.
+Figure 3 illustrates the activations of HyperPG’s similarity function on the surface of a $3D$ hypersphere with anchor $\bm{{\alpha}}=(0,0,1)$, fixed std $\sigma=0.1$ and various mean values $\mu\in[0,1]$.
 The anchor $\bm{{\alpha}}$ defines a prototypical direction vector in latent space $\mathcal{Z}$, similar to other hyperspherical prototypes, and is visualized as a red arrow.
@@
 For $\mu=1$, the distribution centers around the anchor, as the cosine similarity is $1$ if two vectors point in the same direction.
+This corresponds to the von Mises-Fisher (vMF) distribution (Appendix B).
 For $\mu=-1$, the distribution is the on opposite side of the hypersphere, as the cosine similarity is $-1$ for vectors pointing in opposite directions.
@@
 Since no specific direction is indicated, the entire hyperplanar segment orthogonal to the anchor has the highest activation.
-This activation pattern for the cosine similarity would typically require an infinite mixture of prototype vectors pointing in all directions in this hyperplane.
+These ring-like activation patterns would require an infinite mixture of prototype vectors pointing in all directions in this hyperplane.
 HyperPG achieves the same effect by learning only one prototype vector (the anchor) and just two additional scalar parameters.
-This significantly increases HyperPg’s representational power compared to standard hyperspherical and Gaussian prototypes and is a major difference to the von Mises-Fisher distribution (see also Appendix A).
-Combining the strengths of hyperspherical and Gaussian prototypes, HyperPg can learn more complex structures in the latent space, such as human-defined concepts, and provide a more meaningful similarity measure.
+This significantly increases HyperPG’s representational power compared to Gaussian prototypes without increasing the computational complexity.
+This potential capability is also a major difference to the vMF distribution.
 HyperPG can be easily adapted to other probability distributions with additional desirable properties.
-Possible candidate distributions are elaborated on in Appendix A.
-Similarly, it is possible to exchange the cosine similarity to other similarity measures or functions, and learn an untruncated PDF over their output, making the HyperPg idea transferable to applications outside of prototype learning.
+Possible candidate distributions are elaborated on in Appendix B.
+Similarly, it is possible to exchange the cosine similarity to other similarity measures or functions, and learn an untruncated PDF over their output, making the HyperPG idea transferable to other manifolds and applications outside of prototype learning.
 4 Training
 The original ProtoPNet implementation uses three loss functions: a task specific loss like crossentropy for classification, a cluster loss to increase compactness within a class’s cluster, and a separation loss to increase distances between different prototype clusters.
-However, the prototypes of ProtoPNet and its successors are learned in a black-box manner.
-HyperPgNet is a new inherently interpretable deep learning approach built on HyperPg prototypes.
-It introduces a “Right for the Right Concept” loss, inspired by “Right for the Right Reasons” (Ross et al., 2017), to restrict the learned prototypes to human-defined concepts.
-This focus enhances the interpretability and minimizes the influence of confounding factors.
-This section first provides an overview of the used prototype learning losses, then introduces the Right for the Right Concept loss, and finally discusses the overall network architecture and final multi-objective loss.
 4.1 Prototype Losses
@@
 where $s(\cdot,\cdot)$ is some similarity measure.
-The $L_{\mathrm{Clst}}$-Loss function increases compactness by increasing the similarity between prototypes ${\bm{p}}_{c}$ of class $c$ latent embeddings ${\bm{z}}_{c}$ of class $c$ over all samples.
+The $L_{\mathrm{Clst}}$-Loss function increases compactness by increasing the similarity between prototypes ${\bm{p}}_{c}$ and latent embeddings ${\bm{z}}_{c}$ of class $c$ over all samples.
 An additional separation loss increases the margin between different prototypes. The separation loss function is defined as
@@
 The $L_{\mathrm{Sep}}$ function punishes high similarity values between a latent vector ${\bm{z}}_{c}$ of class c and prototypes ${\bm{p}}_{\neg c}$ not belonging to $c$, thereby separating the clusters in latent space.
-Please note, ProtoPNet (Chen et al., 2019) use a slightly different notation by working with the $L_{2}$ distance, instead of a similarity measure.
-In HyperPgNet the learned prototypes are not class exclusive, but shared among different classes.
-Instead, each prototype is assigned to one human-defined concept $k\in K$.
-The Concept Activation Regions (CAR, Crabbé & van der Schaar, 2022) model proposes a concept density loss defined as
-$L_{Density}$ $=-\frac{1}{|K|}\sum_{k\in K}\frac{1}{|{\bm{P}}_{k}|}\sum_{{\bm{p}}_{k}\in{\bm{P}}_{k}}\phi({\bm{p}}_{k},{\bm{Z}}_{k})-\phi({\bm{p}}_{k},{\bm{Z}}_{\neg k})$ (8)
-$\phi({\bm{p}},{\bm{Z}})$ $=\frac{1}{|{\bm{Z}}|}\sum_{{\bm{z}}\in{\bm{Z}}}s({\bm{p}},{\bm{z}}).$ (9)
-The density loss uses the similarity aggregation $\phi({\bm{p}},{\bm{Z}})$ which computes the mean response of a prototype ${\bm{p}}$ with a set of latent features ${\bm{Z}}$.
-The density loss therefore computes the mean response of correctly assigned prototypes minus the mean response of incorrectly assigned prototypes.
-The loss functions proposed by ProtoPNet compute the maximum response of correctly and incorrectly assigned prototypes instead.
-4.2 Right Concept Loss
-To ensure prototypes actually correspond to the input pixels containing the concept, and do not respond to other factors in the background, HyperPgNet introduces the “Right for the Right Concept” (RRC) -Loss inspired by “Right for the Right Reasons” (RRR, Ross et al., 2017).
-The RRC loss is defined as
-$L_{\mathrm{RRC}}=\frac{1}{N}\sum_{i=1}^{N}\sum_{k\in K}\left({\bm{A}}_{{\bm{x}}_{i},k}\frac{\partial}{\partial{\bm{x}}_{i}}\sum_{{\bm{p}}_{k}\in{\bm{P}}_{k}}s\left({\bm{p}}_{k},\mathrm{Enc}\left({\bm{x}}_{i}\right)\right)\right)^{2},$ (10)
-with binary annotation matrix ${\bm{A}}_{{\bm{x}}_{i},k}\in\{0,1\}^{N\times\times W\times H}$ for each input sample ${\bm{x}}_{i}$ and concept $k$.
-This annotation matrix defines for each input image, which pixels contain which concept.
-In the original RRR paper, the annotation matrix specified relevant regions for the classification task, steering the model’s activations away from confounding factors in the background.
-In HyperPgNet the RRC-Loss further strengthens the prototype-concept association.
+Please note, the original ProtoPNet paper uses a slightly different notation by working with the $L_{2}$ distance, instead of a similarity measure.
 4.2 Multi-Objective Loss Function
 To train a prototype learning network like HyperPGNet for downstream tasks like image classification, a multi-objective loss function is employed. This multi-objective loss function is defined as
-$L=L_{\mathrm{CE}}+\lambda_{\mathrm{Clst}}L_{\mathrm{Clst}}+\lambda_{\mathrm{Sep}}L_{\mathrm{Sep}}+\lambda_{\mathrm{RRC}}L_{\mathrm{RRC}},$
+$L=L_{\mathrm{CE}}+\lambda_{\mathrm{Clst}}L_{\mathrm{Clst}}+\lambda_{\mathrm{Sep}}L_{\mathrm{Sep}},$
 where $L_{\mathrm{CE}}$ is the cross-entropy loss over network predictions and ground truth image labels.
-ProtoPNet uses $\lambda_{\mathrm{Clst}}=0.8$ and $\lambda_{\mathrm{Sep}}=0.08$ (Chen et al., 2019).
-For HyperPgNet, the different loss terms are weighted equally, meaning all $\lambda=1$.
+Our experiments uses $\lambda_{\mathrm{Clst}}=0.8$ and $\lambda_{\mathrm{Sep}}=0.08$ as proposed by ProtoPNet (Chen et al., 2019).
 4.3 Network Architecture
-Figure 3: HyperPgNet Architecture.
-The HyperPg module can be easily exchanged to other prototype formulations such as ProtoPNet.
-HyperPgNet uses the truncated Gaussian distribution as density estimator, but other PDFs are possible.
-Fig.
-3 illustrates HyperPgNet’s Architecture for interpretable image classification.
-HyperPgNet uses a pretrained feature encoder such as ConvNext-tiny (Liu et al., 2022) or MiT-B4 (Xie et al., 2021) as a backbone model.
-A neck consisting of two $1\times 1$ convolution layers with ReLU activation in between projects the high dimensional feature map of the backbone into a lower dimensional feature space.
-The HyperPg prototype module is implemented with two layers: First, a prototype learning layer computes the cosine similarity of the learnable HyperPg anchors $\bm{{\alpha}}$ to the latent vectors produced by the neck.
+For our experiments, we use the network architecture proposed by ProtoPNet (Chen et al., 2019), which we deem the basic deep prototype learning architecture for image classification.
+Figure 4 illustrates this network architecture with the HyperPG prototype module.
+In the encoding stage, a pretrained image encoder such as a ResNet model (He et al., 2016) serves as model backbone.
+Then, a convolution neck projects the high dimensional feature map of the backbone into the lower dimensional prototype space.
+This neck consists of two $1\times 1$ convolution layers with ReLU activation in between followed by a Sigmoid activation.
+In the prototype stage, the prototype module computes the similarity scores.
+The original ProtoPNet uses Euclidian prototypes, which compute the $L_{2}$ distance between the encoded feature maps and the prototype parameters.
+Figure 4 shows the make up of the HyperPG module: First, the prototype module computes the cosine similarity of the learnable HyperPG anchors $\bm{{\alpha}}$ to the latent vectors produced by the neck.
 Second, a Density Estimation layer with learnable parameters mean $\mu$ and std $\sigma$ computes the Gaussian PDF over the activations of the previous layer.
-Future extensions can adapt both components independently, for example by implementing a hyperbolic similarity measure or a multi-modal probability distribution.
-Finally, the output of the HyperPg Module is passed through a single fully connected layer (FCL) to produce the output logits or class scores.
-5 Concept Extraction at Scale
-To learn prototypes based on human-defined concepts with RRC-Loss for image classification, pixel level annotations are required.
-Foundation models are leveraged to generate these annotations, reducing the need for human labeling.
-The CUB-200-2011 (Wah et al., 2011) annotates each image with up to 15 part locations (e.g., head, tail, wing).
-Each part is labeled with some attribute describing the coloring, pattern or shape.
-A Segment Anything 2 (SAM2) model (Ravi et al., 2024) creates pixel-level segmentation masks based on the part locations.
-The Stanford Cars dataset (Krause et al., 2013) does not provide part level annotations.
-We defined a list of 10 car parts (e.g., wheel, headlight, radiator) and use a Grounding DINO model (Liu et al., 2023) to generate bounding boxes for each part.
-Finally, SAM2 creates pixel-level segmentation masks based on the bounding boxes.
-This automated pipeline labeled the 16k images of Stanford Cars within 2h on a NVIDIA 4060Ti with 16GB VRAM, and can be easily adapted to other datasets.
-Fig.
-4 presents some examples for CUB-200-2011 and Stanford Cars.
-(a) Concept Segmentation Masks for CUB-200-2011 based on provided positive and negative part annotations.
-(b) Concept Segmentation Mask for Stanford Cars generated from a list of human-defined car parts.
-Figure 4: Examples for Automated Concept Extraction.
+Future extensions could adapt both components independently, for example by implementing a hyperbolic similarity measure or a multi-modal probability distribution.
+The final classification stage produces the output logits based on the prototype module’s similarity scores. ProtoPNet uses a single linear layer due to its inherent interpretability.
+4.3.1 Comparison to Recent Works
+Our experiments use only the basic ProtoPNet architecture and focus on the effect of different prototype formulations.
+This section provides a short overview of the required changes to this basic architecture in order to implement the models described in Table 1.
+LucidPPN (Pach et al., 2024) replaces the encoding stage with a hybrid encoder specializing in color-only and texture-only feature extraction.
+Deformable ProtoPNet (Donnelly et al., 2021) makes changes towards the input of the prototype module to allow for dynamic spatial arrangements of the prototypes, i.e., “deformations”.
+ProtoPool (Rymarczyk et al., 2022) makes changes towards the output of the prototype module, by also learning how to dynamically pool the different prototype activations.
+ProtoTree (Nauta et al., 2021) and ProtoKNN (Ukai et al., 2023) propose to use a different inherently interpretable model for the classification stage.
+ProtoPFormer (Xue et al., 2024) builds exclusively on a Vision Transformer backbone and uses different network branches for the global information encoded in the class token and the local information in the image tokens.
 5 Experiments
-The experiments were performed on two datasets: CUB-200-2011 (CUB) with 200 bird species Wah et al.
-(2011) and Stanford Cars (CARS) with 196 car models Krause et al.
-(2013).
-Implementation details such as data preprocessing or model implementation are provided in Appendix B.
-Birds Cars
-# Prototypes MiT-B4 ConvNeXt # Prototypes MiT-B4 ConvNeXt
-BB Baseline - 17.7 74.2 - 1.9 57.5
-CBM - 75.7 77.9 - 79.9 81.0
-PP ProtoPNet (PPN) 2000 68.0 68.1 1960 86.4 87.0
-PPN + HyperPG 2000 70.5 65.0 1960 87.4 77.3
-CAP HyperPgNet $-L_{\mathrm{RRC}}$ 300 76.5 76.9 180 88.6 88.9
-(Ours) HyperPgNet $+L_{\mathrm{RRC}}$ 300 74.1 71.5 180 81.2 86.0
-Table 1: Test Top-1 Accuracy. Models are grouped by their reasoning process. BB: Black Box. PP: Prototypical Parts. CAP: Concept Aligned Prototypes.
+The experiments were performed on CUB-200-2011 (CUB) with 200 bird species (Wah et al., 2011) with ResNet50 (He et al., 2016), DenseNet121 (Huang et al., 2017) and ViT-B16 (Dosovitskiy, 2020) backbones.
+Later experiments also explore the Stanford Cars (Krause et al., 2013) and Oxford Flowers (Nilsback & Zisserman, 2008) datasets.
+Implementation details such as data preprocessing or model implementation are provided in Appendix D.
+Table 2: Mean Top-1 Test Accuracies and Std for different prototype formulations with ResNet50 Backbones averaged over 3 random seeds.
+ResNet50 Dense121 ViT-B16
+Baseline 79.2 $\pm$ 0.2 75.5 $\pm$ 0.2 78.5 $\pm$ 0.8
+Euclidian (ProtoPNet) 35.3 $\pm$ 1.4 33.6 $\pm$ 3.0 9.4 $\pm$ 0.4
+Gaussian 61.1 $\pm$ 0.2 45.9 $\pm$ 5.1 4.2 $\pm$ 1.3
+Cosine 71.7 $\pm$ 0.4 68.7 $\pm$ 0.5 69.2 $\pm$ 3.9
+HyperPG (Ours) 74.3 $\pm$ 0.6 70.7 $\pm$ 0.3 72.9 $\pm$ 0.1
 5.1 Quantitative Results
-Figure 5: CUB Test Accuracy per Epoch.
-Table 1 reports the top-1 accuracy for the tested models on the CUB and CARS datasets with different backbones.
-HyperPgNet achieves comparable results with the black box CBM and baseline models.
-Notably, HyperPgNet outperforms CBM on CARS, as it can differentiate between different modes for the same concept by learning multiple prototypes per concept.
-CBM is restricted to a binary concept presence prediction.
-Including ${\mathcal{L}}_{\mathrm{RRC}}$ slightly lowers HyperPgNet’s performance but is still comparable to the black-box models.
-Fig.
-5 shows the test accuracy over epochs for CUB with ConvNeXt backbone.
-The concept aligned models CBM and HyperPgNet are the fastest to converge at around 100 epochs.
-ProtoPNet requires over 400 training epochs until the best performance level is reached.
-However, switching from the standard $L_{2}$ based prototypes to HyperPg prototypes effectively halves the required training epochs.
-6.2 Qualitative Comparison
-Fig.
-7 shows the pixel attribution of the highest activated prototype on each image for the tested prototype learning models with ConvNeXt backbone.
-In contrast to prior work (e.g., Chen et al., 2019; Ukai et al., 2023), this visualization is based on the prototype gradients, similar to saliency maps such as GradCAM (Selvaraju et al., 2017), which should avoid spatial misalignment often observed in prototype explanations (Sacha et al., 2024).
-The concept aligned prototypes learned by HyperPgNet focus more on the relevant parts of the image subject than the class-based prototypes of ProtoPNet.
-The network seems to focus correctly on either the overall color of the bird and ignoring the background (first three examples) or on specific markings (last three examples).
-The addition of the ${\mathcal{L}}_{\mathrm{RRC}}$ loss seems to further increase the main focus of the prototype activation on the relevant image regions.
-For ProtoPNet, no qualitative difference can be detected between the standard $L_{2}$ or HyperPg prototypes.
-With both prototype formulations, ProtoPNet is likely to also activate on parts of the image background.
-This is a fundamental weakness of training ProtoPNet in an opaque manner without alignment to relevant concepts, which HyperPgNet addresses.
-On the CARS dataset, the effect of ${\mathcal{L}}_{\mathrm{RRC}}$ is more apparent.
-The defined concepts for this dataset are limited to car parts without fine-grained differentiation between patterns or shapes, such as the concepts on CUB.
-For this dataset, one of the most predictive image regions across models seems to be the front bumper of the cars, which has not been annotated with its own concept.
-As shown in Fig.
-6, the concept-aligned prototypes still activate in this region.
-However, by training the model with the ${\mathcal{L}}_{\mathrm{RRC}}$ loss, the prototypes focus more on the annotated image regions.
-Figure 6: Concept prototype activation after training with and without ${\mathcal{L}}_{\mathrm{RRC}}$ with MiT-B4 backbone.
-Figure 7: Highest prototype activation per image with ConvNeXt backbone.
-6.3 Ablation
-# Prototypes CUB Accuracy
-CBM 71.6
-1 70.5
-5 68.9
-10 68.0
-20 67.3
-Table 2: Test Top-1 Accuracy on CUB with ResNet50 backbone with varying numbers of prototypes per concept.
-We conduct an ablation on the number of prototypes per concept.
-Table 2 presents the test accuracies on CUB for HyperPgNet with ResNet50 backbone.
-The difference in task performance over the different model configurations is negligible.
-However, looking at some qualitative examples as depicted in Fig.
-8 a difference in prototype quality can be observed.
-With fewer prototypes per concept, the model is more likely to also react to parts of the mage background in order to still achieve a low training error.
-With too many prototypes per concept, the latent space becomes too segmented.
-The prototypes start to focus on smaller subparts of the image subject, potentially decreasing the interpretability.
-The chosen configuration of 10 prototypes per concept seems to be a sweet spot regarding predictive performance and prototype quality for the experiments.
-Figure 8: Qualitative Results for HyperPgNet with ResNet50 backbone and varying numbers of prototypes per concept.
+Table 2 reports the mean top-1 accuracy and standard deviation for the tested models on the CUB dataset with backbone models ResNet50 (He et al., 2016), DenseNet121 (Huang et al., 2017) and ViT-B16 (Dosovitskiy, 2020) pretrained on ImageNet.
+Please note that in contrast to prior work, we do not report scores from model ensembles or using pretrained iNaturalist weights.
+The black-box baselines perform the best across all experiments, which we see as a theoretical upper bound for the performance.
+The classical Euclidean prototypes perform the worst across all experiments.
+A first model improvement can be achieved by switching to probabilistic Gaussian prototypes.
+The hyperspherical Cosine prototypes further improve the performance.
+HyperPG combines the advantages of Gaussian and Cosine prototypes, further closing the gap to the upper bound presented by the black-box baselines.
+Notably, both Euclidian and Gaussian prototypes are unable to learn on the latent features of ViT-B16.
+We theorize that the attention mechanism of the Transformer architecture produces a near-hyperspherical latent space.
+This highlights the importance of choosing the right prototype formulation for the used backbone.
+For example, language aligned models such as (Radford et al., 2021) use the cosine similarity for their alignment.
+We expect that only hyperspherical prototypes can produce meaningful results in this setting.
+Figure 5: CUB-200-2011 Test Accuracy per Epoch with ResNet50 backbone. Mean and std over 3 random seeds.
+Figure 5 shows the test accuracy over epochs for CUB with ResNet50 backbone.
+The standard Euclidean prototypes take the longest time (200 Epochs) until convergence while also achieving the lowest accuracy.
+By changing only the prototype formulation to a Gaussian one, the model achieves higher performance and converges at 100 epochs.
+Both the cosine and HyperPG formulation outperform the other formulations in model accuracy and training time, requiring only 30 training epochs.
+5.2 Ablation: Number of Prototypes
+Figure 6: Ablation: Test Accuracy on CUB with different numbers of prototypes per class.
+We conduct an ablation study over the number of prototypes per class and present the test accuracy scores in Figure 6.
+Notably, the predictive performance of Euclidean and Gaussian prototypes degrades when more prototypes are added.
+The prototype formulations with a hyperspherical structure, Cosine and HyperPG prototypes converge to similar performance levels regardless of the number of prototypes per class.
+5.3 Ablation: Number of Dimensions
+We conduct an ablation study over the number of prototype dimensions and present the test accuracy scores in Figure 7.
+Hyperspherical prototype formulations such as Cosine and HyperPG are robust regarding the choice of prototype dimensions, although HyperPG performs slightly better with increased dimensionality.
+The Gaussian prototypes, like HyperPG prototypes, can adapt their variance $\sigma^{2}$ to deal with the increase in dimensionality.
+The performance of Euclidean prototypes suffers most with changes in dimensionality.
+Figure 7: Ablation: Test Accuracy on CUB with different numbers of prototype dimensions.
+5.4 Ablation: Additional Datasets
+We perform additional experiments on two different datasets.
+Stanford Cars (Cars) with 196 car models (Krause et al., 2013) and Oxford Flowers (Flowers) with 102 different species (Nilsback & Zisserman, 2008).
+The relative differences in performance seem to hold across the different datasets, with Euclidean prototypes performing the worst.
+Interestingly, Gaussian prototypes are able to match the performance of the hyperspherical formulations more closely.
+Table 3: Top-1 Test Accuracy with Resnet50 Backbone.
+Cars Flowers
+Baseline 85.2 90.0
+Euclidean (ProtoPNet) 36.2 15.0
+Gaussian 73.0 74.8
+Cosine 79.6 85.5
+HyperPG (Ours) 79.0 87.8
+6 Interpretability Analysis
+We perform a qualitative analysis of the different prototype formulations with ResNet50 backbone (see Figure 8).
+One common visualization technique originally proposed by ProtoPNet is a a direct overlay of the prototype activation map (PAM).
+However, this visualization technique has several shortcomings, which we discuss in more detail in Appendix A.
+In order to offer an alternative to the PAM visualization, we also present GradCAM (Selvaraju et al., 2017) overlays.
+Both Euclidian and HyperPG prototypes do not actually learn class-specific prototypes, in contrast to Gaussian or Cosine prototypes.
+Interestingly, all prototype formulations seem to activate in similar regions in the test image, but differ with the known training examples.
+HyperPG could already leverage class-shared prototypes, which would explain its robustness to changes in the number of prototypes per class.
+There seems to be no qualitative difference between the visualizations produced with the various prototype formulations using the CNN based backbones.
+As expected, the ViT backbone (see Figure 9) suffers from the “distraction problem” as observed by ProtoPFormer (Xue et al., 2024).
+We present additional examples in Appendix A.
+Refer to caption Refer to caption
+(a) Euclidean (R50)
+Refer to caption Refer to caption
+(b) Gaussian (R50)
+Refer to caption Refer to caption
+(c) Cosine (R50)
+Refer to caption Refer to caption
+(d) HyperPG (R50)
+Figure 8: Left: Test Image. Right: 3 closest Training Patches to the Prototype. Top Row: PAM visualization. Bottom Row: GradCAM visualization. R50: Resnet50 Backbone. ViT: ViT-B16 Backbone.
+Figure 9: Cosine Protoypes with ViT. The overlayed heatmap show the distraction problem caused by the attention mechanism.
 7 Conclusion
-This work introduces HyperPg, a new prototype representation learning a probability distribution on the surface of a hypersphere in latent space.
-HyperPg prototypes adapt to the variance of clusters in latent space and improve training time and accuracy compared to other prototype formulations.
-HyperPgNet leverages HyperPg to learn human-defined concept prototypes instead of black-box optimized prototypes.
-The combination of probabilistic prototypes on the hypersphere and concept-aligned prototypes allows HyperPgNet to outperform other prototype learning approaches with regards to accuracy and interpretability.
-One limitation HyperPgNet faces are slightly higher computational requirements due to the inclusion of concept annotations and $L_{\mathrm{RRC}}$ during training.
-However, this is offset by faster convergence.
-Coupled with the increased transparency and interpretability of the model, this makes HyperPgNet a strong contender for scenarios with higher requirements for model trust and safety, like medical applications or human-robot interaction.
-Reproducibility Statement
-For reproducing the experiments, please refer to the following sections: Sec.
-4 describes the required loss functions and the overall network architecture.
-Subsection 3.3 details the activation function of the newly introduced HyperPg prototypes.
-Appendix B provides detailed information on the experiment implementations, including training hyperparameters, processing of the datasets and model implementation details.
-Furthermore, we will publish the code on GitHub after the double-blind review period.
-Appendix A Adapting HyperPg to other Probability Distributions
-Subsection 3.3 defines HyperPg prototypes ${\bm{p}}^{H}=(\bm{{\alpha}},\mu,\sigma)$ as a Gaussian Distribution with mean $\mu$ and std $\sigma$ of cosine similarities around an anchor vector $\alpha$.
+This work provides comprehensive overview of different prototype formulations, with both probabilistic and point-based measurements, in an Euclidean or hyperspherical latent space.
+We introduce HyperPG, a new prototype representation based on Gaussian distributions on the surface of a hypersphere.
+Our experiments showcase the differences in performance and robustness to hyperparameters across different datasets and backbone architectures for these prototype formulations.
+Our results demonstrate the performance advantage of hypersherical prototypes and probabilistic prototypes in contrast to the commonly used Euclidean formulation, with no loss to the inherent interpretability.
+Future work could explore further refinements to the probabilistic prototype formulations, including adaptive mechanisms for prototype selection or the integration of additional probabilistic models such as Mixture Models or Bayesian approaches like MGProto (Wang et al., 2024b) to further enhance interpretability and performance.
+Impact Statement
+This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
+Appendix A Extended Interpretability Analysis
+A.1 Potential Problems with Prototype Activation Maps
+One common visualization technique originally proposed by ProtoPNet is a a direct overlay of the prototype activation map (PAM).
+In this manner, the latent similarity map for each prototype is scaled back up to the original input resolution and overlayed over the image.
+A box is drawn around the 95th quantile to mark the highest activated image region.
+This technique has some obvious flaws: The latent similarity map is usually very low resolution, e.g., for ResNet50 $7\times 7$, unless a segmentation model such as DeeplabV3 (Chen et al., 2017) is used.
+Any perceived gradient in the overlay is likely an artifact from the upscaling operation.
+Additionally, there exist robustness problems with this visualization (Sacha et al., 2024).
+By modifying parts of the background or adding noise to the image, the activation map can change, most likely due to the spatial pooling in the convolution pyramids.
+Appendix B Adapting HyperPG to other Probability Distributions
+We define HyperPG prototypes ${\bm{p}}^{H}=(\bm{{\alpha}},\mu,\sigma)$ as a Gaussian Distribution with mean $\mu$ and variance $\sigma^{2}$ of cosine similarities around an anchor vector $\alpha$.
 This idea of learning a distribution of cosine similarity values around an anchor $\bm{{\alpha}}$ can be adapted to other distributions.
@@
 As early experiments on the CUB-200-2011 dataset showed no significant difference in performance, these sections are relegated to the appendix.
-A.1 Cauchy Distribution
+B.1 Cauchy Distribution
 One theoretical disadvantage of the Gaussian distribution is the fast approach to zero, which is why a distribution with heavier tails such as the Cauchy distribution might be desirable.
@@
 with median $x_{0}$ and average absolute deviation $\gamma$. The HyperPG prototypes with Cauchy are defined as accordingly as ${\bm{p}}^{\mathrm{Cauchy}}=(\bm{{\alpha}},x_{0},\gamma)$.
-Fig.
-9 illustrates the PDF of the Gaussian and Cauchy distributions with $\mu=x_{0}=1$ and $\sigma=\gamma=0.2$, i.e., the main probability mass is aligned with the anchor $\bm{{\alpha}}$.
+Figure 18 illustrates the PDF of the Gaussian and Cauchy distributions with $\mu=x_{0}=1$ and $\sigma=\gamma=0.2$, i.e., the main probability mass is aligned with the anchor $\bm{{\alpha}}$.
 The Gaussian distributions PDF quickly approaches zero and stays near constant.
@@
 Figure 18: PDF for the Gaussian and Cauchy distribution of cosine similarity values. The Cauchy distribution has heavier tails, avoiding vanishing gradients issues.
-A.2 Truncated Distributions
+B.2 Truncated Distributions
 The cosine similarity is defined only on the interval $[-1,1]$.
@@
 $\mathcal{T}_{\mathrm{Cauchy}}(x,x_{0},\gamma,a,b)=\frac{1}{\gamma}\left(1+\left(\frac{x-x_{0}}{\sigma}\right)^{2}\right)^{-1}\left(\arctan\left(\frac{b-x_{0}}{\gamma}\right)-\arctan\left(\frac{a-x_{0}}{\gamma}\right)\right)^{-1}.$ (14)
-Fig.
-10 shows the test accuracy of three HyperPgNet models with Gaussian, truncated Gaussian and truncated Cauchy distribution on the CUB-200-2011 dataset.
+Figure 19 shows the test accuracy of three HyperPGNet models with Gaussian, truncated Gaussian and truncated Cauchy distribution on the CUB-200-2011 dataset.
 While difference in test performance and learning speed were minimal on the CUB-200-2011 dataset, further exploration is necessary, as other experiments showed that the concept-alignment on CUB-200-2011 dominates the learning process, lessening the influence of the prototypes.
-Figure 10: CUB-200-2011 Test Accuracy of HyperPgNet with different probability distributions. The difference in performance is only marginal.
-A.3 von Mises-Fisher Distribution
-The von Mises-Fisher distribution (vMF) is the analogue of the Gaussian distribution on the surface of a hypersphere Hillen et al.
-(2017).
+Figure 19: CUB-200-2011 Test Accuracy with HyperPG prototypes using different PDFs.
+B.3 von Mises-Fisher Distribution
+The von Mises-Fisher distribution (vMF) is the analogue of the Gaussian distribution on the surface of a hypersphere (Hillen et al., 2017).
 The density function $f_{d}$ of the vMF distribution for a $D$-dimensional unit-length vector ${\bm{v}}$ is defined as
@@
 with mean vector $\bm{{\alpha}}$, scalar concentration parameter $\kappa$ and normalization constant $C_{d}(\kappa)$.
-The normalization constant $C_{d}(\kappa)$ is a complex function and difficult to compute for higher dimensions, which is why, for example, Tensorflow [^1] Tensorflow API Documentation - Accessed 2024-09-20 only supports the vMF distribution for $D\leq 5$.
+The normalization constant $C_{d}(\kappa)$ is a complex function and difficult to compute for higher dimensions, which is why, for example, Tensorflow [^1] Tensorflow API Documentation - Accessed 2025-01-10 only supports the vMF distribution for $D\leq 5$.
 However, the vMF distribution is a viable similarity measure when using the unnormalized density function with $C_{d}(\kappa)=1$.
@@
 $\sim\exp\left(\kappa{\bm{v}}^{\top}\bm{{\alpha}}\right).$ (24)
-Eq.
-23 also shows the relationship to the HyperPg similarity with an untruncated Gaussian distribution and prototype mean activation $\mu=1$.
-Fig.
-11 presents a simulation of the vMF distribution on a 3D sphere.
-While both the vMF distribution and HyperPg activation can produce a spherical, gaussian-like activation pattern on the surface of a hypersphere, the vMF distribution cannot produce the ring pattern shown in Fig.
-2.
+Equation 21 also shows the relationship to the HyperPG similarity with an untruncated Gaussian distribution and prototype mean activation $\mu=1$.
+Figure 20 presents a simulation of the vMF distribution on a 3D sphere.
+While both the vMF distribution and HyperPG activation can produce a spherical, gaussian-like activation pattern on the surface of a hypersphere, the vMF distribution cannot produce the ring pattern shown in Figure 3.
 The ring pattern produced by adapting HyperPG’s mean similarity $\mu$ could be approximated by a mixture of vMF distributions.
 Figure 20: Changing the concentration parameter $\kappa$ is akin to changing HyperPG’s std $\sigma$.
-A.4 Fisher-Bingham Distribution
+B.4 Fisher-Bingham Distribution
 As the vMF distribution is the equivalent of an isotropic Gaussian distribution on the surface of a hypersphere, the Fisher-Bingham (FB) distribution is the equivalent of a Gaussian with full covariance matrix.
@@
 For a $D$ dimensional space, the FB distribution is by a $D\times D$ matrix ${\bm{A}}$ of orthogonal vectors $(\bm{{\alpha}}_{1},\bm{{\alpha}}_{2},\ldots,\bm{{\alpha}}_{D})$, concentration parameter $\kappa$ and ellipticity factors $[\beta]_{2:D}$ where $\sum_{j=2}^{D}\beta_{j}=1$ and $0\leq 2|\beta_{j}|<\kappa$.
-The FB unnormalzied PDF is defined as
+The FB’s unnormalized PDF is defined as
 $b({\bm{v}}|{\bm{A}},\kappa,\beta)=\exp\left(\kappa\bm{{\alpha}}_{1}^{\top}{\bm{v}}+\sum_{j=2}^{D}\beta_{j}\left(\bm{{\alpha}}_{j}^{\top}{\bm{v}}\right)^{2}\right).$ (25)
-The FB distribution’s main advantage is the elliptic form of the distribution on the surface of the hypersphere, offering higher adaptability than the other formulations (see Fig.
-12.
+The FB distribution’s main advantage is the elliptic form of the distribution on the surface of the hypersphere, offering higher adaptability than the other formulations (see Figure 21.
 However, the parameter count and constraints are higher.
 Figure 21: Illustration of the Fisher-Bingham Distribution in $D=3$
-A.5 Mixture Models
+B.5 Mixture Models
 HyperPG’s probabilistic nature lends itself to a mixture formulation.
@@
 First experiments with mixture of HyperPG prototypes did not show any improvement over the standard formulation. However, this might change with other datasets.
-Appendix B Implementation Details
-B.1 Data Preprocessing
+Appendix C Scaled Dot-Product Prototype
+This section does not reiterate the full attention mechanism.
+We recommend to the interested reader to refer to the original work (Vaswani et al., 2017).
+Rather, we will take a high-level view of the main equation and explain how we apply it to a prototype formulation.
+Vaswani et al. (2017) define the scaled dot-product attention for some input matrices query ${\bm{Q}}$, key ${\bm{K}}$ and value ${\bm{V}}$ as
+$\mathrm{Attention}({\bm{Q}},{\bm{K}},{\bm{V}})=\mathrm{softmax}\left(\frac{{\bm{Q}}{\bm{K}}^{\top}}{\sqrt{d_{k}}}\right){\bm{V}}$ (28)
+with $d_{k}$ being the number of dimensions for the key matrix ${\bm{K}}$.
+We can use the scaled dot-product to define a new prototype similarity function
+$s_{\mathrm{sdot}}({\bm{z}}|{\bm{p}})=\frac{{\bm{z}}^{\top}{\bm{p}}}{\sqrt{D}}$ (29)
+for a latent patch ${\bm{z}}$ and prototype vector ${\bm{p}}$ with dimensionality $D$.
+In contrast to the cosine similarity, the vectors are not normalized to unit length, which preserves more information.
+However, dividing by the square root of the number of dimensions $D$ acts as a heuristic regarding vector length.
+Intuitively we can think about it as instead of a hard projection onto the surface of the unit-radius hypersphere, all points are moved towards the surface.
+Figure 22 illustrates this scaling.
+In this manner, the angle between vectors becomes a meaningful measurement, while also retaining some information encoded by the vector length.
+Figure 22: Illustration of Cosine Similarity and Scaled Dot-Product
+Appendix D Implementation Details
+D.1 Data Preprocessing
 In contrast to prior work (e.g.
 Chen et al., 2019; Rymarczyk et al., 2020; Ukai et al., 2023) the experiment used an online augmentation process, resulting in 30 training images per class and epoch.
-The input images were first resized to a resolution of $224\times 224$ without cropping to bounding box annotations.
-The augmentations consisted of RandomPerspective, RandomHorizontalFlip and RandomAffine.
-B.2 Hyperparameters
-The prototypical part networks ProtoPNet and HyperPgNet use a convolutional neck after the feature encoder and work on a latent feature map of size $7\times 7\times 256$.
-The models were trained with a minibatch size of 96 images using AdamW optimizer with learning rate 1e-4 and weight decay 1e-4.
-B.3 Compute Resources
-All experiments were performed on a workstation with a single NVIDIA RTX 3090 GPU (24 GB VRAM) per model.
-B.4 Model Implementation
-The models were implement with different feature encoding backbones.
-ConvNeXt-tiny (Liu et al., 2022) and ResNet50 (He et al., 2016) with pretrained weights on ImageNet provide CNN based backbones.
-As a transformer based backbone MiT-B4 (Xie et al., 2021) as provided by HuggingFace is used.
-MiT-B4 is pretrained on ImageNet and finetuned on ADE20k, Cityscapes and COCO-stuff.
-The models are implemented as follows:
-  • The black box baseline model uses the pretrained feature encoder as backbone followed by the classification head: a single linear output layer and softmax activation.
-  • Concept Bottleneck Models (Koh et al., 2020) integrate the concept bottleneck, a linear layer with one neuron per concept, between the backbone and classification head.
-  • ProtoPNet (Chen et al., 2019) use a $L_{2}$ based prototype layer between backbone and classification head.
-  • ProtoPNet + HyperPg use a HyperPg based prototype module between backbone and classification head.
-  • HyperPgNet $-{\mathcal{L}}_{\mathrm{RRC}}$ uses a HyperPg prototype module between backbone and classification head.
-The prototypes are trained with concept alignment, but without the ${\mathcal{L}}_{\mathrm{RRC}}$ loss.
-  • HyperPgNet $+{\mathcal{L}}_{\mathrm{RRC}}$ uses a HyperPg prototype module between backbone and classification head.
-The prototypes are trained with concept alignment and including the ${\mathcal{L}}_{\mathrm{RRC}}$ loss.
-Appendix C Extended Interpretability Analysis
-C.1 Latent Space Structure
-Prototype learning is based on learning structures in the latent space.
-HyperPg specifically learns Gaussian distributions on the surface of a hypersphere in a high-dimensional latent space.
-Dimensionality reduction techniques like UMAP (McInnes et al., 2018) aim to preserve global and local structures from a high dimensional space when projecting into a low dimensional one.
-UMAP supports multiple distance metrics, including hyperspheric manifold distances, thus retaining some of the high dimensional structure when projecting onto a 3D sphere.
-Fig.
-13 illustrates UMAP projections of HyperPg concept-aligned prototypes trained on Stanford Cars.
-This visualization indicates that HyperPgNet is able to disentangle the different concepts in latent space, as the projection shows no overlap of the different clusters.
-When trained with $L_{\mathrm{RRC}}$, the prototypes are packed closer together.
-This could indicate, that the chosen hyper parameter of 20 prototypes per concept is higher than required as not all concept embeddings have the same diversity in the latent space.
-Fig.
-14 illustrates one UMAP projection of HyperPg concept-aligned prototypes trained on CUB-200-2011.
-In comparison to Stanford Cars, the latent space appears less structured.
-This could explain the higher difficulty associated with this dataset.
-(a) HyperPgNet - $L_{\mathrm{RRC}}$
-(b) HyperPgNet + $L_{\mathrm{RRC}}$
-Figure 13: UMAP projection of HyperPg concept prototypes learned on Stanford Cars. Each dot represents one prototype with the color indicating a concept.
-Figure 14: UMAP Projection of HyperPg concept prototypes learned on CUB-200-2011.
+The input images were cropped to the bounding box annotations, then resized to a resolution of $224\times 224$.
+The augmentations consisted of RandomPerspective, RandomHorizontalFlip and ColorJitter.
+D.2 Hyperparameters
+The prototypical part networks ProtoPNet and HyperPGNet use a convolutional neck after the feature encoder and work on a latent feature map of size $7\times 7\times 128$.
+The models were trained with a minibatch size of 48 images using AdamW optimizer with learning rate 1e-4 and weight decay 1e-4.
+D.3 Compute Resources
+All experiments were performed on a desktop workstation with a single NVIDIA RTX 3070Ti GPU (8 GB VRAM) per model.
+On CUB-200-2011 the training duration was roughly 40 seconds per epoch, regardless of the used prototype formulation.
+D.4 Model Implementation
+The experiments used various feature encoding backbones such as ResNet50 (He et al., 2016), DenseNet121 (Huang et al., 2017) and ViT-B16 (Dosovitskiy, 2020) with pretrained weights on ImageNet.
+All models use a single linear output layer as a classification head.
+The prototypical networks follow the architecture proposed by ProtoPNet (Chen et al., 2019): After the pretrained feature encoder, a bottleneck consisting of a Convolution Layer, ReLU activation, Convolution Layer and Sigmoid activation reduces the dimensions of the feature map to the prototype space.
+For example, the ResNet50 encoder produces a feature map of $H\times W\times D$ shape $7\times 7\times 2048$.
+The bottleneck produces a dimension reduction to shape $7\times 7\times 128$.
+The prototype modules produce the prototype similarity scores, which are passed through a global max pooling, before being passed to the classification head.
+Appendix E Parameter Distribution of HyperPG
+To evaluate, if HyperPG actually uses the probabilstic components, we look at the learned mean $\mu$ and standard deviation $\sigma$ values after training on CUB200-2011.
+Figure 23 plots the mean and sigma values.
+The red star marks the theoretical solution if it would learn something similar to the pure Cosine prototypes.
+The pattern demonstrates, that HyperPG actually learns different parameters for the Gaussian components of each prototype.
+Interestingly, two modes at $\sigma=0.5$ and $\sigma=0.4$ appear with a distinct line shape, which warrant future exploration.
+Figure 23: HyperPG learned mean and sigma values.

### v4

@@
 Abstract
-Prototypical part networks offer interpretable alternatives to black-box deep learning models.
-However, many of these networks rely on Euclidean prototypes, which may limit their flexibility.
-This work provides a comprehensive overview of various prototype formulations.
-Experiments conducted on the CUB-200-2011, Stanford Cars, and Oxford Flowers datasets demonstrate the effectiveness and versatility of these different formulations.
+Prototypical part networks offer interpretable alternatives to black-box deep learning models by learning visual prototypes for classification.
+This work provides a comprehensive analysis of prototype formulations, comparing point-based and probabilistic approaches in both Euclidean and hyperspherical latent spaces.
+We introduce HyperPG, a probabilistic prototype representation using Gaussian distributions on hyperspheres.
+Experiments on CUB-200-2011, Stanford Cars, and Oxford Flowers datasets show that hyperspherical prototypes outperform standard Euclidean formulations.
+Critically, hyperspherical prototypes maintain competitive performance under simplified training schemes, while Euclidean prototypes require extensive hyperparameter tuning.
 XAI, Prototype Learning, Image Classification, Deep Learning
@@
 However, the decision-making processes of these models lack transparency and interpretability, making deployment in safety-critical areas challenging.
-Explainable Artificial Intelligence (XAI) seeks to develop interpretability methods to open the black-box reasoning processes of such deep models and increase trust in their decisions.
-XAI methods can be broadly divided into two categories: First, Post-Hoc Methods like LIME (Ribeiro et al., 2016), SHAP (Lundberg & Lee, 2017) or GradCAM (Selvaraju et al., 2017) offer explanations for predictions without requiring retraining.
-While applicable in many scenarios, post-hoc methods may not actually align with the models’ decision-making processes, potentially leading to interpretations that are not entirely faithful (Rudin, 2019).
-Second, inherently interpretable methods provide built-in, case-based reasoning processes.
-For instance, small decision trees are inherently interpretable because their reasoning can be easily understood as a series of if-else statements (Molnar, 2020).
-However, they are constrained in their representational power.
+Explainable Artificial Intelligence (XAI) seeks to develop interpretability methods to open the black-box reasoning processes of models and increase trust in their decisions.
+XAI methods can be broadly divided into two categories: First, post-hoc methods like LIME (Ribeiro et al., 2016), SHAP (Lundberg and Lee, 2017) or GradCAM (Selvaraju et al., 2017), which explain model predictions without retraining.
+While broadly applicable, post-hoc methods may not align with the models’ actual decision-making processes (Rudin, 2019).
+Second, inherently interpretable methods provide built-in, case-based reasoning.
+For instance, small decision trees are inherently interpretable through their if-else structure (Molnar, 2020), but are limited in representational power.
 Figure 1: Different Prototype Formulations. HyperPG is a novel formulation for probabilistic prototypes on a Hypersphere.
-Deep Prototype Learning Architectures such as ProtoPNet (Chen et al., 2019) and its derivatives (e.g., Rymarczyk et al., 2020; Donnelly et al., 2021; Sacha et al., 2023) integrate inherent interpretability into deep learning models through a prototype layer.
-Each neuron in this layer represents a prototype, storing a latent feature vector.
-The model’s predictions are based on the distances between sample features and prototype parameters, for example, by computing the $L_{2}$-distance.
-This makes prototype learning essentially a clustering task in latent space (Zhou et al., 2022).
-Different prototype formulations, such as prototypes using the cosine similarity, induce a hyperspherical structure to the latent space with performance advantages to classification (Mettes et al., 2019).
-Probabilistic formulations such as Gaussian prototypes enable further downstream tasks relying on probability values such as Bayesian approaches.
-With the increased popularity of prototypical part models, there has been a corresponding rise in the number of architectural choices.
-However, to the best of our knowledge, there is no comprehensive overview of the design choices and prototype formulations for learning interpretable prototypes and their effect on predictive performance.
-Table 1: Overview of 14 related models and their prototype learning configurations. 8/14 models use Euclidean prototypes.
+Deep prototype learning architectures such as ProtoPNet (Chen et al., 2019) and its derivatives (Rymarczyk et al., 2020; Donnelly et al., 2021; Sacha et al., 2023) integrate inherent interpretability into deep learning through a prototype layer.
+Each prototype neuron stores a latent feature vector, with predictions based on distances between sample features and prototypes.
+This makes prototype learning a clustering task in latent space (Zhou et al., 2022).
+Different formulations induce different latent structures: cosine similarity induces hyperspherical structure with classification advantages (Mettes et al., 2019), while probabilistic formulations like Gaussian prototypes enable Bayesian downstream tasks.
+Table 1: Overview of 15 related models and their prototype learning configurations.
 Model Similarity Shape Assignment Clf. Head
-Prototype Decoder Euclidean Entire Image Class Exclusive FCL Li et al. (2018)
-ProtoPNet Euclidean Patch Class Exclusive FCL Chen et al. (2019)
-Def. ProtoPNet Cosine Spatial Arrangement Class Exclusive FCL Donnelly et al. (2021)
-ProtoPShare Euclidean Patch Merged after Training FCL Rymarczyk et al. (2020)
-ProtoPool Euclidean Patch Pooling Shared FCL Rymarczyk et al. (2022)
-TesNet Grassman Patch Class Exclusive FCL Wang et al. (2021)
-ProtoTree Euclidean Patch Shared Decision Tree Nauta et al. (2021)
-ProtoKNN Cosine Patch Class Exclusive KNN Clf Ukai et al. (2023)
-PIPNet Cosine Patch Shared FCL Nauta et al. (2023)
-LucidPPN Euclidean Separate Color & Texture Class Exclusive Branch Aggregation Pach et al. (2024)
-ProtoSeg Cosine Patch Class Exclusive FCL Zhou et al. (2022)
-ProtoGMM Gaussian Patch Class Exclusive FCL Moradinasab et al. (2024)
-MGProto Gaussian Patch Class Exclusive Bayesian Likelihood Wang et al. (2024b)
-ProtoPFormer Euclidean Transformer Token Class Exclusive Branch Aggregration Xue et al. (2024)
-This paper explores different prototype formulations. Its contributions are as follows:
-Prototype Formulation Overview: A comprehensive overview of different, learnable prototype definitions, including both point based and probabilistic prototypes as shown in Figure 1.
-HyperPG: A new probabilistic prototype representation with learned parameters anchor $\alpha$, mean $\mu$ and variance $\sigma^{2}$.
-This representation models a Gaussian distribution over cosine similarities, thereby projecting a Gaussian distribution on the surface of a hypersphere.
-HyperPG’s similarity score is based on the Gaussian’s probability density function and adapts its size through a learned standard deviation.
-Benchmarking Protoypes: Extensive image classification experiments based on prototypical part network architectures like ProtoPNet on the CUB-200-2011 (Wah et al., 2011), Stanford Cars (Krause et al., 2013) and Oxford Flowers (Nilsback & Zisserman, 2008) datasets.
+Prototype Decoder Euclidean Entire Image Class Exclusive FCL (Li et al., 2018)
+ProtoPNet Euclidean Patch Class Exclusive FCL (Chen et al., 2019)
+Def. ProtoPNet Cosine Spatial Arrangement Class Exclusive FCL (Donnelly et al., 2021)
+ProtoPShare Euclidean Patch Merged after Training FCL (Rymarczyk et al., 2020)
+ProtoPool Euclidean Patch Pooling Shared FCL (Rymarczyk et al., 2022)
+TesNet Grassman Patch Class Exclusive FCL (Wang et al., 2021)
+ProtoTree Euclidean Patch Shared Decision Tree (Nauta et al., 2021)
+ProtoKNN Cosine Patch Class Exclusive KNN Clf (Ukai et al., 2023)
+PIPNet Cosine Patch Shared FCL (Nauta et al., 2023)
+ST-ProtoPNet Cosine Patch Class Exclusive Branch Aggregation (Wang et al., 2023)
+LucidPPN Euclidean Separate Color & Texture Class Exclusive Branch Aggregation (Pach et al., 2024)
+ProtoSeg Cosine Patch Class Exclusive FCL (Zhou et al., 2022)
+ProtoGMM Gaussian Patch Class Exclusive FCL (Moradinasab et al., 2024)
+MGProto Gaussian Patch Class Exclusive Bayesian Likelihood (Wang et al., 2024b)
+ProtoPFormer Euclidean Transformer Token Class Exclusive Branch Aggregration (Xue et al., 2024)
+The increased popularity of prototypical part models increases architectural choices, yet there is no comprehensive overview of prototype formulations and their effect on predictive performance.
+This paper explores different prototype formulations with the following contributions:
+  • Prototype Formulation Overview: A comprehensive overview of different learnable prototype definitions, including both point-based and probabilistic prototypes as shown in Figure 1.
+  • HyperPG: A probabilistic prototype representation modeling Gaussian distributions over cosine similarities with learned anchor $\alpha$, mean $\mu$, and variance $\sigma^{2}$, projecting Gaussians onto hypersphere surfaces.
+  • Benchmarking and Robustness Analysis: Image classification experiments on CUB-200-2011 (Wah et al., 2011), Stanford Cars (Krause et al., 2013), and Oxford Flowers (Nilsback and Zisserman, 2008) datasets, evaluating both full optimization and simplified training procedures to reveal differences in hyperparameter sensitivity.
+Our code is publicly available on Github [^1] https://github.com/LiXiling/prob-proto.
 2 Related Work
@@
 TesNet (Wang et al., 2021) proposes to compute the prototype similarity on the Grassman Manifold.
+Figure 2: Illustration of similarity computation between prototype $\bm{p}$} and latent vector $\bm{z}$} for different formulations.
+Euclidean: $L_{2}$ distance in latent space.
+Hyperspherical: cosine similarity of normalized vectors (angle on hypersphere).
+Gaussian: PDF of Gaussian distribution in Euclidean space.
+HyperPG: PDF of Gaussian distribution over cosine similarities (Gaussian on hypersphere surface).
 All these approaches share the use of a linear classification layer.
 However, ProtoTree (Nauta et al., 2021) builds a Decision Tree on the learned Euclidean prototypes and ProtoKNN (Ukai et al., 2023) proposes to use a $k$-nearest-neighbor classifier with cosine prototype similarities.
+ST-ProtoPNet (Wang et al., 2023) learn the prototypes close to the decision boundary, similar to support vectors.
 Recent work proposes to learn more interesting or robust features for the prototypes from the image encoders.
@@
 LucidPPN (Pach et al., 2024) propose a hybrid architecture with texture only grayscale prototypes and low resolution color only prototypes.
-Zhou et al.
-(2022) propose to view prototype learning as a clustering task in the network’s latent space for image segmentation.
-In this setting, the clustering can be done outside of the network’s backpropagation path with an external clustering algorithm, so-called non-learnable prototypes.
+“Rethinking Semantic Segmentation” (Zhou et al., 2022) view prototype learning as a clustering task in the network’s latent space.
+The clustering is done outside of the network’s backpropagation path with an external clustering algorithm, so-called non-learnable prototypes.
 Later work such as ProtoGMM (Moradinasab et al., 2024) and MGProto (Wang et al., 2024b) build on this idea and model the prototypes as Gaussian distributions.
-ProtoPFormer (Xue et al., 2024) adapts Euclidean prototypical part networks to Transformer based architectures by using a VisionTransformer (ViT, Dosovitskiy, 2020) backbone.
-To avoid the distraction problem due to the global attention mechanism of ViT architectures, ProtoPFormer uses a hybrid approach with global and local prototypes.
+ProtoPFormer (Xue et al., 2024) adapts Euclidean ProtoPNets to Transformers through a ViT (Dosovitskiy, 2020) backbone.
+To avoid the distraction problem due to the global attention mechanism of ViTs, ProtoPFormer uses a hybrid approach with global and local prototypes.
 The local prototypes are encouraged to focus on different parts of the image by modeling the spatial prototype activations as a 2D Gaussian.
-Table 1 presents a systematic overview of the discussed works.
-The cosine similarity has been shown to perform well in classification tasks (Mettes et al., 2019), but only few prototype learning works have used them so far (Donnelly et al., 2021; Zhou et al., 2022; Ukai et al., 2023).
-There is also a lack for probabilistic prototypes in a hyperspherical space.
-We introduce HyperPG, which combines the predictive performance gains from the cosine similarity with the probabilistic nature of Gaussian distributions.
+Table 1 presents a systematic overview, where the cosine similarity performs well in classification tasks (Mettes et al., 2019), but only few prototype learning works have used them so far (Donnelly et al., 2021; Zhou et al., 2022; Ukai et al., 2023).
+There is a lack for probabilistic prototypes in a hyperspherical space.
+We introduce HyperPG, that combines the performance gains from the cosine similarity with the probabilistic nature of Gaussian distributions.
 3 Overview: Prototype Formulations
-Figure 2: Illustration on how the different prototype formulations compute the similarity between a prototype ${\bm{p}}$ and latent vector ${\bm{z}}$.
-Euclidean prototypes compute the $L_{2}$ distance between two points in latent space.
-Hyperspherical prototypes use the cosine similarity of normalized vectors, which corresponds to the angle between two points on a hypersphere.
-Gaussian prototypes model a Gaussian distribution in Euclidean space and compute the probability density function (PDF).
-HyperPG prototypes learn a Gaussian distribution of cosine similarities, thereby projecting a Gaussian distribution onto the surface of a hypersphere.
 Prototype Learning is an inherently interpretable machine learning method.
@@
 Notable exceptions include Deformable ProtoPNet (Donnelly et al., 2021), where each prototype has is a $\rho=3\times 3$ arrangement of smaller patches, and MCPNet (Wang et al., 2024a), which learns concept prototypes across the latent features.
-The prediction is computed by comparing each prototype ${\bm{p}}$ to the latent feature map ${\bm{z}}$.
-For simplicity’s sake lets assume the spatial dimensions $\rho=\zeta=1$.
+The prediction is done by comparing each prototype $\bm{p}$} to the latent feature map $\bm{z}$}.
+For simplicity, lets assume the spatial dimensions $\rho=\zeta=1$.
 The following equations can be adapted for higher spatial dimensions by summing over the height and width $\sum_{\rho_{w}}\sum_{\rho_{h}}$ for each patch of the latent map.
@@
 By normalizing $D$ dimensional vectors to unit length, they are projected onto the surface of a $D$ dimensional hypershere.
-The cosine similarity is defined on the interval $[-1,1]$ and measures: $1$ for two vectors pointing in the same direction, 00 for orthogonal vectors, and $-1$ for vectors pointing in opposite directions.
+The cosine similarity is defined on the interval [1,1][-1,1] and measures: $1$ for two vectors pointing in the same direction, 0 for orthogonal vectors, and 1-1 for vectors pointing in opposite directions.
 Without the normalization to unit length, we can define a scaled dot-product prototype inspired by the Attention mechanism (Vaswani et al., 2017).
 However, this similarity measure is not bounded to any interval, as the vector length is proportionally preserved.
-For more details see Appendix C.
+For more details see Appendix E.
 Like the $L_{2}$ similarity, the cosine similarity is a point-based measure comparing two vectors directly.
@@
 $=\frac{1}{(2\pi)^{\frac{D}{2}}|\bm{\mathrm{\Sigma}}|^{\frac{1}{2}}}$ $\exp\left(-\frac{1}{2}({\bm{z}}-\bm{{\mu}})^{\top}\bm{\mathrm{\Sigma}}^{-1}({\bm{z}}-\bm{{\mu}})\right).$ (4)
-It is worth noting that the term inside the exponent is based on the $L_{2}$ distance between the latent vector ${\bm{z}}$ and the mean $\bm{{\mu}}$.
-This formulation as a PDF has several advantages: A) The similarity can be interpreted as the likelihood of being sampled from the Gaussian prototypes, which is more meaningful than a distance metric in a high-dimensional latent space.
+The term inside the exponent is based on the $L_{2}$ distance between the latent vector $\bm{z}$} and the mean $\bm{{\mu}}$.
+The PDF formulation has several advantages: A) The similarity can be interpreted as the likelihood of being sampled from the Gaussian prototypes, which is more meaningful than a distance metric in a high-dimensional latent space.
 B) Prototypes can adapt their shape using a full covariance matrix, allowing different variances along various feature dimensions, offering more flexibility in shaping the latent space.
-However, learning a full covariance matrix increases computational requirements, especially with EM clustering.
+However, learning a full covariance matrix increases computational requirements.
 (a) $\mu=1$
@@
 Figure 3: HyperPG prototypes learn a distribution of cosine similarities.
-They are parameterized by a learnable anchor vector $\bm{{\alpha}}$, scalar mean $\mu$ and scalar variance $\sigma^{2}$.
-HyperPG projects a Gaussian distribution of cosine similarities on the surface of a hypersphere, resulting in ring shaped activation patterns around the anchor vector.
+They use a learnable anchor vector $\bm{{\alpha}}$, scalar mean $\mu$ and variance $\sigma^{2}$.
+They project a Gaussian distribution of cosine similarities on the surface of a hypersphere, resulting in ring shaped activation patterns around the anchor vector.
 Figure 4: Prototype Learning Architecture.
@@
 HyperPG prototypes learn a $1D$ Gaussian distribution over the cosine similarities to the anchor vector $\bm{{\alpha}}$.
-Because the cosine similarity is bounded to $[-1,1]$, HyperPG’s similarity measure is defined as the PDF of the truncated Gaussian distribution within these bounds.
+Because the cosine similarity is bounded to [1,1][-1,1], HyperPG’s similarity measure is defined as the PDF of the truncated Gaussian distribution within these bounds.
 Let $\mathcal{G}(x,\mu,\sigma)$ be the cumulative Gaussian distribution function.
@@
 For $\mu=1$, the distribution centers around the anchor, as the cosine similarity is $1$ if two vectors point in the same direction.
-This corresponds to the von Mises-Fisher (vMF) distribution (Appendix B).
-For $\mu=-1$, the distribution is the on opposite side of the hypersphere, as the cosine similarity is $-1$ for vectors pointing in opposite directions.
-For values of $11>\mu>-$ 1, the distribution forms a hollow ring around the anchor vector $\bm{{\alpha}}$.
+This corresponds to the von Mises-Fisher (vMF) distribution (Appendix D).
+For $\mu=-1$, the distribution is the on opposite side of the hypersphere, as the cosine similarity is 1-1 for vectors pointing in opposite directions.
+For values of $1>\mu>-1$, the distribution forms a hollow ring around the anchor vector $\bm{{\alpha}}$.
 This occurs because the cosine similarity for these $\mu$ values expects the activating vectors to point in a different direction than the anchor, without specifying the direction.
@@
 This potential capability is also a major difference to the vMF distribution.
+One drawback of the HyperPG formulation, is that it cannot be “pushed” to the nearest training sample as with the other formulations.
+However, the prototype’s interpretability can be maintained by following the approach of prototypical concept balls (Ma et al., 2023).
+Instead of only using the pushed training sample to interpret the prototype, all training sample with similar activation are equally good explanations.
+By presenting multiple sample patches the user is able to infer a learned concept.
 HyperPG can be easily adapted to other probability distributions with additional desirable properties.
-Possible candidate distributions are elaborated on in Appendix B.
+Possible candidate distributions are elaborated on in Appendix D.
 Similarly, it is possible to exchange the cosine similarity to other similarity measures or functions, and learn an untruncated PDF over their output, making the HyperPG idea transferable to other manifolds and applications outside of prototype learning.
-4 Training
+4 Training Prototypical Part Networks
+Prototypical Part networks like ProtoPNet have a reputation for being difficult to tune and hyperparameter-sensitive.
+To better explore the design choices, we categorize them into three groups: (1) Network architecture governs information flow from input to the output, including pretrained backbone network and neck for feature extraction.
+(2) Prototype hyperparameters shape the prototypical latent space through the choice of prototype formulation, but also number of prototypes, prototype dimensions and additional shaping losses.
+(3) Network optimization settings such as warmup-epochs, learning rate schedules or number of optimizers.
+ProtoPNet and its successors often rely on complex optimization schemes with warm-up phases and stage-specific optimizers and schedules.
+Detailed hyperparameter values are provided in subsection B.2.
+4.1 Network Architecture
+For our experiments, we use the ProtoPNet’s architecture (Chen et al., 2019) as the foundational deep prototype learning model for image classification.
+Figure 4 illustrates this network architecture with the HyperPG prototype module.
+The encoding stage employs a pretrained backbone (e.g., ResNet (He et al., 2016)) .
+A subsequent projection neck consisting of two $1\times 1$ convolution layers with ReLU between them and Sigmoid activation at the end, reduce dimensionality to create the latent prototype space.
+In the prototype stage, similarity scores are computed.
+Original ProtoPNet uses Euclidean prototypes measuring $L_{2}$ distance between encoded features and prototype parameters.
+The HyperPG module, shown in Figure 4, first calculates cosine similarity between learnable anchors $\bm{{\alpha}}$ and neck-produced latent vectors.
+Then, a Density Estimation layer with parameters $\mu$ and $\sigma$ computes Gaussian PDF over the previous layer’s activations.
+Both components could be independently modified in future work, implementing alternative similarity measures (e.g., hyperbolic) or multi-modal probability distributions.
+The classification stage produces output logits from similarity scores through a single linear layer, preserving interpretability.
+4.1.1 Comparison to Recent Works
+Our experiments use the basic ProtoPNet architecture, focusing on the effect of different prototype formulations.
+This section briefly outlines modifications required to implement the models in Table 1.
+LucidPPN (Pach et al., 2024) replaces the encoding stage with a hybrid encoder for color-only and texture-only feature extraction.
+Deformable ProtoPNet (Donnelly et al., 2021) modifies the prototype module’s input to allow dynamic spatial arrangements (“deformations”).
+ProtoPool (Rymarczyk et al., 2022) alters the prototype module’s output by learning to dynamically pool prototype activations.
+ProtoTree (Nauta et al., 2021) and ProtoKNN (Ukai et al., 2023) use a different inherently interpretable models for the classification stage.
+ProtoPFormer (Xue et al., 2024) utilizes a Vision Transformer backbone with separate network branches for global information (class token) and local information (image tokens).
+4.2 Prototype Losses
 The original ProtoPNet implementation uses three loss functions: a task specific loss like crossentropy for classification, a cluster loss to increase compactness within a class’s cluster, and a separation loss to increase distances between different prototype clusters.
-4.1 Prototype Losses
 ProtoPNet defines a cluster loss function to shape the latent space such that all latent vectors ${\bm{z}}_{c}\in{\bm{Z}}_{c}$ with class label $c$ are clustered tightly around the semantically similar prototypes ${\bm{p}}_{c}\in{\bm{P}}_{c}$.
@@
 $L_{\mathrm{Sep}}=\frac{1}{N}\sum_{i=1}^{N}\frac{1}{|C|}\sum_{c\in C}\max_{{\bm{p}}_{\neg c}\notin{\bm{P}}_{c}}\max_{{\bm{z}}_{c,i}\in{\bm{Z}}_{c,i}}s({\bm{p}}_{\neg c},{\bm{z}}_{c,i}),$ (8)
-The $L_{\mathrm{Sep}}$ function punishes high similarity values between a latent vector ${\bm{z}}_{c}$ of class c and prototypes ${\bm{p}}_{\neg c}$ not belonging to $c$, thereby separating the clusters in latent space.
-Please note, the original ProtoPNet paper uses a slightly different notation by working with the $L_{2}$ distance, instead of a similarity measure.
+The $L_{\mathrm{Sep}}$ function punishes high similarity values between a latent vector ${\bm{z}}_{c}$ of class c and prototypes ${\bm{p}}_{\neg c}$ not belonging to $c$, thereby separating the clusters in latent space.Please note, the original ProtoPNet implementation is optimized for minimizing the $L_{2}$ distance, instead of maximizing the similarity measure.
+Using the $L_{2}$ similarity would require tuning of the associated hyperparameters.
 4.3 Multi-Objective Loss Function
@@
 Our experiments uses $\lambda_{\mathrm{Clst}}=0.8$ and $\lambda_{\mathrm{Sep}}=0.08$ as proposed by ProtoPNet (Chen et al., 2019).
-4.3 Network Architecture
-For our experiments, we use the network architecture proposed by ProtoPNet (Chen et al., 2019), which we deem the basic deep prototype learning architecture for image classification.
-Figure 4 illustrates this network architecture with the HyperPG prototype module.
-In the encoding stage, a pretrained image encoder such as a ResNet model (He et al., 2016) serves as model backbone.
-Then, a convolution neck projects the high dimensional feature map of the backbone into the lower dimensional prototype space.
-This neck consists of two $1\times 1$ convolution layers with ReLU activation in between followed by a Sigmoid activation.
-In the prototype stage, the prototype module computes the similarity scores.
-The original ProtoPNet uses Euclidian prototypes, which compute the $L_{2}$ distance between the encoded feature maps and the prototype parameters.
-Figure 4 shows the make up of the HyperPG module: First, the prototype module computes the cosine similarity of the learnable HyperPG anchors $\bm{{\alpha}}$ to the latent vectors produced by the neck.
-Second, a Density Estimation layer with learnable parameters mean $\mu$ and std $\sigma$ computes the Gaussian PDF over the activations of the previous layer.
-Future extensions could adapt both components independently, for example by implementing a hyperbolic similarity measure or a multi-modal probability distribution.
-The final classification stage produces the output logits based on the prototype module’s similarity scores. ProtoPNet uses a single linear layer due to its inherent interpretability.
-4.3.1 Comparison to Recent Works
-Our experiments use only the basic ProtoPNet architecture and focus on the effect of different prototype formulations.
-This section provides a short overview of the required changes to this basic architecture in order to implement the models described in Table 1.
-LucidPPN (Pach et al., 2024) replaces the encoding stage with a hybrid encoder specializing in color-only and texture-only feature extraction.
-Deformable ProtoPNet (Donnelly et al., 2021) makes changes towards the input of the prototype module to allow for dynamic spatial arrangements of the prototypes, i.e., “deformations”.
-ProtoPool (Rymarczyk et al., 2022) makes changes towards the output of the prototype module, by also learning how to dynamically pool the different prototype activations.
-ProtoTree (Nauta et al., 2021) and ProtoKNN (Ukai et al., 2023) propose to use a different inherently interpretable model for the classification stage.
-ProtoPFormer (Xue et al., 2024) builds exclusively on a Vision Transformer backbone and uses different network branches for the global information encoded in the class token and the local information in the image tokens.
+Table 2: Mean Top-1 Test Accuracies (%) CUB-200-2011.
+ProtoPNet Optim. Simplified Optim. Add. Datasets (R50)
+Method R50 D121 R50* D121* ViT* Cars* Flowers*
+Baseline $79.2$ p m 0.2 $75.5$ p m 0.2 $79.2$ p m 0.2 $75.5$ p m 0.2 $78.5$ p m 0.8 $85.2$ $90.0$
+Euclidean (PPN) $73.9$ p m 0.4 $76.4$ p m 1.2 $61.4$ p m 1.4 $57.8$ p m 3.0 $23.4$ p m 0.4 $75.4$ $70.7$
+Gaussian $6.4$ p m 0.3 $75.7$ p m 0.2 $61.1$ p m 0.2 $45.9$ p m 5.1 $4.2$ p m 1.3 $73.0$ $74.8$
+Cosine $78.5$ p m 0.2 $80.4$ p m 0.2 $71.7$ p m 0.4 $68.7$ p m 0.5 $69.2$ p m 3.9 $79.6$ $85.5$
+HyperPG (Ours) $77.8$ p m 0.2 $78.7$ p m 0.1 $74.3$ p m 0.6 $70.7$ p m 0.3 $72.9$ p m 0.1 $79.0$ $87.8$
+*Without warm-up epochs and training optimizations. R50: ResNet50, D121: DenseNet121
+4.4 Network Optimization
+ProtoPNet proposes a sophisticated network optimization scheme to achieve its high performance.
+Initially, for five warm-up epochs, only the convolution neck and prototype module are trained using Adam (Kingma and Ba, 2015) with learning rate $3\text{\times}{10}^{-3}$, while the classification layer uses $1\text{\times}{10}^{-4}$.
+During the warm-up phase, the pretrained image encoder remains frozen.
+After the warm-up phase, a new Adam optimizer trains the encoder with learning rate $1\text{\times}{10}^{-4}$, while the neck and prototype module continue at $3\text{\times}{10}^{-3}$.
+A scheduler reduces learning rates by $1\text{\times}{10}^{-1}$ every 5 epochs.
+A preliminary Bayesian hyperparameter optimization using Weights and Biases (Biewald, 2020) yielded no significant improvements, suggesting the original ProtoPNet implementation is already highly optimized.
 5 Experiments
-The experiments were performed on CUB-200-2011 (CUB) with 200 bird species (Wah et al., 2011) with ResNet50 (He et al., 2016), DenseNet121 (Huang et al., 2017) and ViT-B16 (Dosovitskiy, 2020) backbones.
-Later experiments also explore the Stanford Cars (Krause et al., 2013) and Oxford Flowers (Nilsback & Zisserman, 2008) datasets.
-Implementation details such as data preprocessing or model implementation are provided in Appendix D.
-Table 2: Mean Top-1 Test Accuracies and Std for different prototype formulations with ResNet50 Backbones averaged over 3 random seeds.
-ResNet50 Dense121 ViT-B16
-Baseline 79.2 $\pm$ 0.2 75.5 $\pm$ 0.2 78.5 $\pm$ 0.8
-Euclidian (ProtoPNet) 35.3 $\pm$ 1.4 33.6 $\pm$ 3.0 9.4 $\pm$ 0.4
-Gaussian 61.1 $\pm$ 0.2 45.9 $\pm$ 5.1 4.2 $\pm$ 1.3
-Cosine 71.7 $\pm$ 0.4 68.7 $\pm$ 0.5 69.2 $\pm$ 3.9
-HyperPG (Ours) 74.3 $\pm$ 0.6 70.7 $\pm$ 0.3 72.9 $\pm$ 0.1
-5.1 Quantitative Results
-Table 2 reports the mean top-1 accuracy and standard deviation for the tested models on the CUB dataset with backbone models ResNet50 (He et al., 2016), DenseNet121 (Huang et al., 2017) and ViT-B16 (Dosovitskiy, 2020) pretrained on ImageNet.
-Please note that in contrast to prior work, we do not report scores from model ensembles or using pretrained iNaturalist weights.
-The black-box baselines perform the best across all experiments, which we see as a theoretical upper bound for the performance.
-The classical Euclidean prototypes perform the worst across all experiments.
-A first model improvement can be achieved by switching to probabilistic Gaussian prototypes.
-The hyperspherical Cosine prototypes further improve the performance.
-HyperPG combines the advantages of Gaussian and Cosine prototypes, further closing the gap to the upper bound presented by the black-box baselines.
-Notably, both Euclidian and Gaussian prototypes are unable to learn on the latent features of ViT-B16.
-We theorize that the attention mechanism of the Transformer architecture produces a near-hyperspherical latent space.
-This highlights the importance of choosing the right prototype formulation for the used backbone.
-For example, language aligned models such as (Radford et al., 2021) use the cosine similarity for their alignment.
-We expect that only hyperspherical prototypes can produce meaningful results in this setting.
-Figure 5: CUB-200-2011 Test Accuracy per Epoch with ResNet50 backbone. Mean and std over 3 random seeds.
-Figure 5 shows the test accuracy over epochs for CUB with ResNet50 backbone.
-The standard Euclidean prototypes take the longest time (200 Epochs) until convergence while also achieving the lowest accuracy.
-By changing only the prototype formulation to a Gaussian one, the model achieves higher performance and converges at 100 epochs.
-Both the cosine and HyperPG formulation outperform the other formulations in model accuracy and training time, requiring only 30 training epochs.
-5.2 Ablation: Number of Prototypes
-Figure 6: Ablation: Test Accuracy on CUB with different numbers of prototypes per class.
-We conduct an ablation study over the number of prototypes per class and present the test accuracy scores in Figure 6.
-Notably, the predictive performance of Euclidean and Gaussian prototypes degrades when more prototypes are added.
-The prototype formulations with a hyperspherical structure, Cosine and HyperPG prototypes converge to similar performance levels regardless of the number of prototypes per class.
-5.3 Ablation: Number of Dimensions
-We conduct an ablation study over the number of prototype dimensions and present the test accuracy scores in Figure 7.
-Hyperspherical prototype formulations such as Cosine and HyperPG are robust regarding the choice of prototype dimensions, although HyperPG performs slightly better with increased dimensionality.
-The Gaussian prototypes, like HyperPG prototypes, can adapt their variance $\sigma^{2}$ to deal with the increase in dimensionality.
-The performance of Euclidean prototypes suffers most with changes in dimensionality.
-Figure 7: Ablation: Test Accuracy on CUB with different numbers of prototype dimensions.
-5.4 Ablation: Additional Datasets
-We perform additional experiments on two different datasets.
-Stanford Cars (Cars) with 196 car models (Krause et al., 2013) and Oxford Flowers (Flowers) with 102 different species (Nilsback & Zisserman, 2008).
-The relative differences in performance seem to hold across the different datasets, with Euclidean prototypes performing the worst.
-Interestingly, Gaussian prototypes are able to match the performance of the hyperspherical formulations more closely.
-Table 3: Top-1 Test Accuracy with Resnet50 Backbone.
-Cars Flowers
-Baseline 85.2 90.0
-Euclidean (ProtoPNet) 36.2 15.0
-Gaussian 73.0 74.8
-Cosine 79.6 85.5
-HyperPG (Ours) 79.0 87.8
+Table 2 reports the mean top-1 accuracy and standard deviation for the tested models across three random seeds.
+Unlike prior work, we use only ImageNet weights, omitting model ensembles and pretrained iNaturalist weights.
+We first conduct image classification experiments on CUB-200-2011 (Wah et al., 2011) with the full architecture and training implementation as proposed by ProtoPNet (Chen et al., 2019), then with a simplified network optimization, and finally on two additional image classification datasets.
+Implementation details such as data preprocessing or model implementation are provided in Appendix B.
+5.1 ProtoPNet Optimization
+We experiment on the CUB dataset using ResNet50 (He et al., 2016) and DenseNet121 (Huang et al., 2017) backbones pretrained on ImageNet with ProtoPNet’s full optimization scheme (Chen et al., 2019).
+We also test a blackbox baseline without a prototype module.
+Table 2 (ProtoPNet Optim.) presents the top-1 test accuracies on CUB-200-2011.
+In this experiment, we evaluate different prototype formulations without adjusting hyperparameters.
+Cosine prototypes perform best, even outperforming the Dense121 blackbox baseline.
+While probabilistic formulations perform slightly worse than their point-based counterparts, Gaussian prototypes show high sensitivity to hyperparameters and backbone selection.
+They fail to learn meaningful representations with the ResNet50 backbone.
+5.2 Simplified Optimization
+To evaluate the robustness and practical applicability of different prototype formulations, we test them under a drastically simplified training scheme that removes ProtoPNet’s multi-stage optimization, warm-up epochs, and learning rate scheduling in favor of a single AdamW optimizer (learning rate $1\text{\times}{10}^{-4}$, no scheduling) for the entire model.
+Table 2 (Simplified Optim.) shows the top-1 accuracies on CUB-200-2011.
+The Euclidean prototype formulation shows heavily reduced performance, suggesting that the original ProtoPNet implementation was highly finetuned.
+While hyperspherical prototypes also perform worse, remain competitive with the fully-optimized Euclidean ProtoPNet.
+Figure 5: CUB-200-2011 Test Accuracy per Epoch with ResNet50 backbone and simplified optimization scheme. Mean and std over 3 random seeds.
+Figure 5 illustrates the Top-1 Test Accuracy on CUB-200-2011 over training epochs.
+With the simplified training scheme, the hyperspherical prototype formulations converge much faster than $L_{2}$-based formulations.
+Cosine and HyperPg converge within 20 training epochs, whereas the Euclidean prototypes require over 200 epochs.
+5.3 Ablation: Transformer Backbone
+Using the simplified optimization scheme, we test a Vision Transformer backbone.
+Notably, both Euclidean and Gaussian prototypes fail to learn a latent structure from ViT-B16 features.
+We hypothesize that the Transformer’s attention mechanism produces a near-hyperspherical latent space.
+This underscores the importance of matching prototype formulation to backbone architecture.
+For instance, language aligned models like CLIP (Radford et al., 2021) use the cosine similarity for modality alignment.
+We anticipate that only hyperspherical prototypes can produce meaningful results when using CLIP as a feature extractor.
+5.4 Ablation: Datasets
+We apply our simplified training scheme to two additional datasets: Stanford Cars (Cars) with 196 car models (Krause et al., 2013) and Oxford Flowers (Flowers) with 102 species (Nilsback and Zisserman, 2008).
+Table 2 (Add.
+Datasets) shows top-1 test accuracies.
+The performance pattern among prototype formulations remains consistent across datasets, with Euclidean prototypes consistently underperforming.
 6 Interpretability Analysis
-We perform a qualitative analysis of the different prototype formulations with ResNet50 backbone (see Figure 8).
-One common visualization technique originally proposed by ProtoPNet is a a direct overlay of the prototype activation map (PAM).
-However, this visualization technique has several shortcomings, which we discuss in more detail in Appendix A.
-In order to offer an alternative to the PAM visualization, we also present GradCAM (Selvaraju et al., 2017) overlays.
-Both Euclidian and HyperPG prototypes do not actually learn class-specific prototypes, in contrast to Gaussian or Cosine prototypes.
-Interestingly, all prototype formulations seem to activate in similar regions in the test image, but differ with the known training examples.
-HyperPG could already leverage class-shared prototypes, which would explain its robustness to changes in the number of prototypes per class.
-There seems to be no qualitative difference between the visualizations produced with the various prototype formulations using the CNN based backbones.
-As expected, the ViT backbone (see Figure 9) suffers from the “distraction problem” as observed by ProtoPFormer (Xue et al., 2024).
-We present additional examples in Appendix A.
-Refer to caption Refer to caption
+We qualitatively analyze different prototype formulations with a ResNet50 backbone and the simplified optimization scheme.
+ProtoPNet originally proposed directly overlaying the prototype activation map (PAM).
+However, this visualization technique has several shortcomings, discussed in Appendix A.
+As an alternative, we also present GradCAM (Selvaraju et al., 2017) visualizations in Appendix A.
+Figure 6 shows PAMs for different prototype formulations on the same test image (left), with highest activations marked by yellow rectangles.
+Following the prototypical concepts approach (Ma et al., 2023), we present the two most highly activated training patches, as not all prototype formulations support a “push” onto training samples.
+Interestingly, all prototype formulations activate in similar test image regions but differ in their associated training examples.
+The Euclidean prototype formulation seems to produce less visually consistent prototype associations than the other formulations using the simplified optimization scheme.
+We observe no qualitative differences between visualizations produced by various prototype formulations using CNN-based backbones (see Appendix A).
+While ViT backbone performs well with cosine prototypes, it exhibits the ”distraction problem” (Figure 7) first observed by ProtoPFormer (Xue et al., 2024).
+For an extensive overview of prototype examples, GradCAM activations, and additional backbones, refer to Appendix A.
 (a) Euclidean (R50)
-Refer to caption Refer to caption
 (b) Gaussian (R50)
-Refer to caption Refer to caption
 (c) Cosine (R50)
-Refer to caption Refer to caption
 (d) HyperPG (R50)
-Figure 8: Left: Test Image. Right: 3 closest Training Patches to the Prototype. Top Row: PAM visualization. Bottom Row: GradCAM visualization. R50: Resnet50 Backbone. ViT: ViT-B16 Backbone.
+Figure 6: Left: Test Image. Right: Three closest Training Patches to the Prototype.
 Figure 7: Cosine Protoypes with ViT. The overlayed heatmap show the distraction problem caused by the attention mechanism.
 7 Conclusion
-This work provides comprehensive overview of different prototype formulations, with both probabilistic and point-based measurements, in an Euclidean or hyperspherical latent space.
+This work provides a comprehensive overview of different prototype formulations, with both probabilistic and point-based measurements, in Euclidean or hyperspherical latent spaces.
 We introduce HyperPG, a new prototype representation based on Gaussian distributions on the surface of a hypersphere.
-Our experiments showcase the differences in performance and robustness to hyperparameters across different datasets and backbone architectures for these prototype formulations.
-Our results demonstrate the performance advantage of hypersherical prototypes and probabilistic prototypes in contrast to the commonly used Euclidean formulation, with no loss to the inherent interpretability.
-Future work could explore further refinements to the probabilistic prototype formulations, including adaptive mechanisms for prototype selection or the integration of additional probabilistic models such as Mixture Models or Bayesian approaches like MGProto (Wang et al., 2024b) to further enhance interpretability and performance.
+Our experiments demonstrate substantial differences in robustness across prototype formulations.
+Hyperspherical prototypes maintain competitive performance under drastically simplified training (single optimizer, no warm-up epochs or learning rate scheduling), whereas Euclidean prototypes show severe performance degradation under identical conditions.
+This indicates that the original ProtoPNet’s strong performance depends heavily on extensive hyperparameter tuning, whereas hyperspherical formulations offer greater optimization stability.
+This has practical implications for deployment: hyperspherical prototypes provide improved performance with reduced sensitivity to training configuration, making prototype-based models more accessible for practitioners without extensive hyperparameter search resources.
+Our results show the advantages of hyperspherical and probabilistic prototypes over the commonly used Euclidean formulation, with no loss to inherent interpretability.
+Future work could explore further refinements to probabilistic prototype formulations, including adaptive mechanisms for prototype selection or the integration of additional probabilistic models such as Mixture Models or Bayesian approaches like MGProto (Wang et al., 2024b).
+Acknowledgements
+The research presented in this paper was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 448648559.
 Impact Statement
@@
 Appendix A Extended Interpretability Analysis
+Figure 8: DenseNet121 Euclidean. Top: PAM. Bottom: GradCAM
+Figure 9: DenseNet121 Gaussian. Top: PAM. Bottom: GradCAM
+Figure 10: DenseNet121 Cosine. Top: PAM. Bottom: GradCAM
+Figure 11: DenseNet121 HyperPG. Top: PAM. Bottom: GradCAM
+Figure 12: ViT-B16 Euclidean. Top: PAM. Bottom: GradCAM
+Figure 13: ViT-B16 Gaussian. Top: PAM. Bottom: GradCAM
+Figure 14: ViT-B16 Cosine. Top: PAM. Bottom: GradCAM
+Figure 15: ViT-B16 HyperPG. Top: PAM. Bottom: GradCAM
 A.1 Potential Problems with Prototype Activation Maps
@@
 By modifying parts of the background or adding noise to the image, the activation map can change, most likely due to the spatial pooling in the convolution pyramids.
-Appendix B Adapting HyperPG to other Probability Distributions
+Appendix B Implementation Details
+Table 3: Hyperparameters used in our experiments
+Hyperparameter PPN HyperPG
+Data and Model Configuration
+Image size 224 224
+Number of prototypes per class 10 10
+Training batch size 80 48
+Testing batch size 100 48
+Training Schedule
+Number of training epochs 20 (offline augmentation $\times 30$) 200 (online augmentation)
+Number of warmup epochs 5 N.A.
+Warm Optimizer Learning Rates
+Add-on layers $3\times 10^{-3}$ N.A.
+Prototype vectors $3\times 10^{-3}$ N.A.
+Optimizer Learning Rates
+Features $\times 10^{-4} 1×1$$\times 10^{-4}$
+Add-on layers $3\times 10^{-3}$ $1\times 10^{-4}$
+Prototype vectors $3\times 10^{-3}$ $1\times 10^{-4}$
+Joint learning rate step size 5 N.A.
+Final Layer Training
+Last layer optimizer learning rate $\times 10^{-4} 1×1$$\times 10^{-4}$
+Compute Resources
+GPU Model NVidia RTX 4090 NVidia RTX 3070Ti
+GPU VRAM 24 GB 8 GB
+Our experiments can be divided into two implementations: “PPN”, built on the original ProtoPNet (Chen et al., 2019) implementation published on Github [^2] https://github.com/cfchen-duke/ProtoPNet with a sophisticated network optimization scheme.
+And “HyperPG”, our own implementation [^3] https://github.com/LiXiling/prob-proto built from scratch, originally designed for proposing and testing the HyperPG prototypes.
+B.1 Data Preprocessing
+B.1.1 PPN: Data Preprocessing
+ProtoPNet uses an offline data augmentation process, resulting in 30 training images per class and epoch.
+The input images are cropped to the bounding box annotations.
+The images are augmented by applying Rotation XOR Skewing XOR Shearing, followed by a 50% chance of left-right flipping.
+B.1.2 HyperPG: Data Preprocessing
+In contrast to prior work (e.g.
+Chen et al., 2019; Rymarczyk et al., 2020; Ukai et al., 2023) the experiment used an online augmentation process, resulting in 30 training images per class and epoch.
+The input images were cropped to the bounding box annotations, then resized to a resolution of $224\times 224$.
+The augmentations consisted of RandomPerspective, RandomHorizontalFlip and ColorJitter.
+B.2 Hyperparameters
+For both implementations we provide the full hyperparameters in Table 3. PPN uses the exact hyperparameters as used by ProtoPNet according to the publically available Github implementation.
+Our HyperPG implementation does not use a training scheme with warm-up epochs or learn-rate schedulings, making several hyperparameters not applicable, marked as N.A.
+The batchsizes are optimized for the different hardware configurations.
+B.3 Compute Resources
+The PPN experiments were performed on a computing node with four NVidia RTX 4090 GPUs (24 GB VRAM each), with each model accessing one GPU.
+Please note, that the batch size could have been further increased on this hardware configuration.
+On CUB-200-2011 the training duration was roughly 20 minutes per epoch, with no meaningful difference between prototype formulations.
+The HyperPG experiments with simplified training and online data augmentation were performed on a desktop workstation with a single NVIDIA RTX 3070Ti GPU (8 GB VRAM).
+On CUB-200-2011 the training duration was roughly 40 seconds per epoch, with no meaningful difference between prototype formulations.
+We estimate 160 GPU hours for running the PPN experiments plus 120 GPU hours for the HyperPG experiments.
+With an additional, generous 15% overhead estimation for debugging, analyses, etc.
+we feel confident that this work took less than 320 GPU hours to develop.
+B.4 Model Implementation
+The experiments used various feature encoding backbones such as ResNet50 (He et al., 2016), DenseNet121 (Huang et al., 2017) and ViT-B16 (Dosovitskiy, 2020) with pretrained weights on ImageNet.
+All models use a single linear output layer as a classification head.
+The prototypical networks follow the architecture proposed by ProtoPNet (Chen et al., 2019): After the pretrained feature encoder, a bottleneck consisting of a Convolution Layer, ReLU activation, Convolution Layer and Sigmoid activation reduces the dimensions of the feature map to the prototype space.
+For example, the ResNet50 encoder produces a feature map of $H\times W\times D$ shape $7\times 7\times 2048$.
+The bottleneck produces a dimension reduction to shape $7\times 7\times 128$.
+The prototype modules produce the prototype similarity scores, which are passed through a global max pooling, before being passed to the classification head.
+Appendix C Additional Ablations
+C.1 Ablation: Number of Prototypes
+Figure 16: Ablation: Test Accuracy on CUB with different numbers of prototypes per class.
+We conduct an ablation study over the number of prototypes per class using the simplified training scheme and present the test accuracy scores in Figure 16.
+Notably, the predictive performance of Euclidean and Gaussian prototypes degrades when more prototypes are added.
+The prototype formulations with a hyperspherical structure, Cosine and HyperPG prototypes converge to similar performance levels regardless of the number of prototypes per class.
+C.2 Ablation: Number of Dimensions
+We conduct an ablation study over the number of prototype dimensions using the simplified training scheme and present the test accuracy scores in Figure 17.
+Hyperspherical prototype formulations such as Cosine and HyperPG are robust regarding the choice of prototype dimensions, although HyperPG performs slightly better with increased dimensionality.
+The Gaussian prototypes, like HyperPG prototypes, can adapt their variance $\sigma^{2}$ to deal with the increase in dimensionality.
+The performance of Euclidean prototypes suffers most with changes in dimensionality.
+Figure 17: Ablation: Test Accuracy on CUB with different numbers of prototype dimensions.
+Appendix D Adapting HyperPG to other Probability Distributions
 We define HyperPG prototypes ${\bm{p}}^{H}=(\bm{{\alpha}},\mu,\sigma)$ as a Gaussian Distribution with mean $\mu$ and variance $\sigma^{2}$ of cosine similarities around an anchor vector $\alpha$.
@@
 As early experiments on the CUB-200-2011 dataset showed no significant difference in performance, these sections are relegated to the appendix.
-B.1 Cauchy Distribution
+D.1 Cauchy Distribution
 One theoretical disadvantage of the Gaussian distribution is the fast approach to zero, which is why a distribution with heavier tails such as the Cauchy distribution might be desirable.
@@
 Figure 18: PDF for the Gaussian and Cauchy distribution of cosine similarity values. The Cauchy distribution has heavier tails, avoiding vanishing gradients issues.
-B.2 Truncated Distributions
-The cosine similarity is defined only on the interval $[-1,1]$.
+D.2 Truncated Distributions
+The cosine similarity is defined only on the interval [1,1][-1,1].
 This makes it attractive to also use truncated probability distributions, which are also only defined on this interval.
@@
 The truncated Gaussian pdf $\mathcal{T}_{\mathrm{Gauss}}$ requires the cumulative probability function $\mathcal{G}$ and error function $f_{\mathrm{err}}$, and is defined as
-[…]
+$f_{\mathrm{err}}(x)$ $=\frac{2}{\sqrt{\pi}}\int_{0}^{x}\exp\left(-z^{2}\right)dz,$ (10)
 ${\mathcal{N}}(x;\mu,\sigma)$ $=\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp\left(-\frac{(x-\mu)^{2}}{2\sigma^{2}}\right),$ (11)
@@
 $\mathcal{T}_{\mathrm{Gauss}}(x,\mu,\sigma,a,b)$ $=\frac{{\mathcal{N}}\left(s_{\cos}({\bm{z}}|\bm{{\alpha}});\mu,\sigma\right)}{\mathcal{G}(1,\mu,\sigma)-\mathcal{G}(-1,\mu,\sigma).},$ (13)
-with lower bound $a$ and upper bound $b$, e.g., for the cosine similarity $a=-1$ and $b=1$. Similarly, the truncated Cauchy distribution can be applied, which is defined as
+with lower bound $a$ and upper bound $b$, e.g., for the cosine similarity a=1a=-1 and $b=1$. Similarly, the truncated Cauchy distribution can be applied, which is defined as
 $\mathcal{T}_{\mathrm{Cauchy}}(x,x_{0},\gamma,a,b)=\frac{1}{\gamma}\left(1+\left(\frac{x-x_{0}}{\sigma}\right)^{2}\right)^{-1}\left(\arctan\left(\frac{b-x_{0}}{\gamma}\right)-\arctan\left(\frac{a-x_{0}}{\gamma}\right)\right)^{-1}.$ (14)
@@
 Figure 19: CUB-200-2011 Test Accuracy with HyperPG prototypes using different PDFs.
-B.3 von Mises-Fisher Distribution
+D.3 von Mises-Fisher Distribution
 The von Mises-Fisher distribution (vMF) is the analogue of the Gaussian distribution on the surface of a hypersphere (Hillen et al., 2017).
@@
 with mean vector $\bm{{\alpha}}$, scalar concentration parameter $\kappa$ and normalization constant $C_{d}(\kappa)$.
-The normalization constant $C_{d}(\kappa)$ is a complex function and difficult to compute for higher dimensions, which is why, for example, Tensorflow [^1] Tensorflow API Documentation - Accessed 2025-01-10 only supports the vMF distribution for $D\leq 5$.
+The normalization constant $C_{d}(\kappa)$ is a complex function and difficult to compute for higher dimensions, which is why, for example, Tensorflow [^4] Tensorflow API Documentation - Accessed 2026-01-07 only supports the vMF distribution for $D\leq 5$.
 However, the vMF distribution is a viable similarity measure when using the unnormalized density function with $C_{d}(\kappa)=1$.
@@
 Figure 20: Changing the concentration parameter $\kappa$ is akin to changing HyperPG’s std $\sigma$.
-B.4 Fisher-Bingham Distribution
+D.4 Fisher-Bingham Distribution
 As the vMF distribution is the equivalent of an isotropic Gaussian distribution on the surface of a hypersphere, the Fisher-Bingham (FB) distribution is the equivalent of a Gaussian with full covariance matrix.
 Similar to the vMF, the normalization constant is difficult to compute for higher dimensions, but the unnormalized density function remains feasible.
-For a $D$ dimensional space, the FB distribution is by a $D\times D$ matrix ${\bm{A}}$ of orthogonal vectors $(\bm{{\alpha}}_{1},\bm{{\alpha}}_{2},\ldots,\bm{{\alpha}}_{D})$, concentration parameter $\kappa$ and ellipticity factors $[\beta]_{2:D}$ where $\sum_{j=2}^{D}\beta_{j}=1$ and $0\leq 2|\beta_{j}|<\kappa$.
+For a $D$ dimensional space, the FB distribution is by a $D\times D$ matrix $\bm{A}$} of orthogonal vectors $(\bm{{\alpha}}_{1},\bm{{\alpha}}_{2},\ldots,\bm{{\alpha}}_{D})$, concentration parameter $\kappa$ and ellipticity factors $\beta]_{2:D}$ where $\sum_{j=2}^{D}\beta_{j}=1$ and $0\leq 2|\beta_{j}|<\kappa$.
 The FB’s unnormalized PDF is defined as
@@
 Figure 21: Illustration of the Fisher-Bingham Distribution in $D=3$
-B.5 Mixture Models
+D.5 Mixture Models
 HyperPG’s probabilistic nature lends itself to a mixture formulation.
@@
 First experiments with mixture of HyperPG prototypes did not show any improvement over the standard formulation. However, this might change with other datasets.
-Appendix C Scaled Dot-Product Prototype
+Appendix E Scaled Dot-Product Prototype
 This section does not reiterate the full attention mechanism.
@@
 Figure 22: Illustration of Cosine Similarity and Scaled Dot-Product
-Appendix D Implementation Details
-D.1 Data Preprocessing
-In contrast to prior work (e.g.
-Chen et al., 2019; Rymarczyk et al., 2020; Ukai et al., 2023) the experiment used an online augmentation process, resulting in 30 training images per class and epoch.
-The input images were cropped to the bounding box annotations, then resized to a resolution of $224\times 224$.
-The augmentations consisted of RandomPerspective, RandomHorizontalFlip and ColorJitter.
-D.2 Hyperparameters
-The prototypical part networks ProtoPNet and HyperPGNet use a convolutional neck after the feature encoder and work on a latent feature map of size $7\times 7\times 128$.
-The models were trained with a minibatch size of 48 images using AdamW optimizer with learning rate 1e-4 and weight decay 1e-4.
-D.3 Compute Resources
-All experiments were performed on a desktop workstation with a single NVIDIA RTX 3070Ti GPU (8 GB VRAM) per model.
-On CUB-200-2011 the training duration was roughly 40 seconds per epoch, regardless of the used prototype formulation.
-D.4 Model Implementation
-The experiments used various feature encoding backbones such as ResNet50 (He et al., 2016), DenseNet121 (Huang et al., 2017) and ViT-B16 (Dosovitskiy, 2020) with pretrained weights on ImageNet.
-All models use a single linear output layer as a classification head.
-The prototypical networks follow the architecture proposed by ProtoPNet (Chen et al., 2019): After the pretrained feature encoder, a bottleneck consisting of a Convolution Layer, ReLU activation, Convolution Layer and Sigmoid activation reduces the dimensions of the feature map to the prototype space.
-For example, the ResNet50 encoder produces a feature map of $H\times W\times D$ shape $7\times 7\times 2048$.
-The bottleneck produces a dimension reduction to shape $7\times 7\times 128$.
-The prototype modules produce the prototype similarity scores, which are passed through a global max pooling, before being passed to the classification head.
-Appendix E Parameter Distribution of HyperPG
+Appendix F Parameter Distribution of HyperPG
 To evaluate, if HyperPG actually uses the probabilstic components, we look at the learned mean $\mu$ and standard deviation $\sigma$ values after training on CUB200-2011.
@@
 Figure 23: HyperPG learned mean and sigma values.
+Appendix G Asset Licenses
+Our work relies on the publicly available Github Code of ProtoPNet (Chen et al., 2019) which is released under the MIT license.
+Our experiments use the following datasets:
+  • CUB-200-2011 (Wah et al., 2011) by Caltech Vision Lab. The use is restricted to non-commercial research and educational purposes.
+  • Oxford Flowers (Nilsback and Zisserman, 2008) from the Visual Geometry Group at University of Oxford. No license is detailed.
+  • Stanford Cars (Krause et al., 2013) as published on Kaggle. No license is detailed.